### v1

# UrbanVLP: A Multi-Granularity Vision-Language Pre-Trained Foundation Model for Urban Indicator Prediction

Abstract.
Urban indicator prediction aims to infer socio-economic metrics in diverse urban landscapes using data-driven methods. However, prevalent pre-trained models, particularly those reliant on satellite imagery, face dual challenges. Firstly, concentrating solely on macro-level patterns from satellite data may introduce bias, lacking nuanced details at micro levels, such as architectural details at a place. Secondly, the lack of interpretability in pre-trained models limits their utility in providing transparent evidence for urban planning. In response to these issues, we devise a novel Vision-Language Pre-Trained Model (UrbanVLP) in this paper. Our UrbanVLP seamlessly integrates multi-granularity information from both macro (satellite) and micro (street-view) levels, overcoming the limitations of prior pre-trained models. Moreover, it introduces automatic text generation and calibration, elevating interpretability in downstream applications by producing high-quality text descriptions of urban imagery. Rigorous experiments conducted across six socio-economic tasks underscore UrbanVLP’s superior performance. We also deploy a web platform to verify its practicality.
Y. Liang is the corresponding author. Email: yuxliang@outlook.com
1. Introduction
Urban Indicator Prediction, a scholarly pursuit in the realm of urban computing (Zheng et al., 2014), deploys data-driven methodologies to forecast socio-economic metrics (e.g., GDP, population, carbon emission) in diverse urban landscapes. Rooted in the escalating global prominence of urban environments and the imperative for judicious urban planning, this academic discipline finds motivation in enhancing the efficacy of policymaking, optimizing resource allocation, and mitigating challenges endemic to urban agglomerations (Xi et al., 2022; Yan et al., 2023).
Classical machine learning approaches commonly tackle urban indicator prediction via supervised training, leveraging task-specific datasets like satellite imagery. These tasks span a spectrum from predicting poverty levels (Jean et al., 2016a; Han et al., 2020; Perez et al., 2017; Ayush et al., 2021a), crop yields (You et al., 2017; Wang et al., 2018; Martinez et al., 2021; Yeh et al., 2021), land cover (Uzkent et al., 2019), to commercial activity (He et al., 2018; Liu et al., 2023d). Nevertheless, the inherent task specificity of supervised learning, reliant on ample labeled data, may impede the model’s capacity for broad generalization. This constraint poses a conceivable risk to the overarching robustness and effectiveness of the model.
In contrast, Pre-trained Model, also known as Self-Supervised Learning (Jaiswal et al., 2020), empowers the autonomous discovery and optimization of urban region representations from diverse urban data modalities, encompassing satellite images (Jean et al., 2016b; Ayush et al., 2021b; Yao et al., 2022b; Watmough et al., 2019), Points of Interests (POIs) (Fu et al., 2019; Jenkins et al., 2019; Zhang et al., 2021; Li et al., [n. d.]), road networks (Liang et al., 2018, 2021; Wei Chen, 2023) and human mobility (Wang and Li, 2017; Seong et al., 2023; Li et al., 2023b, a). By obviating the necessity for externally labeled datasets, these pre-training techniques enable models to learn intrinsic region representations, fostering adaptability and scalability in comprehending intricate urban dynamics for enhanced generalization across various urban indicator prediction tasks.
Figure 1. Single-granularity vs. Multi-granularity modeling.
Given the widespread accessibility of satellite imagery through platforms such as Google Maps or Baidu Maps (Li et al., 2022b; Liu et al., 2023d; Xi et al., 2022), coupled with its wealth of information on regional features (e.g., road network patterns, building density, vegetation coverage), a predominant majority of initiatives harness satellite imagery [^1] As other modalities (e.g., mobility, POIs) are always inaccessible due to privacy issues. as the foundational modality for learning urban representations. Despite its promise, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture, storefront). As depicted in Figure 1, while the satellite images (a) and (b) possess remarkably similar layouts, their urban indicators considerably differ. Upon zooming into micro levels using the corresponding street-view images (c) and (d), a more nuanced and fine-grained understanding becomes apparent. Exclusively capturing macro-level patterns through satellite imagery may introduce substantial bias, falling short in furnishing more comprehensive information across diverse urban levels. Hence, the integration and alignment of multi-granularity information (macro-level and micro-level) for urban indicator prediction present an ongoing challenge that requires further exploration.
Furthermore, while urban decision-makers rely on explainable evidence for service deployment and urban planning, the pre-trained models for urban indicator prediction mostly lack interpretability. They typically learn direct mapping from single or multi-modal data to the output representations, which poses a significant challenge in comprehending the intrinsic factors influencing the predictive outcomes. In light of the recent surge in Large Language Models (LLMs) (Ouyang et al., 2022; Zhao et al., 2023), UrbanCLIP (Yan et al., 2023) emerges as an innovative solution. This model effectively integrates textual modality as auxiliary knowledge into the realm of urban indicator prediction. Figure 1 shows how these textual data encapsulate key elements within both satellite images and their street-view counterparts. Notably, even in the absence of image-text pairs, UrbanCLIP adeptly harnesses a myriad of textual descriptions generated by pre-trained image-to-text LLMs, thereby facilitating both explainable reasoning and prediction. However, the text generation process of UrbanCLIP raises several critical concerns. a) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery. b) Homogenization: The generated description appears overly simplified and general, potentially leading to homogenization and inadequate differentiation. For instance, the average length of the text description in UrbanCLIP is quantified at 18.1 words, leading to insufficient information. To better align with the intricacies of satellite (or street-view) imagery, we necessitate a more powerful approach that goes beyond the intuitive text generation method in UrbanCLIP, ensuring a more faithful and explainable representation of urban regions.
In this paper, we present a Vision-Language Pre-Trained Model entitled UrbanVLP for urban indicator prediction. Our model elaborately integrates multi-granularity (macro and micro-level) information from both satellite and street-view imagery to produce comprehensive urban region representations, while simultaneously harnessing the interpretability inherent in high-quality textual descriptions. Targeting the first challenge, we introduce a novel Multi-Granularity Cross-Modal Alignment module. It utilizes dual-branch contrastive learning to establish alignment between information derived from two semantic granularities. Compared to standard contrastive learning applied in the satellite branch at the macro level, we employ token-level contrastive learning to capture fine-grained details in street-view imagery. Concurrently, spatial information, enriched with the location coordinates of street-view images, is encoded and injected to provide essential spatial context. Finally, the fusion process aggregates features from street-view images and location-specific features within the same region, seamlessly incorporating fine-grained information into the satellite branch.
To address the second issue (i.e., lack of interpretability), we devise an Automatic Text Generation together with a Calibration mechanism to uphold text quality standards. Given the impracticality of generating a large number of descriptions using GPT-4V (Yang et al., 2023) or Gemini (Fu et al., 2023) due to prohibitive expense, we utilize a closely comparable GPT4V (Yang et al., 2023) model called ShareGPT4V (Chen et al., 2023a), which is a strong caption model trained on 100k GPT4V (Yang et al., 2023) generated captions. To guarantee the quality of LLM-generated descriptions, we propose a reference-free metric PerceptionScore, motivated by the human evaluation system. This metric comprises two components: Clipscore and Cyclescore, working collaborateively to capture both semantic similarity and detail inclusiveness.
Overall, the main contributions of our work are summarized as:
  • Multi-granularity cross-modal alignment. We explored the role of two different visual data modalities at different semantic granularities - street-view and satellite imagery. We introduced street-view data to inject fine-grained semantic information based on the satellite imagery and provided a framework.
  • Automatic text generation and calibration. Powered by image-to-text LLMs, we generate text descriptions and implement a robust evaluation mechanism, ensuring the fidelity between the generated texts and the corresponding image content.
  • A new benchmark & Empirical evidence. We plan to open-source the first vision-text and multi-granularity urban dataset upon paper notification, consisting of six downstream tasks across the socio-economy. Extensive experiments are conducted on them. As illustrated in Figure 2, UrbanVLP outperforms existing approaches by an average improvement of 3.55% on the $R^{2}$ metric.
Figure 2. $R^{2}$ performance on Beijing and Shenzhen dataset.
Figure 3. Overall framework of our proposed UrbanVLP.
2. Preliminaries
2.1. Formulation
Definition 1 (Urban Region) We follow previous research (Xi et al., 2022; Yan et al., 2023) and evenly divide the target area (e.g. a city) into $L$ urban regions.
Definition 2 (Urban Imagery) Diverse visual images, including satellite and street-view images, facilitate an intuitive analysis of urban regions. Specifically, satellite images provide a coarse-grained overview of the structural composition of a region. Each input satellite image w.r.t the urban area $g$ can be represented as $I^{sa}_{g}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the length and width. Street-view images are captured by vehicles from map providers during urban traversals, offering detailed visual information on the fine-grained aspects within a region. By querying the latitude and longitude range of satellite images, multiple street-view images are associated together w.r.t the urban region $g$ based on their locations $\mathcal{L}_{g}$. Each individual street-view image is represented as $I^{{sv}}_{g}\in\mathbb{R}^{H\times W\times 3}$.
Definition 3 (Text Description) Textual descriptions of urban area $g$ (including satellite text $T^{sa}_{g}$ and street-view text $T^{sv}_{g}$) provide precise means for urban area analysis. Such text can be manually generated or produced using image annotation tools. Particularly, leveraging advanced LLMs enables detailed and insightful analytical descriptions of given area images. Notably, incorporating spatial context (e.g. vegetation coverage and POI) significantly reflects the functional characteristics of the region.
Definition 4 (Urban Indicators) Urban indicators assess various aspects of a region, including population, economy, society, and public services, from multiple perspectives. A set of $L$ urban regions with $K$ indicators is represented as $\mathbf{Y}\in\mathbb{R}^{L\times K}$. In this paper, we comprehensively use population (count), GDP (in million Chinese Yuan), carbon emissions (in tons), night lights (intensity), house prices (in Chinese Yuan), and points of interest, i.ePOI, (count for different categories) as ground-truth urban indicators.
Problem Statement (Urban Region Representation) We next formally define the problem of urban region representation learning. Given a satellite images $I^{sa}_{g}$, a set of street-view images $\mathcal{I}^{sv}_{g}$ belonging to the target coverage, and their corresponding latitude and longitude pair $\mathcal{L}_{g}$. The objective is to employ a learning function $\mathcal{F}$ to map them to representation vectors $\mathbf{e}_{g}=\mathcal{F}(I^{sa}_{g},\mathcal{I}^{sv}_{g},\mathcal{L}_{g})$. The representation can then be further utilized to infer urban indicators $\mathbf{Y}_{g}$ for any given region $g$. We assume that during the training phase, automated additional text descriptions are available.
2.2. Related Work
2.2.1. Urban Region Representation Learning
In literature, many studies have learned task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility (Li et al., 2022c; Liu et al., 2023d; Xi et al., 2022). For example, Jean et al. (2016b) used nighttime light images as weak labels to learn area representation and assess regional economic development, while Urban2Vec (Wang et al., 2020) integrated street-view imagery and POI data to learn neighborhood embeddings. Some contrastive learning approaches like PG-SimCLR (Xi et al., 2022) and UrbanCLIP (Yan et al., 2023) have shown success in representing urban regions through satellite image. However, the aforementioned works exclusively considered a single type of urban imagery and overlooked the potential synergy between satellite imagery and street-view imagery, which can complement each other. Recently, some efforts (Barbierato et al., 2020; Huang et al., 2023a; Li et al., 2022c; Cao et al., 2023, 2018a; Law et al., 2019; Cao et al., 2018b; Huang et al., 2023b) have begun to explore the simple combination of satellite imagery and street-view imagery. For instance, Vision-LSTM (Huang et al., 2023b) integrates diverse street-level images into the satellite image domain using an LSTM-based network for precise urban village identification. Li et al. (2022c) propose utilizing street segments for region modeling and predict socio-economic indicators by merging street-view and satellite images through a multi-stage pipeline. Nevertheless, prior studies overlook distinct features in satellite and street-view images for varied granularities. Additionally, few studies incorporate text as a supervisory signal.
2.2.2. Large Multimodal Model
As LLMs undergo rapid evolution, a faction within the research community is increasingly focused on integrating visual knowledge into these models, such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021). Early studies (Li et al., 2022a, 2023c) enhanced CLIP with refined data strategies for more diverse datasets, proving effective for fundamental visual tasks (Li et al., 2022d; Liu et al., 2023c; Zhang et al., 2022) but less so for complex tasks like visual question answering. Recently, new models like MiniGPT-4 (Chen et al., 2023b), LLaVA (Liu et al., 2023b), and InstructBLIP (Dai et al., 2023) have improved the understanding of complex problems by expanding parameters and training data, called large multimodal models (LMMs). In this paper, we comprehensively evaluate the text generation quality of representative LMMs model and propose an automated text description generation and calibration technique to ensure effective image description capabilities with LMMs.
2.2.3. Vision-Language Pre-training (VLP)
VLP aims to jointly encode vision and language in a fusion model. Early works on VLP can be broadly categorized into single-stream (Chen et al., 2020; Kim et al., 2021; Li et al., 2020) and two-stream (Tan and Bansal, 2019; Lu et al., 2019) methodologies, with the former using a unified architecture for modality integration and the latter employing separate encoders before merging modalities. Milestone work CLIP (Radford et al., 2021) and its variants (Li et al., 2021; Yao et al., 2022a) highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval. Recent works (Tsimpoukelli et al., 2021; Alayrac et al., 2022) shift towards leveraging LLMs knowledge for vision-language representation learning. In urban indicator prediction, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information. Our groundbreaking work introduces multiple types of text descriptions as supplements, aiming to learn more comprehensive and interpretable urban region representations.
3. Methodology
We present an overview of our proposed framework in Figure 3. There are two principal stages in our framework:
  • Stage 1: We first devise an automatic text generation and calibration module. Leveraging ShareGPT4V (Chen et al., 2023a) (current state-of-the-art LMM captioner model), textual descriptions are created for each street-view image, incorporating geographical and visual prompts. PerceptionScore is also introduced to ensure high-quality text retention from both semantic and visual perspectives. Building upon the foundation of high-quality image-text pairs, we further develop a multi-granularity cross-modal alignment framework, achieving multi-granular alignment and multimodal correlation through multi-level contrastive learning and fine-grained information injection.
  • Stage 2: In the urban indicator prediction stage, we leverage the frozen encoder pre-trained in Stage 1 to extract input data features. Subsequently, a lightweight multi-layer perceptron (MLP) is fine-tuned to yield precise urban indicator predictions. Additionally, optional information such as street-view imagery and geographical coordinates is considered.
3.1. Automatic Text Generation and Calibration
Text Generation. For each satellite image, we initiate a range query to obtain the corresponding set of available street-view images and their respective geospatial coordinates within the designated area. Existing works (Lai et al., 2023; Fan et al., 2023a; Yan et al., 2023) proved that the quality of the text is crucial for enhancing performance. Therefore, advanced LMMs, owing to their robust cross-modal knowledge storage capabilities, offer detailed textual comprehension and descriptions for each street-view image. However, this confronts two primary challenges. Firstly, due to substantial financial implications associated with API usage, employing closed source LMMs like GPT4V (Yang et al., 2023) or Gemini (Fu et al., 2023) for generating tens of thousands of text descriptions is impractical. Secondly, street-view images often encompass diverse detailed information, necessitating a well-designed prompt template for obtaining high-quality textual descriptions.
To address these issues, we first utilize a closely related GPT4V model, known as ShareGPT4V (Chen et al., 2023a), a powerful language model with vision capabilities trained on 100k GPT4V generated captions, to generate street-view image descriptions. In the process of designing text prompt templates, the proportion of segmented elements in street-view images reflects the distribution of elements, aiding LMMs in evaluating the importance of each element. To achieve this, we employ a pre-trained segmentation model to decouple visual elements and calculate the proportion of segmentation for each element. Simultaneously, urban and geospatial coordinates of street-view images are included as prompts to guide the expression of additional urban information. Finally, the prompt template is formulated as follows: Analyze the street-view panoramic image in [City] in a comprehensive and detailed manner. The coordinate of the street-view image is [Longitude, Latitude]. The segmentation ratio of the street-view image is [Segment Information] . More prompt examples and segmentation ratio details can be found in Appendix C.
Text Calibration. In contrast to previous research (Yan et al., 2023) that employed simplistic rules and manual refinement processes for text description of satellite images, street view images often contain rich fine-grained information, leading to potential hallucination issues (Rawte et al., 2023) in generated text descriptions that require further calibration. Rule-based methods are overly general and fail to address specific issues in each image description as they are tailored to particular cases. On the other hand, manual rewriting is labor-intensive and demands significant effort. Therefore, an effective method for evaluating text quality is crucial.
[…]
$\begin{split}&I^{\prime}=\operatorname{SDXL}(I),\\ &\texttt{CycleScore}(I,I^{\prime})=1-\operatorname{MAE}(\operatorname{Seg}(I),\operatorname{Seg}(I^{\prime})),\\ &\texttt{PerceptionScore}(I,T)=(\texttt{CLIPScore}(I,T)+\mathop{\texttt{CycleScore}(I,I^{\prime})})/2.\end{split}$
Figure 4. The procedure of Cycle Score calculation.
3.2. Multi-Granularity Cross-modal Alignment
Modality Representation. Utilizing automated generation techniques, we obtain a dataset of high-quality image-text pairs, represented as $D=(I,T)$. Here, $I$ denotes either a satellite image $I^{sa}_{g}$ or a street view image $I^{sv}_{g}$, while $T$ represents either a satellite text description $T^{sa}_{g}$ or a street view text description $T^{st}_{g}$.
we first deploy Vision Transformer (ViT) (Dosovitskiy et al., 2021) as an image encoder to process images derived from satellites and street-views, proficient in deconstructing images into patches for feature extraction. In particular, we reshape the input image $I\in\mathbb{R}^{H\times W\times 3}$ into a sequence of flattened 2D patches ${I}_{P}\in\mathbb{R}^{N\times({P^{2}}\cdot{C})}$, where C is channel number, $(P,P)$ is the resolution of each image patch, and $N=HW/P^{2}$ is the patch number, which is the input sequence length of ViT. Notebly, the ViT utilizes a consistent latent vector size (denoted as $d$) across all its layers. Thus, we can obtain the outputs as the patch embeddings through a trainable linear projection. We also prepend a learnable embedding to the patch embedding sequence (i.e., $\mathbf{z}_{0}^{0}={I}_{cls}$). Besides, we add learnable position embedding $\mathbf{E}_{pos}$ to retain positional information. The process is as follows:
(1) $\mathbf{z}_{0}=[{I}_{cls};{I}_{p}^{1}\mathbf{W};{I}_{p}^{2}\mathbf{W};\cdots;{I}_{p}^{N}\mathbf{W}]+\mathbf{E}_{pos},$
where $\mathbf{W}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$ and $\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times d}$ are learnable parameter. $\mathbf{z}_{0}$ is the linear projection result and then sent to ViT.
Consistent with the standard Transformer encoder, ViT includes alternating layers of multi-head self-attention $\operatorname{MSA(\cdot)}$ operation (Vaswani et al., 2017). Furthermore, layer normalization (LN) is adopted before every block, whereas residual connections are after every block:
(2) $\mathbf{z}_{l}=\mathrm{MSA}(\mathrm{LN}(\mathbf{z}_{l-1}))+\mathbf{z}_{l-1},\quad l=1\ldots L.$
Finally, we use the output $\mathbf{z}_{l}^{0}$ corresponding to the special tokens ${I}_{cls}$ inserted at the beginning as the final output of the image encoder, rewrite as $\mathbf{z}_{I}$ for image representation.
Concurrently, we use basic Transformer-Encoder (Vaswani et al., 2017) as a text encoder by default to comprehend and encode the semantics of textual descriptions in parallel. The text encoder follows a similar $\operatorname{MSA}$ mechanism of the visual encoder. The input text sequence is bracketed with $[SOS]$ and $[EOS]$ tokens, and the activation of the highest layer of Transformer at $[EOS]$ token is considered the global representation $\mathbf{z}_{T}$ of text.
Moreover, depending on whether it is a satellite image-text pair or a street-view image-text pair, we categorize them into two branches with separate visual and text encoders – namely, the satellite branch and the street-view branch.
Furthermore, considering the longitude and latitude information corresponding to street view images, which aids in identifying the geospatial location of the area, we leverage the open-source GeoCLIP’s (Vivanco et al., 2023) location encoder. This encoder captures longitude and latitude information at different resolutions, generating semantically rich high-dimensional features. Here, a set of longitude and latitude $\mathcal{L}_{g}$ is aggregated into a single vector, denoted as $\mathbf{z}_{L}$, thereby incorporating geographical positioning.
Modality Alignment. Our goal is to capture the coarse-grained global regional information from the satellite branch and the fine-grained local positional information from the street-view branch.
For capturing global information, even though we have obtained overall visual features for the region from images in the satellite branch, we can further incorporate fine-grained information from images in the street-view branch. Specifically, the visual features of streets within each satellite region are aggregated, simultaneously fused with the corresponding geographical feature, aiding in learning a robust and comprehensive representation of urban areas. The formalized process is as follows:
(3) $\mathbf{z}_{g_{sa}}=\textit{f}~{}(\mathbf{z}_{I_{sa}},~{}\operatorname{Aggr}(\mathbf{z}_{I_{sv}}^{1},\mathbf{z}_{I_{sv}}^{2},...,\mathbf{z}_{I_{sv}}^{m}),~{}\operatorname{Aggr}(\mathbf{z}_{L}^{1},\mathbf{z}_{L}^{2},...,\mathbf{z}_{L}^{m})),$
where $\mathbf{z}_{I_{sa}}$ represents the satellite visual features for the queried region, $\mathbf{z}_{I_{sv}}^{k}$ denotes the $k$-th street-view visual feature in that region, and $\mathbf{z}_{L}^{k}$ represents the $k$-th street view location feature in the same region. The variable $m$ represents the count of street views and their corresponding locations within the region. $\operatorname{Aggr}(\cdot)$ denotes the aggregation method for all street views in the region, utilizing additive aggregation in this case. The function f  signifies the feature fusion method.
To establish a global-level semantic correspondence between satellite images and their related descriptions, we employ a joint optimization approach for the image and text modal encoders in the satellite branch. For the $i$-th satellite image-report pair $(I_{g_{i}},T_{g_{i}})$ in a mini-batch, we regard two modality data as queries and keys alternatively to learn the correct image-report pairings, to maximally preserve the mutual information between the true pairs in latent space. The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms: $\mathcal{L}_{{CG}}^{Image\rightarrow Text}$ and $\mathcal{L}_{{CG}}^{Text\rightarrow Image}$, which measure the similarity between the image and text embeddings, respectively, defined as:
(4) $\mathcal{L}_{{CG}}^{Image\rightarrow Text}=-\sum_{i}^{N}{\log\frac{\exp(I_{g_{i}}^{\top}T_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(I_{g_{i}}^{\top}T_{g_{j}}/\tau)}}},$
(5) $\mathcal{L}_{{CG}}^{Text\rightarrow Image}=-\sum_{i}^{N}{\log\frac{\exp(T_{g_{i}}^{\top}I_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(T_{g_{i}}^{\top}I_{g_{j}}/\tau)}}},$
(6) $\mathcal{L}_{CG}=\frac{1}{N}(\mathcal{L}_{{CG}}^{Image\rightarrow Text}+\mathcal{L}_{{CG}}^{Text\rightarrow Image}),$
where $I_{g_{i}}$ and $T_{g_{j}}$ are the normalized embedding of satellite image representation $\mathbf{z}_{g_{sa}}$ in the $i$-th pair and that of text representation $\mathbf{z}_{T_{sa}}$ in the $j$-th pair, respectively. Besides, $N$ is the batch size, and $\tau$ is the temperature for contrastive learning.
For capturing local information, given that the street-view local branch is expected to capture fine-grained information to provide ground-level details for region embedding, approaches like (Radford et al., 2021; Jia et al., 2021) rely exclusively on the similarity of each modality’s global feature, thus lacking the capability to capture finer-level information such as the relationship between visual objects and textual tokens (Yao et al., 2022a). We utilize a fine-grained interaction mechanism to implement alignment. Specifically, we utilize token-level maximum similarity between visual and textual tokens to direct the contrastive objective. Initially, we compute the similarity between each visual token and all textual tokens, and then leverage the maximum value to calculate the average similarity of all image tokens to textual tokens. The similar approach is also applied to text-to-image.
(7) $\operatorname{SIM}(v_{i},t_{i})=\frac{1}{l_{1}}(\sum_{k_{1}=1}^{l_{1}}v_{ik_{1}}^{\top}{\operatorname{argmax}_{k_{2}\in[0,l_{2})}(t_{ik_{2}})}),$
(8) $\operatorname{SIM}(t_{i},v_{i})=\frac{1}{l_{2}}(\sum_{k_{2}=1}^{l_{1}}t_{ik_{1}}^{\top}{\operatorname{argmax}_{k_{2}\in[0,l_{1})}(v_{ik_{2}})}),$
where $v_{i}$ denotes the $i$-th token of visual features and $t_{j}$ represents the $j$-th token of textual features. $l_{1}$ and $l_{2}$ denote the token length of visual and textual features in street-view branch respectively. The fine-grained token-level representation can be optimized through:
(9) $\mathcal{L}_{{CL}}^{Image\rightarrow Text}=-\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(v_{i}^{\top},t_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(v_{i}^{\top},t_{j})/\tau)}}},$
(10) $\mathcal{L}_{{CL}}^{Text\rightarrow Image}=-\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(t_{i}^{\top},v_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(t_{i}^{\top},v_{j})/\tau)}}},$
(11) $\mathcal{L}_{CL}=\frac{1}{N}(\mathcal{L}_{{CL}}^{Image\rightarrow Text}+\mathcal{L}_{{CL}}^{Text\rightarrow Image}).$
3.3. Urban Indicator Prediction
Pre-training Stage. The total objective of UrbanVLP can be defined as the sum of the above two losses:
(12) $\mathcal{L}_{Total}=\alpha\mathcal{L}_{{CG}}+\beta\mathcal{L}_{{CL}}.$
where $\alpha$ and $\beta$ are weight hyperparameter. Through backpropagation optimization, We achieve multi-granularity cross-modal urban image-text input data alignment, resulting in a robust encoder.
Prediction Stage. During the fine-tuning stage, we employ a linear probing approach (He et al., 2022). This process begins by extracting frozen features $\mathbf{e}_{sa},\mathbf{e}_{sv},\mathbf{e}_{t}$ from the pre-trained encoder for satellite images, street view images, and street view positions (The latter two are optional). Subsequently, these features are fused, and a minimalist classifier (MLP) is trained on top to fine-tune the prediction of urban metrics, denoted as $\mathbf{Y}_{i}=MLP(\mathbf{e}_{sa},\mathbf{e}_{sv},\mathbf{e}_{t})$. As depicted in Figure 3, these encoders have adapted to capturing high-level abstractions from extensive visual data, enabling the MLP to manipulate static features and study the model’s inference capabilities.
4. Experiments
In this section, we conduct extensive experiments to investigate the following Research Questions (RQ):
  • RQ1: Can UrbanVLP outperform previous baseline methods and generalize well to various urban indicator tasks?
  • RQ2: How does each component (e.g., satellite branch, street-view branch, location encoding branch) contribute to UrbanVLP?
  • RQ3: How do automated text generation and calibration affect the generated text quality?
  • RQ4: How do we envision the practicality of UrbanVLP?
Given the limits of page number, RQ4 is detailed in Appendix D.
4.1. Datasets & Tasks
Our dataset comprises a dual-category structure encompassing both satellite and street-view image components, each paired with corresponding textual descriptions. For the satellite image-text pairs, we follow and refine the setting of (Yan et al., 2023) to utilize core area data of 4 cities in China.
4.1.1. UrbanSV Dataset
We construct the UrbanSV Dataset by street-view geo-tagged images collected from Baidu Map API and generated text descriptions by LMM.
  • Street-View Panoramic Image Collection. The Street-View image sampling process utilized the Baidu Maps API [^3] https://lbsyun.baidu.com/, facilitating the extraction of road network data within our predefined sampling area. Using the ArcGIS [^4] https://www.esri.com/en-us/arcgis/products/develop-with-arcgis/overview tool, we exported the road network data and established a sampling interval of 500 meters, which allowed for the comprehensive collection of street-view images at every half-kilometer point along the road network, ensuring thorough coverage that complements the satellite imagery’s scope. Following the road network sampling, panoramic street-view images were collected based on the sampling points. These images, with a resolution of 2048x664, also far exceed the 256x256 resolution of satellite imagery, ensuring clarity and ample detail for further analysis. As is shown in Table 1, a total of 47,582 street-level images were collected, whose data volume is 2.7 times that of the satellite dataset, capturing the intricacies of the urban landscape and providing a street-level view that enhances the overall urban analysis.
  • Corresponding Text Description Generation. The textual data corresponding to each satellite image is generated from ShareGPT4V (Chen et al., 2023a), which is recognized for its superior text generation capabilities compared to other contemporary open-source LMM models (Liu et al., 2023a; Dai et al., 2023; Ye et al., 2023; Gao et al., 2023; Chen et al., 2023b, a), offering comprehensive and high-quality textual information. The generated quality can be guaranteed by our perception scores. We filter out the texts with PerceptionScores less than 0.5.
Table 1. Urban imagery dataset statistics.
Dataset Coverage #Street -view Image #Satellite Image
Bottom-left
Top-right
Beijing 39.75°N, 116.03°E 40.15°N, 116.79°E 14,164 4,592
Shanghai 30.98°N, 121.10°E 31.51°N, 121.80°E 15,616 5,244
Guangzhou 22.94°N, 113.10°E 23.40°N, 113.68°E 7,635 3,402
Shenzhen 22.45°N, 113.75°E 22.84°N, 114.62°E 10,167 4,324
4.1.2. Downstream Task Dataset
In our research, the Downstream Task Dataset plays a crucial role in gauging the practical implications of our multimodal framework. We collect six representative indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI. Each indicator illuminates a significant dimension of urban vitality and sustainability. The detailed introduction of each downstream indicator is provided in the Appendix A.
Table 2. Urban indicators prediction results in four datasets. The best results are in bold and the second-best results are underlined.
Methods UrbanVLP StructualUrban (Li et al., 2022c) Vision-LSTM (Huang et al., 2023a) UrbanCLIP-SV (Yan et al., 2023) UrbanCLIP (Yan et al., 2023) PG-SimCLR (Xi et al., 2022) ResNet-18 (He et al., 2016) Autoencoder (Kramer, 1991)
Metric $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
Beijing Carbon 0.787 0.457 0.353 0.745 0.500 0.397 0.692 0.502 0.421 0.504 0.697 0.536 0.716 0.528 0.528 0.430 0.797 0.632 0.383 0.821 0.598 0.294 0.851 0.572
Population 0.725 0.513 0.404 0.699 0.536 0.422 0.649 0.583 0.511 0.447 0.727 0.574 0.661 0.569 0.452 0.476 1.228 0.963 0.251 0.874 0.657 0.164 0.918 0.673
GDP 0.586 0.650 0.416 0.560 0.670 0.418 0.536 0.690 0.454 0.208 0.899 0.553 0.555 0.674 0.439 0.270 1.679 1.067 0.242 1.253 0.733 0.165 1.521 0.838
Night Light 0.531 0.629 0.394 0.471 0.668 0.447 0.409 0.712 0.466 0.352 0.733 0.476 0.420 0.700 0.457 0.367 0.728 0.404 0.337 0.740 0.581 0.276 0.784 0.643
House Price 0.506 0.640 0.479 0.494 0.647 0.481 0.470 0.655 0.493 0.453 0.672 0.512 0.503 0.641 0.482 0.341 0.718 0.555 0.284 0.721 0.576 0.332 0.714 0.576
POI 0.319 0.706 0.383 0.277 0.712 0.385 0.236 0.718 0.388 0.227 0.724 0.410 0.253 0.731 0.388 - - - 0.209 0.747 0.415 0.164 0.759 0.427
Shanghai Carbon 0.716 0.529 0.392 0.677 0.556 0.431 0.568 0.613 0.458 0.574 0.473 0.453 0.671 0.569 0.426 0.270 0.742 0.551 0.243 0.763 0.528 0.230 0.771 0.532
Population 0.593 0.607 0.471 0.548 0.630 0.486 0.521 0.657 0.518 0.402 0.793 0.590 0.536 0.648 0.507 0.288 1.051 0.831 0.274 0.833 0.632 0.166 0.898 0.742
GDP 0.310 0.816 0.595 0.305 0.841 0.613 0.304 1.012 0.664 0.315 1.214 0.741 0.326 0.807 0.587 0.270 1.679 1.067 0.241 1.237 0.740 0.197 1.511 0.844
Night Light 0.457 0.667 0.459 0.412 0.689 0.494 0.355 0.731 0.546 0.276 0.763 0.554 0.387 0.709 0.511 0.228 0.596 0.466 0.202 0.789 0.573 0.164 0.810 0.617
POI 0.413 0.673 0.363 0.387 0.681 0.375 0.247 0.699 0.389 0.243 0.707 0.399 0.269 0.704 0.394 - - - 0.221 0.723 0.411 0.186 0.729 0.425
Guangzhou Carbon 0.698 0.514 0.385 0.654 0.550 0.506 0.551 0.564 0.462 0.423 0.520 0.478 0.585 0.603 0.444 0.413 0.951 0.716 0.362 0.673 0.536 0.254 0.733 0.570
Population 0.665 0.565 0.441 0.638 0.586 0.453 0.617 0.615 0.506 0.497 0.685 0.612 0.633 0.592 0.467 0.294 1.046 2.112 0.249 0.834 0.678 0.195 0.928 0.753
GDP 0.436 0.764 0.541 0.419 0.791 0.549 0.405 1.011 0.625 0.321 1.004 0.689 0.440 0.762 0.546 0.263 2.162 1.553 0.234 1.147 0.732 0.176 1.463 0.811
Night Light 0.577 0.573 0.418 0.534 0.596 0.442 0.405 1.011 0.625 0.335 0.672 0.514 0.483 0.633 0.478 0.436 0.549 0.427 0.225 0.720 0.562 0.176 0.798 0.620
POI 0.396 0.752 0.344 0.325 0.771 0.364 0.143 0.822 0.411 0.136 0.823 0.415 0.150 0.819 0.408 - - - 0.108 0.920 0.433 0.084 0.956 0.469
Shenzhen Carbon 0.659 0.568 0.418 0.613 0.570 0.434 0.564 0.576 0.471 0.483 0.586 0.488 0.562 0.571 0.483 0.234 0.975 0.746 0.207 0.735 0.597 0.189 0.887 0.634
Population 0.790 0.448 0.343 0.724 0.476 0.384 0.624 0.562 0.471 0.589 0.604 0.483 0.727 0.510 0.392 0.294 1.046 2.112 0.264 0.871 0.694 0.175 0.921 0.772
GDP 0.532 0.676 0.448 0.489 0.695 0.471 0.462 0.720 0.517 0.433 0.754 0.577 0.508 0.693 0.464 0.294 1.046 2.112 0.182 1.245 0.836 0.119 1.733 0.884
Night Light 0.457 0.667 0.459 0.421 0.694 0.502 0.404 0.702 0.508 0.320 0.717 0.531 0.387 0.709 0.511 0.454 0.199 0.158 0.243 0.726 0.543 0.166 0.758 0.608
POI 0.461 0.759 0.347 0.321 0.780 0.374 0.254 0.795 0.396 0.147 0.849 0.440 0.185 0.838 0.427 - - - 0.126 0.968 0.469 0.092 0.973 0.479
$1^{\text{st}}$ount 53 0 0 1 5 3 0 0
Table 3. Ablation study on UrbanSV Beijing dataset. The best results are in bold.
Model Carbon Population GDP Night Light House Price POI
$\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
w/o SV branch 0.754 0.491 0.387 0.690 0.545 0.428 0.547 0.680 0.433 0.468 0.670 0.445 0.520 0.635 0.465 0.263 0.714 0.387
w/o ST branch 0.518 0.687 0.512 0.465 0.715 0.559 0.229 0.887 0.543 0.423 0.698 0.423 0.295 0.764 0.575 0.265 0.716 0.390
w/o LE branch 0.776 0.469 0.358 0.717 0.520 0.414 0.589 0.652 0.409 0.518 0.633 0.400 0.508 0.639 0.472 0.301 0.710 0.385
w/o SV&ST branch 0.716 0.528 0.324 0.626 0.598 0.425 0.258 0.871 0.511 0.482 0.661 0.402 0.493 0.654 0.513 0.259 0.721 0.395
UrbanVLP 0.788 0.455 0.455 0.720 0.517 0.410 0.585 0.651 0.420 0.523 0.634 0.398 0.523 0.634 0.407 0.319 0.706 0.383
4.2. Experimental Setup
4.2.1. Baselines
We compare our method with 7 baselines in the field of urban imagery-based socioeconomic prediction. The single granularity imagery-based urban region representation learning methods include: AutoEncoder (Kramer, 1991), ResNet-18 (He et al., 2016), PG-SimCLR (Xi et al., 2022), UrbanCLIP (Yan et al., 2023), UrbanCLIP-SV (Yan et al., 2023). We also apply multiple baselines for multi-granularity urban imagery-based representation learning: Vision-LSTM (Huang et al., 2023b) and StructualUrban
(Li et al., 2022c)
. The in-depth overview of each baseline method is presented in the Appendix B.
4.2.2. Metrics and Implementation
The performance of our predictions is evaluated using a trio of standard metrics: the coefficient of determination denoted as $R^{2}$, the root mean squared error (RMSE), and the mean absolute error (MAE). An increase in $R^{2}$, along with a reduction in RMSE and MAE values, is indicative of enhanced model accuracy. Adam optimizer is chosen to minimize the training loss during parameter learning. A grid search on hyperparameters is conducted, where search ranges for learning rate and batch size are set as 2e-6, 2e-5, 2e-4, 2e-3, 2e-2 and 4, 8, 16, 32, 64, respectively. The data augmentation strategy follows the setting of (Radford et al., 2021). The weight hyperparameter $\alpha$ and $\beta$ of $\mathcal{L}_{Total}$ are set to 0.5, 0.5 respectively. We run all the models on NVIDIA A800 GPUs.
We implement the aggregation method $Aggr$ through MLP layers. To ensure a fixed length for network processing, we set the max street-view sequence length within one as 25, and sequences shorter than that are achieved by padding with zeros. The fusion method is Addition. We utilize the current state-of-the-art GeoCLIP (Vivanco et al., 2023) as our location encoder. For the semantic segmentation model to decouple visual elements with street-view images, the architecture for the semantic model is ResNet18dilated + PPM_deepsup. The original dataset contains 150 categories, from which we select 38 categories and group them into 13 street-view features following. (Fan et al., 2023b). The detailed categories can be seen in Appendix C.
4.3. RQ1: Performance Evaluation
To further validate the effectiveness of the proposed UrbanVLP, we conduct comparisons with existing state-of-the-art methods using all datasets. The comprehensive results are presented in Table 2. As we can see, UrbanVLP significantly surpasses the performance of baseline methods employing both single-granularity and multi-granularity imagery, achieving a substantial margin of improvement. It outperforms the best baseline, StructualUrban (Li et al., 2022c), by 4.2%, 2.6%, 2.6%, 6.0%, 1.2% and 4.7% in terms of $R^{2}$ for all 6 indicators: Carbon, Population, GDP, Night Light, House Price and POI in Beijing. The similar performance improvement can also be seen in the dataset of the other three cities. Besides, the average performance gain of UrbanVLP on RMSE and MAE in Beijing data are 2.3% and 2.0%, respectively. In order to provide a more intuitive display, we have created radar Image in Figure 2 to visualize the performance of two cities, Beijing and Shenzhen. The results further prove the effectiveness of our multi-granularity framework for urban region representation learning.
It is evident that pipelines leveraging multi-granularity approaches typically demonstrate superior performance compared to those employing a single granularity level. This enhancement can be rationalized by the incorporation of additional fine-grained information derived from street-view imagery. UrbanVLP exhibits superior performance over StructuralUrban (Li et al., 2022c) and Vision-LSTM (Huang et al., 2023b), illustrating the efficacy of textual mobility, which enhances interpretability. Concurrently, it is observed that UrbanCLIP-SV significantly underperforms compared to UrbanCLIP, demonstrating that street-view imagery, while rich in fine-grained details, lacks macroscopic perspective, leading to suboptimal outcomes when used in isolation. This also underscores the irreplaceability of satellite imagery as a macroscopic visual modality.
4.4. RQ2: Ablation Studies
We conduct ablation studies to investigate the effectiveness of different components in UrbanVLP on UrbanSV Beijing dataset, including the Satellite branch (ST branch), Street-View branch (SV branch), and Location Encoding branch (LE branch). The results are demonstrated in Table 3. It can be observed that the incorporation of street-view imagery results in an average improvement of 3.75% in terms of $R^{2}$. Furthermore, while street-view imagery contributes fine-grained information, its individual impact is still outstripped by that of satellite imagery, aligning with the performance differential between UrbanCLIP and UrbanCLIP-SV in Section 4.3. The role of location coordinates has been underexplored in previous work. Although not as impactful as visual modalities, their inclusion has yielded an average improvement of 0.5% in terms of $R^{2}$.
In our work, we leverage ShareGPT4V (Chen et al., 2023a) for text description generation. Here we conduct experiments on current open-sourced LMM models, and compute the PerceptionScore of them. The results are presented in Table LABEL:table:lmms. As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which are used as the Image-to-Text models (Yan et al., 2023).
Table 4. Perception score of representative LMMs.
Models Average Generated-Description Scores
LLaVA-v1.5 (Liu et al., 2023a) 0.674
mPLUG-Owl (Ye et al., 2023) 0.662
LLaMA-Adapter V2 (Gao et al., 2023) 0.698
MiniGPTv2 (Chen et al., 2023b) 0.693
ShareGPT4V (Chen et al., 2023a) 0.714
Figure 5. Illustration of the quality of our generated description example. The prompt for (c) is :”From scale 1 to 10, decide how similar the image is to the description”, whereas the prompt for (d) is: ”Generate a panoramic street-view image based on the description”.
Figure 6. Perception Score distribution on UrbanSV dataset for four representative cities.
4.5. RQ3: Qualitative Analysis
To visually illustrate the quality of the generated description, we present an example in Figure 5. The generated description of the street-view image (a) is depicted in (b). Subsequently, we employ GPT4V (Yang et al., 2023) to assess the quality score of the generated description, which is rated as 7 out of 10, indicating the effectiveness of the generated text. Furthermore, GPT4V also provides specific areas for improvement in (c). In (d), we utilize GPT4V to generate an image based on the description in (b). It is evident that the generated image bears a resemblance to the original one. In order to ascertain the efficacy of ShareGPT4V, we computed the distribution of PerceptionScores for descriptions generated by ShareGPT4V, as depicted in Figure 6. Notably, only less than 0.02% of image-text pairs scored below 0.5, thereby evidencing the high quality of the descriptions produced.
Figure 7. Examples of predictions from UrbanCLIP and our proposed UrbanVLP.
Here we also show some predicted results in Figure 7 on UrbanSV Beijing dataset. As we can see, satellite image (a) and (b), despite sharing similar layouts, (a) encompasses residential and campus areas, while (b) pertains to industrial zones. Their socio-economic characteristics are entirely different, which could be reflected by socio-enconomic indicators. The carbon emissions and GDP of campus and residential area are significantly lower than those of industrial park, whereas the disparities in population and housing prices are not as pronounced, possibly due to the unique characteristics of school district housing near educational institutions. Night light intensity is similar across the two area due to the fact that the majority of nocturnal economic activities within factories occur indoors, unlike in Central Business Districts (CBDs) where extensive illumination is generated. The same principle applies to school and residential area as well. As observed, the predictions made by UrbanCLIP do not effectively differentiate between the two, with the downstream metrics in Figure 7 (b) still inclined towards predicting a homogenized socio-economic functional area similar to that of (a). In contrast, the results from UrbanVLP are capable of distinctly distinguishing the socio-economic attributes of the two in terms of carbon emission, GDP, and house price.
5. Conclusion and future work
Urban region representation learning plays a significant role in helping us understand the patterns, structures, and dynamics of cities. UrbanVLP, for the first time, explores the differences between street-view and satellite images from a semantic granularity perspective, as well as their roles in modeling urban region representation. A text generation and calibration mechanism is also proposed to ensure high-quality description generation. It has achieved state-of-the-art results on the constructed UrbanSV and downstream dataset. In the future, the further research direction can be two-fold: Add additional modalities such as POIs to enhance information richness and enable the design of better model architectures such as MAE (He et al., 2022) to efficiently utilize existing data.
Appendix A Detailed introduction of downstream dataset
We provide the detailed introduction and data source of our downstream benchmark.
  • Carbon Emissions: Extracted from the Open-source Data Inventory for Anthropogenic CO2 (ODIAC), this dataset offers granular carbon emission figures in 2022 that are in alignment with the spatial resolution of our satellite imagery, which is one square kilometer per image. The unit is ton.
  • Population: Sourced from WorldPop, this data provides population distribution data in 2020. The unit is #citizens.
  • GDP: This data reflects the economic development situation in China in terms of Gross Domestic Product (GDP) figures to add an economic layer to our urban region analysis.
  • Night Light: Nighttime light can, to some extent, characterize the intensity of human activities, which is of great significance for urban development studies. we take the nightlight data from (Zhong et al., 2022) in 2020.
  • House Price: The price of housing is influenced by a variety of socio-economic factors. By observing changes in housing prices, it is possible to discern, to a certain extent, the state of economic development. We collect data from Anjuke (https://www.anjuke.com) in 2023.3. The unit is yuan/$m^{2}$. We take the average of housing prices within an urban region.
  • POI: The distribution of different types of POIs (Points of Interest) within urban areas can serve as a social indicator, as they reflect the social and economic structure, population activities, and lifestyles of the city. We collect POIs (location and category) from the web. The data has 14 categories: Dining and Cuisine, Leisure and Entertainment, Sports and Fitness, Business and Residential, Healthcare, Financial Institutions, Tourist Attractions, Lifestyle Services, Shopping and Consumption, Automobile Related, Hotel Accommodation, Transport Facilities, Science, Education and Culture, Companies and Enterprises. We take aggregation of the number of POIs of each category within each region.
We construct this benchmark to comprehensively evaluate the learned urban region representation. We uniformly partition the downstream dataset into train:val:test=6:2:2. These indicators constitute the downstream task dataset, covering economic (GDP, Night Light, House Price), social (Population, POI), and environmental (Carbon Emission) indicators. The data preprocessing setting adheres to (Yan et al., 2023) with refinement.
Appendix B In-depth overview of baseline methods
The single granularity imagery-based urban region representation learning methods include:
  • AutoEncoder (Kramer, 1991): An architectural framework for neural networks that is designed to extract representations from unlabeled satellite imagery with reconstruction loss.
  • ResNet-18 (He et al., 2016): The well-known residual neural network pre-trained on ImageNet (Deng et al., 2009), which can directly extract visual features of satellite imagery with the knowledge acquired from natural imagery.
  • PG-SimCLR (Xi et al., 2022): A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning. We don’t use PG-SimCLR to predict POI for a fair comparison in our study.
  • UrbanCLIP (Yan et al., 2023): A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
  • UrbanCLIP-SV (Yan et al., 2023): Original UrbanCLIP is pre-trained on satellite image-text pairs. Here we also verify its performance on our proposed UrbanSV dataset, utilizing the same aggregation method to acquire regional features.
We also apply multiple baselines for multi-granularity urban imagery-based representation learning:
  • Vision-LSTM (Huang et al., 2023b): A vision long short-term memory neural network which could process variable-length street-view image within each region. The integration of satellite and street-view imagery enables a more effective capture of comprehensive urban representation, achieving reliable recognition results.
  • StructualUrban(Li et al., 2022c): The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images. We term this method StructualUrban for simplicity.
Appendix C Segmentation Ratio and Prompt Examples
As illustrated in Figure 8, we demonstrate examples of the segmentation ratio of street-view images in our UrbanSV dataset. The segmentation ratio is decoupled by pre-trained segmentation models and is used in both prompt construction and perception score computation. We group the segmentation categories into 13 categories following (Fan et al., 2023b): Light Vehicle; Façade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature. Figure 9 illustrates some exmaples of our prompts used for description generation.
Figure 8. Segmentation Ratio.
Figure 9. Prompt Examples.
Appendix D RQ4: Practicality
Figure 10. UrbanVLP system.
To demonstrate the practical applications of our model in real-life scenarios, we have developed a system bearing the same name UrbanVLP. The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban indicator framework using the Mapbox platform (Mapbox, [n. d.]). This system renders a visual representation of urban areas through satellite imagery, complemented by street-level photographs for an interactive exploration experience. Figure 10 showcases the user-friendly interface of the map, highlighting functionalities that facilitate zooming, searching for locations, and navigating through various sectors.
Each urban region captured by a satellite image, encompassing a 1km by 1km area, is marked by a blue dot, which is encircled by yellow points that denote the locations of street-view images accessible within that region. Engaging with these points allows users to uncover a suite of urban metrics, such as carbon emissions, population, GDP, night light, house price, and POI. The system’s visual components are augmented by a captioning feature that provides clear and informative textual descriptions, offering insights into the spatial characteristics of the designated area.
Furthermore, a street-view panel allows users to peruse individual street-view images alongside pertinent captions. The system also facilitates the discovery of popular POIs within any given area, enriching the understanding of its functional characteristics. Ultimately, UrbanVLP is designed to distill complex urban data into a visual and user-friendly format, thereby offering a holistic and detailed view of urban landscapes and their key indicators.

### v2

@@
-UrbanVLP: A Multi-Granularity Vision-Language Pre-Trained Foundation Model for Urban Indicator Prediction
-Abstract.
-Urban indicator prediction aims to infer socio-economic metrics in diverse urban landscapes using data-driven methods.
-However, prevalent pre-trained models, particularly those reliant on satellite imagery, face dual challenges.
+UrbanVLP: Multi-Granularity Vision-Language Pretraining for Urban Region Profiling
+Abstract
+Urban region profiling aims to learn a low-dimensional representation of a given urban area while preserving its characteristics, such as demographics, infrastructure, and economic activities, for urban planning and development.
+However, prevalent pretrained models, particularly those reliant on satellite imagery, face dual challenges.
 Firstly, concentrating solely on macro-level patterns from satellite data may introduce bias, lacking nuanced details at micro levels, such as architectural details at a place.
-Secondly, the lack of interpretability in pre-trained models limits their utility in providing transparent evidence for urban planning.
-In response to these issues, we devise a novel Vision-Language Pre-Trained Model (UrbanVLP) in this paper.
-Our UrbanVLP seamlessly integrates multi-granularity information from both macro (satellite) and micro (street-view) levels, overcoming the limitations of prior pre-trained models.
+Secondly, the lack of interpretability in pretrained models limits their utility in providing transparent evidence for urban planning.
+In response to these issues, we devise a novel framework entitled UrbanVLP based on Vision-Language Pretraining.
+Our UrbanVLP seamlessly integrates multi-granularity information from both macro (satellite) and micro (street-view) levels, overcoming the limitations of prior pretrained models.
 Moreover, it introduces automatic text generation and calibration, elevating interpretability in downstream applications by producing high-quality text descriptions of urban imagery.
-Rigorous experiments conducted across six socio-economic tasks underscore UrbanVLP’s superior performance.
-We also deploy a web platform to verify its practicality.
-Y. Liang is the corresponding author. Email: yuxliang@outlook.com
-1. Introduction
-Urban Indicator Prediction, a scholarly pursuit in the realm of urban computing (Zheng et al., 2014), deploys data-driven methodologies to forecast socio-economic metrics (e.g., GDP, population, carbon emission) in diverse urban landscapes.
-Rooted in the escalating global prominence of urban environments and the imperative for judicious urban planning, this academic discipline finds motivation in enhancing the efficacy of policymaking, optimizing resource allocation, and mitigating challenges endemic to urban agglomerations (Xi et al., 2022; Yan et al., 2023).
-Classical machine learning approaches commonly tackle urban indicator prediction via supervised training, leveraging task-specific datasets like satellite imagery.
-These tasks span a spectrum from predicting poverty levels (Jean et al., 2016a; Han et al., 2020; Perez et al., 2017; Ayush et al., 2021a), crop yields (You et al., 2017; Wang et al., 2018; Martinez et al., 2021; Yeh et al., 2021), land cover (Uzkent et al., 2019), to commercial activity (He et al., 2018; Liu et al., 2023d).
-Nevertheless, the inherent task specificity of supervised learning, reliant on ample labeled data, may impede the model’s capacity for broad generalization.
-This constraint poses a conceivable risk to the overarching robustness and effectiveness of the model.
-In contrast, Pre-trained Model, also known as Self-Supervised Learning (Jaiswal et al., 2020), empowers the autonomous discovery and optimization of urban region representations from diverse urban data modalities, encompassing satellite images (Jean et al., 2016b; Ayush et al., 2021b; Yao et al., 2022b; Watmough et al., 2019), Points of Interests (POIs) (Fu et al., 2019; Jenkins et al., 2019; Zhang et al., 2021; Li et al., [n.
-d.]), road networks (Liang et al., 2018, 2021; Wei Chen, 2023) and human mobility (Wang and Li, 2017; Seong et al., 2023; Li et al., 2023b, a).
-By obviating the necessity for externally labeled datasets, these pre-training techniques enable models to learn intrinsic region representations, fostering adaptability and scalability in comprehending intricate urban dynamics for enhanced generalization across various urban indicator prediction tasks.
-Figure 1. Single-granularity vs. Multi-granularity modeling.
-Given the widespread accessibility of satellite imagery through platforms such as Google Maps or Baidu Maps (Li et al., 2022b; Liu et al., 2023d; Xi et al., 2022), coupled with its wealth of information on regional features (e.g., road network patterns, building density, vegetation coverage), a predominant majority of initiatives harness satellite imagery [^1] As other modalities (e.g., mobility, POIs) are always inaccessible due to privacy issues.
-as the foundational modality for learning urban representations.
-Despite its promise, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture, storefront).
-As depicted in Figure 1, while the satellite images (a) and (b) possess remarkably similar layouts, their urban indicators considerably differ.
-Upon zooming into micro levels using the corresponding street-view images (c) and (d), a more nuanced and fine-grained understanding becomes apparent.
-Exclusively capturing macro-level patterns through satellite imagery may introduce substantial bias, falling short in furnishing more comprehensive information across diverse urban levels.
-Hence, the integration and alignment of multi-granularity information (macro-level and micro-level) for urban indicator prediction present an ongoing challenge that requires further exploration.
-Furthermore, while urban decision-makers rely on explainable evidence for service deployment and urban planning, the pre-trained models for urban indicator prediction mostly lack interpretability.
-They typically learn direct mapping from single or multi-modal data to the output representations, which poses a significant challenge in comprehending the intrinsic factors influencing the predictive outcomes.
-In light of the recent surge in Large Language Models (LLMs) (Ouyang et al., 2022; Zhao et al., 2023), UrbanCLIP (Yan et al., 2023) emerges as an innovative solution.
-This model effectively integrates textual modality as auxiliary knowledge into the realm of urban indicator prediction.
-Figure 1 shows how these textual data encapsulate key elements within both satellite images and their street-view counterparts.
-Notably, even in the absence of image-text pairs, UrbanCLIP adeptly harnesses a myriad of textual descriptions generated by pre-trained image-to-text LLMs, thereby facilitating both explainable reasoning and prediction.
+Rigorous experiments conducted across six urban indicator prediction tasks underscore its superior performance.
+1 Introduction
+Urban Region Profiling, a scholarly endeavor within the domain of GeoAI [45, 54] and smart city [85], refers to the process of transforming urban regions into a low-dimensional vector representation using deep neural networks.
+Rooted in the escalating global prominence of urban environments and the imperative for urban planning, this academic discipline aims to capture and encode the multifaceted attributes of urban areas, such as demographic information, economic activity, environmental factors, infrastructure, and social dynamics, into a format that can be easily analyzed and utilized by computational models [73, 74].
+Given the widespread accessibility of satellite imagery on platforms such as Google Maps [42, 51, 73], coupled with its wealth of information on regional features (e.g., road networks, building density, and vegetation coverage), a predominant majority of initiatives harness satellite imagery as the foundational modality for learning urban region representations [88].
+Through a comprehensive review of the current literature, we summarize mainstream methods for urban region profiling into two categories: a) Task-specific supervised learning commonly realizes urban region profiling in a fully supervised task, e.g., identifying poverty levels [21, 59, 4, 5], crop yields [80, 70, 63, 57, 53, 78], and commercial activeness [24, 51].
+However, its inherent task specificity (i.e., reliant on ample labeled data), may impede the model’s capacity for broad generalization to other downstream tasks.
+b) Vision pretraining, also known as a type of Self-Supervised Learning [28], aims to learn general visual features from urban satellite imagery, which are subsequently fine-tuned on a specific task for enhancing performance  [29, 6, 73, 51].
+The inclusion of more auxiliary data, e.g., Points of Interests (POIs) [19, 29, 82, 40], road networks [46, 47, 12] and human mobility [71, 64, 44, 39], leads to richer, more accurate, and more useful representations of urban areas.
+For clarity, Figure 1(a-b) depict a sketch of these two streams of approaches.
+Despite the success of using satellite imagery for urban region profiling, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture).
+Previous approaches in Figure 1(a-b) mainly focus on single granularity, neglecting finer-grained visual clues.
+Nonetheless, as the saying goes, “the devil is in the details.”.
+By zooming into micro levels using corresponding street-view images, a more nuanced and fine-grained understanding emerges, as shown in the left hand side of Figure 1(d).
+Therefore, the integration and alignment of multi-granularity information, both macro-level and micro-level, remain ongoing challenges in urban region profiling that require further exploration.
+Figure 1: Frameworks for urban region profiling.
+Compared to existing literature, we present the first attempt to introduce multi-granular visual information and high-quality calibrated texts for this task.
+Furthermore, while urban planners rely on explainable evidence for service deployment and urban planning, the vision-pretraining-based models for urban region profiling mostly lack interpretability.
+They typically learn a mapping from single/multi-modal data to the representations, posing a significant challenge in comprehending the intrinsic factors influencing the prediction outcomes.
+In light of the recent surge in Large Language Models (LLMs) [58, 84, 55], UrbanCLIP [74] stands out as an innovative solution, which effectively integrates textual modality as auxiliary knowledge into urban region profile, as depicted in Figure 1(c).
 However, the text generation process of UrbanCLIP raises several critical concerns.
-a) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery.
-b) Homogenization: The generated description appears overly simplified and general, potentially leading to homogenization and inadequate differentiation.
-For instance, the average length of the text description in UrbanCLIP is quantified at 18.1 words, leading to insufficient information.
+i) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery.
+ii) Homogenization: The generated descriptions appear overly simplified and general, potentially leading to homogenization and inadequate differentiation.
 To better align with the intricacies of satellite (or street-view) imagery, we necessitate a more powerful approach that goes beyond the intuitive text generation method in UrbanCLIP, ensuring a more faithful and explainable representation of urban regions.
-In this paper, we present a Vision-Language Pre-Trained Model entitled UrbanVLP for urban indicator prediction.
-Our model elaborately integrates multi-granularity (macro and micro-level) information from both satellite and street-view imagery to produce comprehensive urban region representations, while simultaneously harnessing the interpretability inherent in high-quality textual descriptions.
-Targeting the first challenge, we introduce a novel Multi-Granularity Cross-Modal Alignment module.
-It utilizes dual-branch contrastive learning to establish alignment between information derived from two semantic granularities.
-Compared to standard contrastive learning applied in the satellite branch at the macro level, we employ token-level contrastive learning to capture fine-grained details in street-view imagery.
-Concurrently, spatial information, enriched with the location coordinates of street-view images, is encoded and injected to provide essential spatial context.
-Finally, the fusion process aggregates features from street-view images and location-specific features within the same region, seamlessly incorporating fine-grained information into the satellite branch.
+In this paper, we present a Vision-Language Pretraining framework (i.e., UrbanVLP) for urban region profiling.
+As depicted in Figure 1(d), our model elaborately integrates multi-granularity information from both satellite (macro-level) and street-view (micro-level) imagery to produce comprehensive urban region representations, while simultaneously harnessing the interpretability inherent in high-quality textual descriptions.
+Targeting the first challenge, we introduce a novel Multi-Granularity Cross-Modal Alignment module, which utilizes dual-branch contrastive learning to establish alignment between information derived from two semantic granularities.
 To address the second issue (i.e., lack of interpretability), we devise an Automatic Text Generation together with a Calibration mechanism to uphold text quality standards.
-Given the impracticality of generating a large number of descriptions using GPT-4V (Yang et al., 2023) or Gemini (Fu et al., 2023) due to prohibitive expense, we utilize a closely comparable GPT4V (Yang et al., 2023) model called ShareGPT4V (Chen et al., 2023a), which is a strong caption model trained on 100k GPT4V (Yang et al., 2023) generated captions.
-To guarantee the quality of LLM-generated descriptions, we propose a reference-free metric PerceptionScore, motivated by the human evaluation system.
-This metric comprises two components: Clipscore and Cyclescore, working collaborateively to capture both semantic similarity and detail inclusiveness.
-Overall, the main contributions of our work are summarized as:
+Given the impracticality of generating a large number of descriptions using GPT4V [75] or Gemini [18] due to prohibitive expense, we utilize a closely comparable GPT4V [75] model called ShareGPT4V [11].
+To guarantee the quality of LLM-generated descriptions, we propose a reference-free metric called PerceptionScore, motivated by the human evaluation system.
+In summary, our major contributions lie in the following aspects:
   • Multi-Granularity Cross-Modal Alignment.
-We explored the role of two different visual data modalities at different semantic granularities - street-view and satellite imagery.
-We introduced street-view data to inject fine-grained semantic information based on the satellite imagery and provided a framework.
+We explore the role of two distinct visual data modalities at various semantic granularities – satellite imagery and street-view imagery.
+We integrate street-view data to inject fine-grained semantic information through token-level contrastive learning.
+By aggregating and fusing local features within a specific region with satellite imagery data, we facilitate the development of a robust and comprehensive representation of urban landscapes.
   • Automatic Text Generation and Calibration.
-Powered by image-to-text LLMs, we generate text descriptions and implement a robust evaluation mechanism, ensuring the fidelity between the generated texts and the corresponding image content.
-  • A new benchmark & Empirical evidence.
-We plan to open-source the first vision-text and multi-granularity urban dataset upon paper notification, consisting of six downstream tasks across the socio-economy.
-Extensive experiments are conducted on them.
-As illustrated in Figure 2, UrbanVLP outperforms existing approaches by an average improvement of 3.55% on the $R^{2}$ metric.
-Figure 2. $R^{2}$ performance on Beijing and Shenzhen dataset.
-Figure 3. Overall framework of our proposed UrbanVLP.
-2. Preliminaries
-2.1. Formulation
-Definition 1 (Urban Region) We follow previous research (Xi et al., 2022; Yan et al., 2023) and evenly divide the target area (e.g. a city) into $L$ urban regions.
+Powered by image-to-text LLMs, we generate text descriptions and implement a robust evaluation mechanism based on a newly-proposed reference-free metric, ensuring the fidelity between the generated texts and the corresponding image content.
+  • Empirical Evidence.
+We evaluate the proposed UrbanVLP on six urban indicator prediction tasks (e.g., GDP, population, and carbon emission) in diverse urban landscapes.
+Extensive experiments demonstrate that our UrbanVLP outperforms existing approaches by an average improvement of 3.95% on the $R^{2}$ metric, while preserving good interpretability based on generated text.
+  • New Benchmark and System.
+We will open-source the first multi-granular vision-text benchmark ($\mathtt{CityScape}$) for urban region profiling upon paper notification, encompassing 6 downstream tasks.
+Moreover, we have deployed a web platform to verify the model’s practicality (see Appendix I).
+2 Related work
+We formally define the problem of urban region profiling.
+Given a satellite images $I^{sa}_{g}$, a set of street-view images $\mathcal{I}^{sv}_{g}$ belonging to the target coverage, and their corresponding latitude and longitude pair $\mathcal{L}_{g}$.
+The objective is to employ a learning function $\mathcal{F}$ to map them to representation vectors $\mathbf{e}_{g}=\mathcal{F}(I^{sa}_{g},\mathcal{I}^{sv}_{g},\mathcal{L}_{g})$.
+The representation can then be further utilized in downstream tasks, such as inferring urban indicators $\mathbf{Y}_{g}$ for any given region $g$.
+We assume that during the training phase, automated additional text descriptions are available.
+The detailed formulation and definitions related to this task are provided in Appendix A.
+Next, we introduce the related work from three aspects.
+Urban Region Profiling.
+In literature, many studies have focused on learning task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility [41, 51, 73].
+For example, Urban2Vec [72] integrated street-view imagery and POI data to learn neighborhood embeddings.
+Some contrastive learning approaches like PG-SimCLR [73] and UrbanCLIP [74] have shown success in representing urban regions through satellite images.
+However, the aforementioned works exclusively considered a single type of urban imagery, overlooking the potential synergy between satellite imagery and street-view imagery, which can complement each other.
+Recently, some efforts [7, 27, 41, 9, 33, 8, 27] have begun to explore the simple combination of satellite imagery and street-view imagery.
+Vision-LSTM [27] integrates diverse street-level images into the satellite image domain using an LSTM-based network for precise urban village identification.
+Li et al.
+[41] propose utilizing street segments for region modeling and predicting socio-economic indicators by merging street-view and satellite images through a multi-stage pipeline.
+Nevertheless, prior studies mostly overlook distinct features in satellite and street-view images for varied granularities.
+To this end, we propose a unified multi-granularity cross-modal alignment framework for comprehensive urban representation in this study.
+Vision-Language Pretraining (VLP).
+VLP aims to jointly encode vision and language in a fusion model.
+Early works on VLP can be broadly categorized into single-stream [13, 31, 35] and two-stream [66, 52] methodologies, with the former using a unified architecture for modality integration and the latter employing separate encoders before merging modalities.
+Milestone work CLIP [61] and its variants [43, 76] highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval.
+Recent works [67, 2] shift towards leveraging LLMs knowledge for vision-language representation learning.
+In urban region profiling, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information.
+Our groundbreaking work introduces multiple types of text descriptions as supplements, aiming to learn more comprehensive and interpretable urban region representations.
+Large Multimodal Model (LMM).
+As LLMs undergo rapid evolution, a faction within the research community is increasingly focused on integrating visual knowledge into these models, such as CLIP [61] and ALIGN [30].
+Early studies [36, 37] enhanced CLIP with refined data strategies for more diverse datasets, which have proven effective for fundamental visual tasks [38, 50, 81] but have shown limitations for complex tasks such as visual question answering.
+Recently, new models like MiniGPT-4 [10], LLaVA [49], and InstructBLIP [14] have improved the understanding of complex problems by expanding parameters and training data, called large multimodal models (LMMs).
+In this paper, we comprehensively evaluate the text generation quality of representative LMMs model and propose an automated text description generation and calibration technique to ensure effective image description capabilities with LMMs.
+Figure 2: Overall framework of our proposed UrbanVLP.
+3 Methodology
+We present an overview of our proposed framework in Figure 2, which is decoupled into two stages.
+In the pretraining stage, we devise an automatic text generation and calibration module using ShareGPT4V [11] to generate textual descriptions for street-view images with geographical and visual prompts.
+Introducing the PerceptionScore to ensure high-quality text, we then develop a multi-granularity cross-modal alignment framework through multi-level contrastive learning and fine-grained information injection.
+In the fine-tuning stage, we use frozen encoders from Stage 1 to extract data features, fine-tune a lightweight MLP for precise predictions, and incorporate optional street-view imagery and geographical coordinates to enrich the analysis.
+3.1 Automatic Text Generation and Calibration
+Text Generation.
+For each street-view image, we utilize advanced Large Multimodal Models (LMMs) due to their robust cross-modal knowledge storage capabilities, to provide comprehensive textual descriptions.
+However, this confronts two primary challenges.
+Firstly, due to substantial financial costs associated with API usage, employing closed source LMMs like GPT4V [75] or Gemini [18] for generating tens of thousands of text descriptions is impractical.
+Secondly, street-view images often encompass diverse detailed information, necessitating a well-designed prompt template for obtaining high-quality textual descriptions.
+To address these issues, we first utilize a closely related GPT4V model, known as ShareGPT4V [11], to generate street-view image descriptions.
+This model is a powerful language model with vision capabilities trained on 100k GPT4V generated captions, demonstrating captioning capabilities comparable to GPT4V.
+In the process of designing text prompt templates, the proportion of segmented elements in street-view images reflects the distribution of elements, aiding LMMs in evaluating the importance of each element.
+To achieve this, we employ a pretrained segmentation model to decouple visual elements and calculate the proportion of segmentation for each element.
+Simultaneously, urban and geospatial coordinates of street-view images are included as prompts to guide the expression of additional urban information.
+Template and segmentation ratio details can be found in Appendix F.
+Text Calibration.
+As existing works [32, 16, 74] demonstrated that the quality of the textural descriptions is crucial for enhancing model performance.
+However, previous research [74] primarily employed simplistic rules and manual refinement processes for text description of satellite images, leading to potential unresolved hallucination issues [62] in generated text descriptions.
+Besides, rule-based methods are overly general and fail to address specific issues in each image description as they are tailored to particular cases.
+On the other hand, manual rewriting is highly labor-intensive.
+Therefore, an effective method for evaluating text quality and calibrating the alignment between text and image content is necessary to ensure a high level of LMM-inherent knowledge injection.
+In contrast to the common evaluation methods used to assess the quality of descriptions generated by LMMs, which typically rely on manually annotated image descriptions as references [65, 83, 34] to measure content overlap, we introduce an automatic mechanism designed to simulate human evaluation systems.
+This mechanism places greater emphasis on semantic similarity rather than merely assessing typical content overlap, thereby introducing PerceptionScore as a novel quality metric.
+Specifically, it comprises two components: text semantic quality and visual recall quality.
+The former approach directly adopts the efficient implementation of CLIPScore [25], which is based on CLIP’s [61] strong zero-shot capabilities and demonstrates high correlation with human semantic judgment.
+However, CLIPScore solely provides a measure of textual quality and fails to indicate whether specific elements in the image are omitted in the text (visual recall) [26].
+Figure 3: The procedure of CycleScore calculation.
+[…]
+$\begin{split}&I^{\prime}=\operatorname{SDXL}(I),\quad\texttt{CycleScore}(I,I^{\prime})=1-\operatorname{MAE}(\operatorname{Seg}(I),\operatorname{Seg}(I^{\prime})),\\ &\texttt{PerceptionScore}(I,T)=(\texttt{CLIPScore}(I,T)+\mathop{\texttt{CycleScore}(I,I^{\prime})})/2.\end{split}$ (1)
+3.2 Multi-Granularity Cross-modal Alignment
+Modality Representation.
+Utilizing automated generation techniques, we obtain a dataset of high-quality image-text pairs, represented as $D=(I,T)$.
+Here, $I$ denotes satellite images $I^{st}$ or street-view images $I^{sv}$, while $T$ represents satellite text descriptions $T^{st}$ or street-view text descriptions $T^{sv}$.
+We first deploy Vision Transformer (ViT) [15] as an image encoder to process images derived from satellite and street-view sources.
+Finally, we use the output $\mathbf{z}_{l}^{0}$ corresponding to the special tokens ${I}_{cls}$ inserted at the beginning as the final output of the image encoder, rewrite as $\mathbf{z}_{I}$ for image representation.
+Concurrently, we use basic Transformer-Encoder [68] as a text encoder by default to comprehend and encode the semantics of textual descriptions in parallel.
+Furthermore, considering that the geospatial location information associated with street-view images can aid in pinpointing the precise geospatial coordinates of the area, we utilize the open-sourced GeoCLIP’s [69] location encoder to effectively instill geospatial information.
+This location encoder captures longitude and latitude $\mathcal{L}_{g}$, generating semantically rich, high-dimensional geospatial features, denoted as $\mathbf{z}_{L}$.
+Modality Alignment.
+The efficient alignment of visual and textual modalities is crucial for accurately modeling regional representations at both the coarse-grained global level and the fine-grained local level.
+To capture global visual information, in addition to the overall visual features for the region derived from satellite images, the clustering aggregation of location-level clues gathered within the region can additionally inject fine-grained details.
+Therefore, we further incorporate fine-grained information from street-view images, which are simultaneously fused with the corresponding geographical features.
+This approach aids in learning a robust and comprehensive representation of urban areas.
+The formalized process is as follows:
+$\mathbf{z}_{g_{st}}=\textit{f}~{}(\mathbf{z}_{I_{st}},~{}\operatorname{Aggr}(\mathbf{z}_{I_{sv}}^{1},\mathbf{z}_{I_{sv}}^{2},...,\mathbf{z}_{I_{sv}}^{m}),~{}\operatorname{Aggr}(\mathbf{z}_{L}^{1},\mathbf{z}_{L}^{2},...,\mathbf{z}_{L}^{m})),$ (2)
+where $\mathbf{z}_{I_{st}}$ represents the satellite visual features for the queried region, $\mathbf{z}_{I_{sv}}^{k}$ denotes the $k$-th street-view visual feature in that region, and $\mathbf{z}_{L}^{k}$ represents the $k$-th street-view location feature in the same region.
+The variable $m$ indicates the count of street-view scenes within the region.
+$\operatorname{Aggr}(\cdot)$ denotes the aggregation method.
+The function f  signifies the feature fusion method.
+To establish a global-level semantic correspondence between satellite visual data and their associated textual descriptions, we employ a joint contrastive optimization approach for the image and text modal encoders in the satellite branch.
+For the $i$-th satellite image-text pair $(I^{st}_{i},T^{st}_{i})$ in a mini-batch, we contrast the image-text pairs against others in the samples, which aims to maximally preserve the mutual information between the authentic pairs in the latent space.
+The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms: $\mathcal{L}_{{CG}}^{Image\rightarrow Text}$ and $\mathcal{L}_{{CG}}^{Text\rightarrow Image}$, which measure the similarity between the visual and textual embeddings, respectively, defined as: The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms which measure the similarity between the visual and textual embeddings, respectively, defined as:
+$\mathcal{L}_{CG}=-\frac{1}{N}(\sum_{i}^{N}{\log\frac{\exp(I_{g_{i}}^{\top}T_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(I_{g_{i}}^{\top}T_{g_{j}}/\tau)}}}+\sum_{i}^{N}{\log\frac{\exp(T_{g_{i}}^{\top}I_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(T_{g_{i}}^{\top}I_{g_{j}}/\tau)}}}),$ (3)
+where $I_{g_{i}}$ and $T_{g_{j}}$ are the normalized embedding of satellite representation $\mathbf{z}_{g_{st}}$ in the $i$-th pair and that of textual representation $\mathbf{z}_{T_{st}}$ in the $j$-th pair, respectively.
+Besides, $N$ is the batch size, and $\tau$ is the temperature for contrastive learning.
+For capturing the local-level information, given that the street-view branch is expected to capture fine-grained details to enrich region embedding at ground level, approaches like [61, 30] rely exclusively on the similarity of each modality’s global feature, thus lacking the capability to capture finer-level information such as the relationship between visual objects and textual tokens [76].
+We leverage a fine-grained interaction mechanism to implement cross-modal alignment.
+Specifically, we utilize token-level maximum similarity between visual and textual tokens to direct the contrastive objective.
+Initially, we compute the similarity between each visual token and all textual tokens, and then leverage the maximum value to calculate the average similarity of all image tokens to textual tokens.
+The similar approach is also applied to text-to-image.
+$\operatorname{SIM}(v_{i},t_{i})=\frac{1}{l_{1}}(\sum_{k_{1}=1}^{l_{1}}{\operatorname{argmax}_{k_{2}\in[0,l_{2})}(v_{ik_{1}}^{\top}t_{ik_{2}})}),\vspace{-0.5em}$ (4)
+$\operatorname{SIM}(t_{i},v_{i})=\frac{1}{l_{2}}(\sum_{k_{2}=1}^{l_{1}}{\operatorname{argmax}_{k_{2}\in[0,l_{1})}(t_{ik_{1}}^{\top}v_{ik_{2}})}),$ (5)
+where $v_{i}$ and $t_{j}$ denote the normalized embedding of street-view representation $\mathbf{z}_{I_{sv}}$ in the $i$-th pair and that of textual representation $\mathbf{z}_{T_{sv}}$ in the $j$-th pair, respectively.
+The fine-grained token-level representation can be optimized through:
+$\mathcal{L}_{CL}=-\frac{1}{N}(\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(v_{i}^{\top},t_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(v_{i}^{\top},t_{j})/\tau)}}}+\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(t_{i}^{\top},v_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(t_{i}^{\top},v_{j})/\tau)}}}).$ (6)
+3.3 Pretraining & Fine-Tuning
+Pretraining Stage. The overall objective of UrbanVLP can be defined as the simultaneous optimization of the aforementioned two losses:
+$\mathcal{L}_{Total}=\alpha\mathcal{L}_{{CG}}+\beta\mathcal{L}_{{CL}}.$ (7)
+where $\alpha$ and $\beta$ are weight hyperparameter.
+Through backpropagation optimization, We achieve multi-granularity cross-modal urban image-text input data alignment, resulting in a robust encoder.
+Fine-Tuning Stage.
+During the fine-tuning stage, we employ a linear probing approach [23].
+This process begins by extracting frozen features $\mathbf{e}_{st},\mathbf{e}_{sv},\mathbf{e}_{p}$ from the pretrained encoder for satellite images, street-view images, and street-view positions (The latter two are optional).
+Subsequently, these features are fused, and a minimalist classifier (MLP) is trained on top to fine-tune the prediction of urban metrics, denoted as $\mathbf{Y}_{i}=MLP(\mathbf{e}_{sa},\mathbf{e}_{sv},\mathbf{e}_{p})$.
+As depicted in Figure 2, these encoders have adapted to capturing high-level abstractions from extensive visual data, enabling the MLP to manipulate static features and study the model’s inference capabilities.
+Note that in downstream tasks, text information is unnecessary as additional text generation time impedes real-time application.
+4 Experiments
+4.1 Experimental Setup
+Dataset & Task Description.
+In this paper, given the current lack of open-sourced datasets in the research community, we introduce a new benchmark dataset named $\mathtt{CityScape}$, which will be released upon paper notification.
+$\mathtt{CityScape}$ is unique in that it comprises a dual-category structure encompassing both satellite and street-view image components, each paired with corresponding high-quality textual descriptions.
+For the satellite image-text pairs, we adhere to [74] to utilize core area data of four cities in China.
+A comprehensive overview of $\mathtt{CityScape}$, along with relevant statistics, is provided in Appendix B.
+Table 1: Urban indicators prediction results in four datasets. The best results are in bold and the second-best results are underlined.
+Methods UrbanVLP StructuralUrban [41] Vision-LSTM [27] UrbanCLIP-SV [74] UrbanCLIP [74] PG-SimCLR [73] ViT [15]
+Metric $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
+Beijing Carbon 0.769 0.477 0.369 0.736 0.518 0.405 0.674 0.519 0.433 0.489 0.713 0.548 0.703 0.541 0.539 0.430 0.797 0.632 0.411 0.810 0.607
+Population 0.714 0.523 0.411 0.691 0.545 0.427 0.640 0.591 0.518 0.435 0.734 0.581 0.655 0.576 0.459 0.476 1.228 0.963 0.442 0.861 0.635
+GDP 0.537 0.684 0.416 0.512 0.694 0.426 0.497 0.717 0.469 0.188 0.910 0.568 0.514 0.694 0.445 0.270 1.679 1.067 0.265 1.073 0.730
+Night Light 0.470 0.668 0.403 0.421 0.696 0.459 0.354 0.747 0.475 0.304 0.769 0.483 0.377 0.741 0.468 0.367 0.728 0.404 0.358 0.733 0.523
+House Price 0.503 0.644 0.482 0.493 0.649 0.485 0.471 0.658 0.495 0.451 0.674 0.515 0.501 0.647 0.486 0.341 0.718 0.555 0.332 0.719 0.569
+POI 0.299 0.723 0.374 0.275 0.725 0.380 0.235 0.728 0.388 0.226 0.729 0.407 0.251 0.734 0.385 - - - 0.215 0.742 0.411
+Shanghai Carbon 0.718 0.520 0.381 0.678 0.550 0.432 0.571 0.606 0.455 0.576 0.465 0.450 0.673 0.560 0.426 0.270 0.742 0.551 0.254 0.758 0.522
+Population 0.589 0.609 0.476 0.545 0.631 0.488 0.518 0.659 0.522 0.400 0.793 0.592 0.533 0.650 0.511 0.288 1.051 0.831 0.279 0.826 0.627
+GDP 0.323 0.805 0.593 0.312 0.840 0.613 0.311 1.008 0.661 0.323 1.214 0.741 0.334 0.804 0.586 0.270 1.679 1.067 0.263 1.221 0.735
+Night Light 0.435 0.683 0.490 0.411 0.690 0.497 0.352 0.734 0.550 0.274 0.765 0.556 0.384 0.710 0.515 0.228 0.756 0.495 0.209 0.768 0.561
+POI 0.381 0.680 0.386 0.375 0.685 0.392 0.246 0.704 0.390 0.240 0.711 0.402 0.265 0.710 0.395 - - - 0.234 0.713 0.406
+Guangzhou Carbon 0.701 0.513 0.372 0.655 0.550 0.504 0.554 0.564 0.457 0.425 0.520 0.473 0.585 0.603 0.440 0.413 0.951 0.716 0.374 0.665 0.528
+Population 0.655 0.573 0.454 0.631 0.587 0.459 0.614 0.619 0.509 0.491 0.693 0.622 0.627 0.596 0.473 0.294 1.046 2.112 0.263 0.821 0.661
+GDP 0.442 0.755 0.541 0.429 0.781 0.549 0.412 0.997 0.624 0.327 0.998 0.690 0.446 0.752 0.546 0.263 1.157 1.553 0.249 1.122 0.719
+Night Light 0.548 0.594 0.428 0.508 0.607 0.448 0.381 1.021 0.629 0.314 0.692 0.524 0.463 0.651 0.483 0.436 0.649 0.432 0.357 0.642 0.505
+POI 0.438 0.732 0.333 0.357 0.758 0.359 0.175 0.811 0.401 0.169 0.826 0.405 0.185 0.802 0.401 - - - 0.136 0.904 0.427
+Shenzhen Carbon 0.625 0.605 0.464 0.587 0.593 0.481 0.534 0.609 0.516 0.460 0.621 0.529 0.541 0.607 0.526 0.234 0.975 0.746 0.221 0.727 0.574
+Population 0.790 0.452 0.348 0.724 0.476 0.386 0.624 0.562 0.471 0.589 0.604 0.484 0.727 0.510 0.392 0.294 1.046 2.112 0.280 0.832 0.654
+GDP 0.533 0.682 0.447 0.489 0.697 0.471 0.462 0.723 0.517 0.433 0.755 0.577 0.508 0.693 0.464 0.294 1.046 2.112 0.254 1.226 0.779
+Night Light 0.457 0.667 0.459 0.421 0.694 0.502 0.404 0.702 0.508 0.320 0.717 0.531 0.387 0.709 0.511 0.247 0.934 0.738 0.254 0.715 0.530
+POI 0.461 0.752 0.370 0.321 0.780 0.389 0.254 0.795 0.412 0.147 0.849 0.453 0.185 0.838 0.436 - - - 0.137 0.954 0.455
+$1^{\text{st}}$ount 56 1 0 1 5 0 0
+Baselines for Comparison.
+We compare our method with seven recent baselines in the field of urban imagery-based socioeconomic prediction.
+The single granularity imagery-based urban region representation learning methods include: ViT [15], PG-SimCLR [73], UrbanCLIP [74], UrbanCLIP-SV [74].
+We also apply multiple baselines for multi-granularity urban imagery-based representation learning: Vision-LSTM [27] and StructuralUrban  [41].
+The in-depth overview of each baseline method is presented in the Appendix C.
+Evaluation Metrics.
+The performance of our predictions is evaluated using a trio of standard metrics: the coefficient of determination denoted as $R^{2}$, the root mean squared error (RMSE), and the mean absolute error (MAE).
+An increase in $R^{2}$, along with a decrease in RMSE and MAE values, signifies improved model accuracy.
+More implementation details are provided in Appendix D.
+4.2 Performance Evaluation
+Figure 4: $R^{2}$ results in Beijing.
+To further validate the effectiveness of the proposed UrbanVLP framework, we conduct comparisons with existing state-of-the-art methods on our proposed datasets.
+The comprehensive results are presented in Table 1, and we can draw the following conclusions: UrbanVLP significantly surpasses the performance of baseline methods employing both single-granularity and multi-granularity imagery, achieving a substantial margin of improvement.
+It can be seen that UrhanVLP outperforms the best baseline, StructuralUrban [41], by 3.3%, 2.3%, 2.5%, 4.9%, 1.0% and 2.4% in terms of $R^{2}$ for all 6 indicators: Carbon, Population, GDP, Night Light, House Price and POI in Beijing.
+The similar performance improvement can also be seen in the dataset of the other three cities.
+Besides, the average performance gain of UrbanVLP on RMSE and MAE in Beijing dataset are 1.8% and 2.1%, respectively.
+In order to provide a more intuitive display, we have created Radar Images in Figure 4 to visualize the performance of Beijing.
+The results further prove the effectiveness of our framework for urban region representation learning.
+Multi-granularity cross-modal alignment effectively incorporates fine-grained information, demonstrating superior performance.
+We can observe that pipelines leveraging multi-granularity approaches typically demonstrate superior performance compared to those employing a single granularity level.
+This enhancement can be rationalized by the incorporation of additional fine-grained information derived from street-view imagery.
+Concurrently, it is noticed that UrbanCLIP-SV underperforms compared to UrbanCLIP, demonstrating that street-view imagery, while rich in fine-grained details, lacks macroscopic perspective, leading to suboptimal outcomes when used in isolation.
+This also underscores the irreplaceability of satellite imagery as a macroscopic visual modality.
+Compared to other modalities, the textual modality facilitates a more explicit and comprehensive understanding of the region.
+In baseline models, there are works such as Vision-LSTM [27], which introduces trajectories, and PGSimCLR [73], which incorporates POI distributions to integrate external information and enhance regional representations.
+However, the superior performance of our UrbanVLP illustrates the efficacy of textual mobility as a high-information-density modality [23], which fully leverages the inherent knowledge of LLMs and explicitly enhances interpretability.
+4.3 Ablation Studies
+Figure 5: Ablation study on $\mathtt{CityScape}$-$\mathtt{Beijing}$.
+We conduct ablation studies to investigate the effectiveness of different components in UrbanVLP on Beijing dataset, including the Satellite branch (ST branch), Street-View branch (SV branch), and Location Encoding branch (LE branch).
+More ablation experiments can be found in Appendix H.
+The effectiveness of incorporating street-view branch.
+One of the main contributions of this work is the incorporation of multi-granularity information that reflects the urban spatial hierarchy.
+The street-view branch specifically represents micro-level modeling.
+It can be observed that the incorporation of street-view branch results in an average improvement of 3.56% in terms of $R^{2}$.
+The presence of the street-view branch introduces the capture of rich detailed information, thereby enhancing the precision of the modeling.
+Furthermore, it is also can be seen that while street-view imagery provides fine-grained clues, its individual impact is still overshadowed by that of satellite imagery.
+This aligns with the performance differential between UrbanCLIP and UrbanCLIP-SV discussed in the Section 4.2, underscoring the indispensable role of macro-level modeling
+The effectiveness of textual modality.
+In Table 1, we demonstrate the function of the textual modality by comparing UrbanVLP with a standard ViT-based model, which shares the same configuration as the unimodal visual encoder of UrbanVLP (encompassing both satellite and street-view branches).
+We then utilize the visual features extracted without textual augmentation to predict downstream urban indicators.
+As we can see, the lack of textual information leads to substantial performance degradation, underscoring the critical role of textual modalities in attaining a comprehensive visual representation.
+Similar findings were also reported in [74].
+Exploration of the impact of location coordinates.
+The role of location coordinates has been underexplored in previous work of urban region representation learning.
+However, its importance has already been emphasized in work related to Geo-Localization [69, 79, 3].
+In Figure 5, we evaluate the performance of the model without the Location Encoding branch.
+Although not as impactful as visual modalities, the inclusion of geospatial locations has yielded an average improvement of 1.5% in $R^{2}$.
+However, when relying solely on location information, the performance is suboptimal.
+This indicates that at the current stage, location data can only serve as an auxiliary enhancement and lacks the capacity for standalone modeling due to its deficiency in contextual semantic information.
+Figure 6: Illustration of the quality of our generated description example.
+Figure 7: Examples of predictions from UrbanCLIP and our proposed UrbanVLP.
+4.4 Qualitative Analysis
+We further investigate the quality of generated texts and the predictive performance of UrbanVLP in practice. Due to page limitations, more experimental analysis can be found in the Appendix.
+Illustration of the quality of our generated descriptions.
+To visually illustrate the quality of the generated description, we present an example in Figure 7.
+The generated description of the street-view image (a) is depicted in (b).
+Subsequently, we employ GPT4V [75] to assess the quality score of the generated description, which is rated as 7 out of 10, indicating the effectiveness of the generated text.
+Furthermore, GPT4V also provides specific areas for improvement in (c).
+In (d), we utilize GPT4V to generate an image based on the description in (b).
+It is evident that the generated image bears a resemblance to the original one.
+Figure 8: Our web platform. Zoom in or see Figure 17 for a better view.
+Case Study for Predicted Results.
+Here we also show some predicted results in Figure 7 on $\mathtt{CityScape}$-$\mathtt{Beijing}$ dataset.
+As we can see, satellite images (a) and (b), although sharing similar layouts, differ significantly in their land use.
+(a) encompasses residential and campus areas, whereas (b) is indicative of industrial zones.
+Their socio-economic characteristics are entirely different, which could be reflected by socio-enconomic indicators.
+The carbon emissions and GDP of campus and residential area are significantly lower than those of industrial park, whereas the disparities in population and housing prices are not as pronounced, possibly due to the unique characteristics of school district housing near educational institutions.
+Night light intensity is similar across the two area due to the fact that the majority of nocturnal economic activities within factories occur indoors, unlike in Central Business Districts (CBDs) where extensive illumination is generated.
+The same principle applies to school and residential area as well.
+As observed, the predictions made by UrbanCLIP do not effectively differentiate between the two, with the downstream metrics in Figure 7(b) still inclined towards predicting a homogenized socio-economic functional area similar to that of (a).
+In contrast, the results from UrbanVLP are capable of distinctly distinguishing the socio-economic attributes of the two in terms of carbon emission, GDP, and house price.
+To showcase practical applications, we further develop a web-based system (Figure 8 and 17), with more details and discussion in Appendix I.
+Visualization of Region Representations.
+In this section, we map the region representations learned in UrbanVLP into a two-dimensional space using the PCA algorithm in Figure 10.
+This process aims to visualize how the semantic meanings of the regions are manifested within the high-dimensional representation space.
+As we can see, the three images on the left and on the right belong to different clusters, each exhibiting intra-cluster similarities and inter-cluster differences.
+Specifically, UrbanVLP effectively models regions into high-dimensional space, wherein satellite images with similar architectural layouts demonstrate spatial similarity.
+Consequently, this leads to a reasonable, accurate, and semantically rich representation of the regions.
+Figure 9: Representation space visualization.
+Figure 10: PCA visualization of urban imagery.
+Figure 10 provides a further visualization with a population indicator, compared to UrbanCLIP.
+We use the dot color to indicate the population (log) in the corresponding regions of Beijing.
+As shown by the red dashed box, the regions modeled by UrbanVLP exhibit more significant clustering characteristics and greater distinction in high-dimensional space.
+Additionally, the lower parts of the two main band-shaped branches in the figure are also more distinctly separated.
+Points that are spatially close also have similar population values.
+5 Conclusion and Future Work
+Urban region representation learning plays a significant role in helping us understand the patterns, structures, and dynamics of cities.
+UrbanVLP, for the first time, explores the differences between street-view and satellite images from a semantic granularity perspective, as well as their roles in modeling urban region representation.
+A text generation and calibration mechanism is also proposed to ensure high-quality description generation.
+It has achieved state-of-the-art results on the constructed $\mathtt{CityScape}$ and downstream dataset.
+In the future, the further research direction can be two-fold: 1) Adding additional modalities such as POIs to enhance information richness, and 2) enabling the design of better model architectures such as MAE [23] to efficiently utilize existing data source.
+Appendix / supplemental material
+Appendix A Formulation
+Definition 1 (Urban Region) We follow previous research [73, 74] and evenly divide the target area (e.g. a city) into $L$ urban regions.
 Definition 2 (Urban Imagery) Diverse visual images, including satellite and street-view images, facilitate an intuitive analysis of urban regions.
@@
 In this paper, we comprehensively use population (count), GDP (in million Chinese Yuan), carbon emissions (in tons), night lights (intensity), house prices (in Chinese Yuan), and points of interest, i.ePOI, (count for different categories) as ground-truth urban indicators.
-Problem Statement (Urban Region Representation) We next formally define the problem of urban region representation learning.
-Given a satellite images $I^{sa}_{g}$, a set of street-view images $\mathcal{I}^{sv}_{g}$ belonging to the target coverage, and their corresponding latitude and longitude pair $\mathcal{L}_{g}$.
-The objective is to employ a learning function $\mathcal{F}$ to map them to representation vectors $\mathbf{e}_{g}=\mathcal{F}(I^{sa}_{g},\mathcal{I}^{sv}_{g},\mathcal{L}_{g})$.
-The representation can then be further utilized to infer urban indicators $\mathbf{Y}_{g}$ for any given region $g$.
-We assume that during the training phase, automated additional text descriptions are available.
-2.2. Related Work
-2.2.1. Urban Region Representation Learning
-In literature, many studies have learned task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility (Li et al., 2022c; Liu et al., 2023d; Xi et al., 2022).
-For example, Jean et al.
-(2016b) used nighttime light images as weak labels to learn area representation and assess regional economic development, while Urban2Vec (Wang et al., 2020) integrated street-view imagery and POI data to learn neighborhood embeddings.
-Some contrastive learning approaches like PG-SimCLR (Xi et al., 2022) and UrbanCLIP (Yan et al., 2023) have shown success in representing urban regions through satellite image.
-However, the aforementioned works exclusively considered a single type of urban imagery and overlooked the potential synergy between satellite imagery and street-view imagery, which can complement each other.
-Recently, some efforts (Barbierato et al., 2020; Huang et al., 2023a; Li et al., 2022c; Cao et al., 2023, 2018a; Law et al., 2019; Cao et al., 2018b; Huang et al., 2023b) have begun to explore the simple combination of satellite imagery and street-view imagery.
-For instance, Vision-LSTM (Huang et al., 2023b) integrates diverse street-level images into the satellite image domain using an LSTM-based network for precise urban village identification.
-Li et al.
-(2022c) propose utilizing street segments for region modeling and predict socio-economic indicators by merging street-view and satellite images through a multi-stage pipeline.
-Nevertheless, prior studies overlook distinct features in satellite and street-view images for varied granularities.
-Additionally, few studies incorporate text as a supervisory signal.
-2.2.2. Large Multimodal Model
-As LLMs undergo rapid evolution, a faction within the research community is increasingly focused on integrating visual knowledge into these models, such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021).
-Early studies (Li et al., 2022a, 2023c) enhanced CLIP with refined data strategies for more diverse datasets, proving effective for fundamental visual tasks (Li et al., 2022d; Liu et al., 2023c; Zhang et al., 2022) but less so for complex tasks like visual question answering.
-Recently, new models like MiniGPT-4 (Chen et al., 2023b), LLaVA (Liu et al., 2023b), and InstructBLIP (Dai et al., 2023) have improved the understanding of complex problems by expanding parameters and training data, called large multimodal models (LMMs).
-In this paper, we comprehensively evaluate the text generation quality of representative LMMs model and propose an automated text description generation and calibration technique to ensure effective image description capabilities with LMMs.
-2.2.3. Vision-Language Pre-training (VLP)
-VLP aims to jointly encode vision and language in a fusion model.
-Early works on VLP can be broadly categorized into single-stream (Chen et al., 2020; Kim et al., 2021; Li et al., 2020) and two-stream (Tan and Bansal, 2019; Lu et al., 2019) methodologies, with the former using a unified architecture for modality integration and the latter employing separate encoders before merging modalities.
-Milestone work CLIP (Radford et al., 2021) and its variants (Li et al., 2021; Yao et al., 2022a) highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval.
-Recent works (Tsimpoukelli et al., 2021; Alayrac et al., 2022) shift towards leveraging LLMs knowledge for vision-language representation learning.
-In urban indicator prediction, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information.
-Our groundbreaking work introduces multiple types of text descriptions as supplements, aiming to learn more comprehensive and interpretable urban region representations.
-3. Methodology
-We present an overview of our proposed framework in Figure 3. There are two principal stages in our framework:
-  • Stage 1: We first devise an automatic text generation and calibration module.
-Leveraging ShareGPT4V (Chen et al., 2023a) (current state-of-the-art LMM captioner model), textual descriptions are created for each street-view image, incorporating geographical and visual prompts.
-PerceptionScore is also introduced to ensure high-quality text retention from both semantic and visual perspectives.
-Building upon the foundation of high-quality image-text pairs, we further develop a multi-granularity cross-modal alignment framework, achieving multi-granular alignment and multimodal correlation through multi-level contrastive learning and fine-grained information injection.
-  • Stage 2: In the urban indicator prediction stage, we leverage the frozen encoder pre-trained in Stage 1 to extract input data features.
-Subsequently, a lightweight multi-layer perceptron (MLP) is fine-tuned to yield precise urban indicator predictions.
-Additionally, optional information such as street-view imagery and geographical coordinates is considered.
-3.1. Automatic Text Generation and Calibration
-Text Generation.
-For each satellite image, we initiate a range query to obtain the corresponding set of available street-view images and their respective geospatial coordinates within the designated area.
-Existing works (Lai et al., 2023; Fan et al., 2023a; Yan et al., 2023) proved that the quality of the text is crucial for enhancing performance.
-Therefore, advanced LMMs, owing to their robust cross-modal knowledge storage capabilities, offer detailed textual comprehension and descriptions for each street-view image.
-However, this confronts two primary challenges.
-Firstly, due to substantial financial implications associated with API usage, employing closed source LMMs like GPT4V (Yang et al., 2023) or Gemini (Fu et al., 2023) for generating tens of thousands of text descriptions is impractical.
-Secondly, street-view images often encompass diverse detailed information, necessitating a well-designed prompt template for obtaining high-quality textual descriptions.
-To address these issues, we first utilize a closely related GPT4V model, known as ShareGPT4V (Chen et al., 2023a), a powerful language model with vision capabilities trained on 100k GPT4V generated captions, to generate street-view image descriptions.
-In the process of designing text prompt templates, the proportion of segmented elements in street-view images reflects the distribution of elements, aiding LMMs in evaluating the importance of each element.
-To achieve this, we employ a pre-trained segmentation model to decouple visual elements and calculate the proportion of segmentation for each element.
-Simultaneously, urban and geospatial coordinates of street-view images are included as prompts to guide the expression of additional urban information.
-Finally, the prompt template is formulated as follows: Analyze the street-view panoramic image in [City] in a comprehensive and detailed manner.
-The coordinate of the street-view image is [Longitude, Latitude].
-The segmentation ratio of the street-view image is [Segment Information] .
-More prompt examples and segmentation ratio details can be found in Appendix C.
-Text Calibration.
-In contrast to previous research (Yan et al., 2023) that employed simplistic rules and manual refinement processes for text description of satellite images, street view images often contain rich fine-grained information, leading to potential hallucination issues (Rawte et al., 2023) in generated text descriptions that require further calibration.
-Rule-based methods are overly general and fail to address specific issues in each image description as they are tailored to particular cases.
-On the other hand, manual rewriting is labor-intensive and demands significant effort.
-Therefore, an effective method for evaluating text quality is crucial.
-[…]
-$\begin{split}&I^{\prime}=\operatorname{SDXL}(I),\\ &\texttt{CycleScore}(I,I^{\prime})=1-\operatorname{MAE}(\operatorname{Seg}(I),\operatorname{Seg}(I^{\prime})),\\ &\texttt{PerceptionScore}(I,T)=(\texttt{CLIPScore}(I,T)+\mathop{\texttt{CycleScore}(I,I^{\prime})})/2.\end{split}$
-Figure 4. The procedure of Cycle Score calculation.
-3.2. Multi-Granularity Cross-modal Alignment
-Modality Representation.
-Utilizing automated generation techniques, we obtain a dataset of high-quality image-text pairs, represented as $D=(I,T)$.
-Here, $I$ denotes either a satellite image $I^{sa}_{g}$ or a street view image $I^{sv}_{g}$, while $T$ represents either a satellite text description $T^{sa}_{g}$ or a street view text description $T^{st}_{g}$.
-we first deploy Vision Transformer (ViT) (Dosovitskiy et al., 2021) as an image encoder to process images derived from satellites and street-views, proficient in deconstructing images into patches for feature extraction.
-In particular, we reshape the input image $I\in\mathbb{R}^{H\times W\times 3}$ into a sequence of flattened 2D patches ${I}_{P}\in\mathbb{R}^{N\times({P^{2}}\cdot{C})}$, where C is channel number, $(P,P)$ is the resolution of each image patch, and $N=HW/P^{2}$ is the patch number, which is the input sequence length of ViT.
-Notebly, the ViT utilizes a consistent latent vector size (denoted as $d$) across all its layers.
-Thus, we can obtain the outputs as the patch embeddings through a trainable linear projection.
-We also prepend a learnable embedding to the patch embedding sequence (i.e., $\mathbf{z}_{0}^{0}={I}_{cls}$).
-Besides, we add learnable position embedding $\mathbf{E}_{pos}$ to retain positional information.
-The process is as follows:
-(1) $\mathbf{z}_{0}=[{I}_{cls};{I}_{p}^{1}\mathbf{W};{I}_{p}^{2}\mathbf{W};\cdots;{I}_{p}^{N}\mathbf{W}]+\mathbf{E}_{pos},$
-where $\mathbf{W}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$ and $\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times d}$ are learnable parameter.
-$\mathbf{z}_{0}$ is the linear projection result and then sent to ViT.
-Consistent with the standard Transformer encoder, ViT includes alternating layers of multi-head self-attention $\operatorname{MSA(\cdot)}$ operation (Vaswani et al., 2017).
-Furthermore, layer normalization (LN) is adopted before every block, whereas residual connections are after every block:
-(2) $\mathbf{z}_{l}=\mathrm{MSA}(\mathrm{LN}(\mathbf{z}_{l-1}))+\mathbf{z}_{l-1},\quad l=1\ldots L.$
-Finally, we use the output $\mathbf{z}_{l}^{0}$ corresponding to the special tokens ${I}_{cls}$ inserted at the beginning as the final output of the image encoder, rewrite as $\mathbf{z}_{I}$ for image representation.
-Concurrently, we use basic Transformer-Encoder (Vaswani et al., 2017) as a text encoder by default to comprehend and encode the semantics of textual descriptions in parallel.
-The text encoder follows a similar $\operatorname{MSA}$ mechanism of the visual encoder.
-The input text sequence is bracketed with $[SOS]$ and $[EOS]$ tokens, and the activation of the highest layer of Transformer at $[EOS]$ token is considered the global representation $\mathbf{z}_{T}$ of text.
-Moreover, depending on whether it is a satellite image-text pair or a street-view image-text pair, we categorize them into two branches with separate visual and text encoders – namely, the satellite branch and the street-view branch.
-Furthermore, considering the longitude and latitude information corresponding to street view images, which aids in identifying the geospatial location of the area, we leverage the open-source GeoCLIP’s (Vivanco et al., 2023) location encoder.
-This encoder captures longitude and latitude information at different resolutions, generating semantically rich high-dimensional features.
-Here, a set of longitude and latitude $\mathcal{L}_{g}$ is aggregated into a single vector, denoted as $\mathbf{z}_{L}$, thereby incorporating geographical positioning.
-Modality Alignment. Our goal is to capture the coarse-grained global regional information from the satellite branch and the fine-grained local positional information from the street-view branch.
-For capturing global information, even though we have obtained overall visual features for the region from images in the satellite branch, we can further incorporate fine-grained information from images in the street-view branch.
-Specifically, the visual features of streets within each satellite region are aggregated, simultaneously fused with the corresponding geographical feature, aiding in learning a robust and comprehensive representation of urban areas.
-The formalized process is as follows:
-(3) $\mathbf{z}_{g_{sa}}=\textit{f}~{}(\mathbf{z}_{I_{sa}},~{}\operatorname{Aggr}(\mathbf{z}_{I_{sv}}^{1},\mathbf{z}_{I_{sv}}^{2},...,\mathbf{z}_{I_{sv}}^{m}),~{}\operatorname{Aggr}(\mathbf{z}_{L}^{1},\mathbf{z}_{L}^{2},...,\mathbf{z}_{L}^{m})),$
-where $\mathbf{z}_{I_{sa}}$ represents the satellite visual features for the queried region, $\mathbf{z}_{I_{sv}}^{k}$ denotes the $k$-th street-view visual feature in that region, and $\mathbf{z}_{L}^{k}$ represents the $k$-th street view location feature in the same region.
-The variable $m$ represents the count of street views and their corresponding locations within the region.
-$\operatorname{Aggr}(\cdot)$ denotes the aggregation method for all street views in the region, utilizing additive aggregation in this case.
-The function f  signifies the feature fusion method.
-To establish a global-level semantic correspondence between satellite images and their related descriptions, we employ a joint optimization approach for the image and text modal encoders in the satellite branch.
-For the $i$-th satellite image-report pair $(I_{g_{i}},T_{g_{i}})$ in a mini-batch, we regard two modality data as queries and keys alternatively to learn the correct image-report pairings, to maximally preserve the mutual information between the true pairs in latent space.
-The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms: $\mathcal{L}_{{CG}}^{Image\rightarrow Text}$ and $\mathcal{L}_{{CG}}^{Text\rightarrow Image}$, which measure the similarity between the image and text embeddings, respectively, defined as:
-(4) $\mathcal{L}_{{CG}}^{Image\rightarrow Text}=-\sum_{i}^{N}{\log\frac{\exp(I_{g_{i}}^{\top}T_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(I_{g_{i}}^{\top}T_{g_{j}}/\tau)}}},$
-(5) $\mathcal{L}_{{CG}}^{Text\rightarrow Image}=-\sum_{i}^{N}{\log\frac{\exp(T_{g_{i}}^{\top}I_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(T_{g_{i}}^{\top}I_{g_{j}}/\tau)}}},$
-(6) $\mathcal{L}_{CG}=\frac{1}{N}(\mathcal{L}_{{CG}}^{Image\rightarrow Text}+\mathcal{L}_{{CG}}^{Text\rightarrow Image}),$
-where $I_{g_{i}}$ and $T_{g_{j}}$ are the normalized embedding of satellite image representation $\mathbf{z}_{g_{sa}}$ in the $i$-th pair and that of text representation $\mathbf{z}_{T_{sa}}$ in the $j$-th pair, respectively.
-Besides, $N$ is the batch size, and $\tau$ is the temperature for contrastive learning.
-For capturing local information, given that the street-view local branch is expected to capture fine-grained information to provide ground-level details for region embedding, approaches like (Radford et al., 2021; Jia et al., 2021) rely exclusively on the similarity of each modality’s global feature, thus lacking the capability to capture finer-level information such as the relationship between visual objects and textual tokens (Yao et al., 2022a).
-We utilize a fine-grained interaction mechanism to implement alignment.
-Specifically, we utilize token-level maximum similarity between visual and textual tokens to direct the contrastive objective.
-Initially, we compute the similarity between each visual token and all textual tokens, and then leverage the maximum value to calculate the average similarity of all image tokens to textual tokens.
-The similar approach is also applied to text-to-image.
-(7) $\operatorname{SIM}(v_{i},t_{i})=\frac{1}{l_{1}}(\sum_{k_{1}=1}^{l_{1}}v_{ik_{1}}^{\top}{\operatorname{argmax}_{k_{2}\in[0,l_{2})}(t_{ik_{2}})}),$
-(8) $\operatorname{SIM}(t_{i},v_{i})=\frac{1}{l_{2}}(\sum_{k_{2}=1}^{l_{1}}t_{ik_{1}}^{\top}{\operatorname{argmax}_{k_{2}\in[0,l_{1})}(v_{ik_{2}})}),$
-where $v_{i}$ denotes the $i$-th token of visual features and $t_{j}$ represents the $j$-th token of textual features.
-$l_{1}$ and $l_{2}$ denote the token length of visual and textual features in street-view branch respectively.
-The fine-grained token-level representation can be optimized through:
-(9) $\mathcal{L}_{{CL}}^{Image\rightarrow Text}=-\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(v_{i}^{\top},t_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(v_{i}^{\top},t_{j})/\tau)}}},$
-(10) $\mathcal{L}_{{CL}}^{Text\rightarrow Image}=-\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(t_{i}^{\top},v_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(t_{i}^{\top},v_{j})/\tau)}}},$
-(11) $\mathcal{L}_{CL}=\frac{1}{N}(\mathcal{L}_{{CL}}^{Image\rightarrow Text}+\mathcal{L}_{{CL}}^{Text\rightarrow Image}).$
-3.3. Urban Indicator Prediction
-Pre-training Stage. The total objective of UrbanVLP can be defined as the sum of the above two losses:
-(12) $\mathcal{L}_{Total}=\alpha\mathcal{L}_{{CG}}+\beta\mathcal{L}_{{CL}}.$
-where $\alpha$ and $\beta$ are weight hyperparameter.
-Through backpropagation optimization, We achieve multi-granularity cross-modal urban image-text input data alignment, resulting in a robust encoder.
-Prediction Stage.
-During the fine-tuning stage, we employ a linear probing approach (He et al., 2022).
-This process begins by extracting frozen features $\mathbf{e}_{sa},\mathbf{e}_{sv},\mathbf{e}_{t}$ from the pre-trained encoder for satellite images, street view images, and street view positions (The latter two are optional).
-Subsequently, these features are fused, and a minimalist classifier (MLP) is trained on top to fine-tune the prediction of urban metrics, denoted as $\mathbf{Y}_{i}=MLP(\mathbf{e}_{sa},\mathbf{e}_{sv},\mathbf{e}_{t})$.
-As depicted in Figure 3, these encoders have adapted to capturing high-level abstractions from extensive visual data, enabling the MLP to manipulate static features and study the model’s inference capabilities.
-4. Experiments
-In this section, we conduct extensive experiments to investigate the following Research Questions (RQ):
-  • RQ1: Can UrbanVLP outperform previous baseline methods and generalize well to various urban indicator tasks?
-  • RQ2: How does each component (e.g., satellite branch, street-view branch, location encoding branch) contribute to UrbanVLP?
-  • RQ3: How do automated text generation and calibration affect the generated text quality?
-  • RQ4: How do we envision the practicality of UrbanVLP?
-Given the limits of page number, RQ4 is detailed in Appendix D.
-4.1. Datasets & Tasks
-Our dataset comprises a dual-category structure encompassing both satellite and street-view image components, each paired with corresponding textual descriptions.
-For the satellite image-text pairs, we follow and refine the setting of (Yan et al., 2023) to utilize core area data of 4 cities in China.
-4.1.1. UrbanSV Dataset
-We construct the UrbanSV Dataset by street-view geo-tagged images collected from Baidu Map API and generated text descriptions by LMM.
+Appendix B More Details about Datasets & Tasks
+B.1 CityScape Dataset
+We construct the $\mathtt{CityScape}$ Dataset using street-view geo-tagged images collected from Baidu Map API and generated text descriptions by LMM.
   • Street-View Panoramic Image Collection.
@@
 Following the road network sampling, panoramic street-view images were collected based on the sampling points.
-These images, with a resolution of 2048x664, also far exceed the 256x256 resolution of satellite imagery, ensuring clarity and ample detail for further analysis.
-As is shown in Table 1, a total of 47,582 street-level images were collected, whose data volume is 2.7 times that of the satellite dataset, capturing the intricacies of the urban landscape and providing a street-level view that enhances the overall urban analysis.
+These images, with a resolution of 2048x664, also far exceed the 256x256 resolution of satellite imagery [74], ensuring clarity and ample details for further analysis.
+    As is shown in Table 2, a total of 47,582 street-level images were collected, whose data volume is 2.7 times larger than that of the satellite dataset, capturing the intricacies of the urban landscape and providing a street-level view that enhances the overall urban analysis.
+    In figure 11, we showcase our sampling process for the road network, which is used to collect street-view imagery at strategically selected locations.
+We initially acquire road network data from the online open-source website and subsequently import it into ArcGIS with a sampling interval of 500 meters.
+Following this, we gather the latitude and longitude coordinates of the sampling points within our dataset area.
+These coordinates are then used to invoke the Baidu API to collect street-view imagery.
+Additionally, not all sampled points have corresponding street view images, thereby resulting in a final count of street view images that is less than the number of sampled points.
   • Corresponding Text Description Generation.
-The textual data corresponding to each satellite image is generated from ShareGPT4V (Chen et al., 2023a), which is recognized for its superior text generation capabilities compared to other contemporary open-source LMM models (Liu et al., 2023a; Dai et al., 2023; Ye et al., 2023; Gao et al., 2023; Chen et al., 2023b, a), offering comprehensive and high-quality textual information.
-The generated quality can be guaranteed by our perception scores.
-We filter out the texts with PerceptionScores less than 0.5.
-Table 1. Urban imagery dataset statistics.
+The textual data corresponding to each street-view image is generated from ShareGPT4V [11], which is recognized for its superior text generation capabilities compared to other contemporary open-source LMM models [48, 14, 77, 20, 10, 11], offering comprehensive and high-quality textual information.
+The quality of the generated data is guaranteed by our proposed PerceptionScore.
+We filter out the texts with PerceptionScore lower than 0.5.
+Table 2: Urban imagery dataset statistics.
 Dataset Coverage #Street -view Image #Satellite Image
@@
 Shenzhen 22.45°N, 113.75°E 22.84°N, 114.62°E 10,167 4,324
-4.1.2. Downstream Task Dataset
-In our research, the Downstream Task Dataset plays a crucial role in gauging the practical implications of our multimodal framework.
-We collect six representative indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI.
+Figure 11: Road Network Sampling for Street-View Images on Beijing.
+B.2 Downstream Dataset & Tasks
+In our research, the Downstream Dataset plays a crucial role in gauging the practical implications of our multimodal framework.
+We collect six representative urban indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI.
 Each indicator illuminates a significant dimension of urban vitality and sustainability.
-The detailed introduction of each downstream indicator is provided in the Appendix A.
-Table 2. Urban indicators prediction results in four datasets. The best results are in bold and the second-best results are underlined.
-Methods UrbanVLP StructualUrban (Li et al., 2022c) Vision-LSTM (Huang et al., 2023a) UrbanCLIP-SV (Yan et al., 2023) UrbanCLIP (Yan et al., 2023) PG-SimCLR (Xi et al., 2022) ResNet-18 (He et al., 2016) Autoencoder (Kramer, 1991)
-Metric $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
-Beijing Carbon 0.787 0.457 0.353 0.745 0.500 0.397 0.692 0.502 0.421 0.504 0.697 0.536 0.716 0.528 0.528 0.430 0.797 0.632 0.383 0.821 0.598 0.294 0.851 0.572
-Population 0.725 0.513 0.404 0.699 0.536 0.422 0.649 0.583 0.511 0.447 0.727 0.574 0.661 0.569 0.452 0.476 1.228 0.963 0.251 0.874 0.657 0.164 0.918 0.673
-GDP 0.586 0.650 0.416 0.560 0.670 0.418 0.536 0.690 0.454 0.208 0.899 0.553 0.555 0.674 0.439 0.270 1.679 1.067 0.242 1.253 0.733 0.165 1.521 0.838
-Night Light 0.531 0.629 0.394 0.471 0.668 0.447 0.409 0.712 0.466 0.352 0.733 0.476 0.420 0.700 0.457 0.367 0.728 0.404 0.337 0.740 0.581 0.276 0.784 0.643
-House Price 0.506 0.640 0.479 0.494 0.647 0.481 0.470 0.655 0.493 0.453 0.672 0.512 0.503 0.641 0.482 0.341 0.718 0.555 0.284 0.721 0.576 0.332 0.714 0.576
-POI 0.319 0.706 0.383 0.277 0.712 0.385 0.236 0.718 0.388 0.227 0.724 0.410 0.253 0.731 0.388 - - - 0.209 0.747 0.415 0.164 0.759 0.427
-Shanghai Carbon 0.716 0.529 0.392 0.677 0.556 0.431 0.568 0.613 0.458 0.574 0.473 0.453 0.671 0.569 0.426 0.270 0.742 0.551 0.243 0.763 0.528 0.230 0.771 0.532
-Population 0.593 0.607 0.471 0.548 0.630 0.486 0.521 0.657 0.518 0.402 0.793 0.590 0.536 0.648 0.507 0.288 1.051 0.831 0.274 0.833 0.632 0.166 0.898 0.742
-GDP 0.310 0.816 0.595 0.305 0.841 0.613 0.304 1.012 0.664 0.315 1.214 0.741 0.326 0.807 0.587 0.270 1.679 1.067 0.241 1.237 0.740 0.197 1.511 0.844
-Night Light 0.457 0.667 0.459 0.412 0.689 0.494 0.355 0.731 0.546 0.276 0.763 0.554 0.387 0.709 0.511 0.228 0.596 0.466 0.202 0.789 0.573 0.164 0.810 0.617
-POI 0.413 0.673 0.363 0.387 0.681 0.375 0.247 0.699 0.389 0.243 0.707 0.399 0.269 0.704 0.394 - - - 0.221 0.723 0.411 0.186 0.729 0.425
-Guangzhou Carbon 0.698 0.514 0.385 0.654 0.550 0.506 0.551 0.564 0.462 0.423 0.520 0.478 0.585 0.603 0.444 0.413 0.951 0.716 0.362 0.673 0.536 0.254 0.733 0.570
-Population 0.665 0.565 0.441 0.638 0.586 0.453 0.617 0.615 0.506 0.497 0.685 0.612 0.633 0.592 0.467 0.294 1.046 2.112 0.249 0.834 0.678 0.195 0.928 0.753
-GDP 0.436 0.764 0.541 0.419 0.791 0.549 0.405 1.011 0.625 0.321 1.004 0.689 0.440 0.762 0.546 0.263 2.162 1.553 0.234 1.147 0.732 0.176 1.463 0.811
-Night Light 0.577 0.573 0.418 0.534 0.596 0.442 0.405 1.011 0.625 0.335 0.672 0.514 0.483 0.633 0.478 0.436 0.549 0.427 0.225 0.720 0.562 0.176 0.798 0.620
-POI 0.396 0.752 0.344 0.325 0.771 0.364 0.143 0.822 0.411 0.136 0.823 0.415 0.150 0.819 0.408 - - - 0.108 0.920 0.433 0.084 0.956 0.469
-Shenzhen Carbon 0.659 0.568 0.418 0.613 0.570 0.434 0.564 0.576 0.471 0.483 0.586 0.488 0.562 0.571 0.483 0.234 0.975 0.746 0.207 0.735 0.597 0.189 0.887 0.634
-Population 0.790 0.448 0.343 0.724 0.476 0.384 0.624 0.562 0.471 0.589 0.604 0.483 0.727 0.510 0.392 0.294 1.046 2.112 0.264 0.871 0.694 0.175 0.921 0.772
-GDP 0.532 0.676 0.448 0.489 0.695 0.471 0.462 0.720 0.517 0.433 0.754 0.577 0.508 0.693 0.464 0.294 1.046 2.112 0.182 1.245 0.836 0.119 1.733 0.884
-Night Light 0.457 0.667 0.459 0.421 0.694 0.502 0.404 0.702 0.508 0.320 0.717 0.531 0.387 0.709 0.511 0.454 0.199 0.158 0.243 0.726 0.543 0.166 0.758 0.608
-POI 0.461 0.759 0.347 0.321 0.780 0.374 0.254 0.795 0.396 0.147 0.849 0.440 0.185 0.838 0.427 - - - 0.126 0.968 0.469 0.092 0.973 0.479
-$1^{\text{st}}$ount 53 0 0 1 5 3 0 0
-Table 3. Ablation study on UrbanSV Beijing dataset. The best results are in bold.
-Model Carbon Population GDP Night Light House Price POI
-$\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
-w/o SV branch 0.754 0.491 0.387 0.690 0.545 0.428 0.547 0.680 0.433 0.468 0.670 0.445 0.520 0.635 0.465 0.263 0.714 0.387
-w/o ST branch 0.518 0.687 0.512 0.465 0.715 0.559 0.229 0.887 0.543 0.423 0.698 0.423 0.295 0.764 0.575 0.265 0.716 0.390
-w/o LE branch 0.776 0.469 0.358 0.717 0.520 0.414 0.589 0.652 0.409 0.518 0.633 0.400 0.508 0.639 0.472 0.301 0.710 0.385
-w/o SV&ST branch 0.716 0.528 0.324 0.626 0.598 0.425 0.258 0.871 0.511 0.482 0.661 0.402 0.493 0.654 0.513 0.259 0.721 0.395
-UrbanVLP 0.788 0.455 0.455 0.720 0.517 0.410 0.585 0.651 0.420 0.523 0.634 0.398 0.523 0.634 0.407 0.319 0.706 0.383
-4.2. Experimental Setup
-4.2.1. Baselines
-We compare our method with 7 baselines in the field of urban imagery-based socioeconomic prediction.
-The single granularity imagery-based urban region representation learning methods include: AutoEncoder (Kramer, 1991), ResNet-18 (He et al., 2016), PG-SimCLR (Xi et al., 2022), UrbanCLIP (Yan et al., 2023), UrbanCLIP-SV (Yan et al., 2023).
-We also apply multiple baselines for multi-granularity urban imagery-based representation learning: Vision-LSTM (Huang et al., 2023b) and StructualUrban
-(Li et al., 2022c)
-. The in-depth overview of each baseline method is presented in the Appendix B.
-4.2.2. Metrics and Implementation
-The performance of our predictions is evaluated using a trio of standard metrics: the coefficient of determination denoted as $R^{2}$, the root mean squared error (RMSE), and the mean absolute error (MAE).
-An increase in $R^{2}$, along with a reduction in RMSE and MAE values, is indicative of enhanced model accuracy.
-Adam optimizer is chosen to minimize the training loss during parameter learning.
+In this section, we provide the detailed introduction and data source of our downstream benchmark.
+  • Carbon Emissions: Extracted from the Open-source Data Inventory for Anthropogenic CO2 (ODIAC), this dataset offers granular carbon emission figures in 2022 that are in alignment with the spatial resolution of our satellite imagery, which is one square kilometer per image.
+The emissions are quantified in tons.
+  • Population: The quantum of population serves as an indicator of a region’s sociodemographic characteristics.
+Sourced from WorldPop (https://hub.worldpop.org), we gather population distribution data in 2020, to provide an in-depth portrayal of demographic patterns.
+The unit is #citizens.
+  • GDP: This dataset encapsulates China’s economic development by incorporating Gross Domestic Product (GDP) figures, thereby infusing an economic dimension into our analysis of urban regions.
+The unit is million Chinese Yuan.
+  • Night Light: Nighttime light can, to some extent, characterize the intensity of human activities, which is of great significance for urban development studies.
+we take the nightlight data from [87] in 2020.
+  • House Price: The price of housing is is subject to the influence of numerous socio-economic factors.
+By observing fluctuations in housing prices, it is possible to discern, to a certain extent, the state of economic development.
+We collect data from Anjuke (https://www.anjuke.com) in March 2023 with yuan/$m^{2}$ unit.
+Individual house price represents a locational attribute.
+To extrapolate a regional-level characteristic, we take the average of housing prices across an urban region.
+  • POI: The spatial distribution of various types of POIs (Points of Interest) within urban areas is a potent social indicator, as it reflects the underlying social and economic architecture, as well as the population’s activities and urban lifestyles.
+Consequently, to glean insights into these urban dynamics, we collect data of POIs from web sources, encompassing both their locations and categories.
+The collected data has 14 categories: Dining and Cuisine, Leisure and Entertainment, Sports and Fitness, Business and Residential, Healthcare, Financial Institutions, Tourist Attractions, Lifestyle Services, Shopping and Consumption, Automobile Related, Hotel Accommodation, Transport Facilities, Science, Education and Culture, Companies and Enterprises.
+To capture the unique attributes of each area, we aggregate the number of POIs by category within each region.
+We construct this benchmark with the aim of conducting a thorough evaluation of the urban region representations that have been derived through our learning framework.
+We uniformly partition the downstream dataset into train:val:test=7:1:2.
+These indicators constitute the downstream task dataset, covering economic (GDP, Night Light, House Price), social (Population, POI), and environmental (Carbon Emission) indicators.
+The data preprocessing setting adheres to [74] with refinement.
+For some indicators data with relatively small differences in values, we have taken the logarithm of these values to facilitate training.
+B.3 Downstream Dataset Analysis
+Figure 12: Spatial distribution of Popultion, GDP and NightLight in Beijing.
+Here we visualize the distribution of some downstream socio-economic indicator data on Beijing in Figure 12.
+The distribution of different indicators exhibits similar trends, indicating that Beijing has a clear ’center-periphery’ structure, with the core area being densely populated and economically developed, while the peripheral areas are relatively sparse.
+Beyond geographical spatial differences, the concentration trends and variability of the data are critical perspectives that merit attention.
+In Figure 14 we demonstrate these patterns by illustrating the normalized values of different socio-economic indicators in Beijing.
+Outliers with extreme maximum and minimum values will be filtered out during our processing.
+As we can see, the values of Carbon Emission and NightLight are relatively concentrated, showing minimal variability.
+The distribution of GDP values is more widespread, while Population distribution exhibits a bimodal characteristic, reflecting the inherent diversity.
+Figure 13: Data distribution of Carbon Emission, Popultion, GDP and NightLight in Beijing.
+Figure 14: PerceptionScore distribution on $\mathtt{CityScape}$ dataset for 4 representative cities.
+Appendix C In-depth Overview of Baseline Methods
+The single granularity imagery-based urban region representation learning methods include:
+  • ViT [15]: ViT is a successful paradigm of applying transformers to computer vision, where images are segmented into patches for fixed-length modeling.
+When sufficient data is available for pre-training, ViT’s performance can surpass that of CNNs [22].
+In our experiment, we leverage ViT-B as the baseline for comparison.
+  • PG-SimCLR [73]: A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning.
+In our study, we do not employ PG-SimCLR to predict POI to ensure a fair comparison.
+  • UrbanCLIP [74]: A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
+  • UrbanCLIP-SV [74]: Original UrbanCLIP is pretrained on satellite image-text pairs.
+Here we also verify its performance on our proposed $\mathtt{CityScape}$ dataset, utilizing the same aggregation method to acquire regional features.
+We also apply multiple baselines for multi-granularity urban imagery-based representation learning:
+  • Vision-LSTM [27]: A vision long short-term memory neural network which could process variable-length street-view image within each region.
+The integration of satellite and street-view imagery, along with trajectories, facilitates a more effective capture of comprehensive urban representations, yielding reliable recognition results.
+  • StructuralUrban[41]: The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images.
+We term this method StructuralUrban for simplicity.
+We present a visual representation of the performance in the main paper Figure 4 to effectively demonstrate the comprehensive capabilities of our model across various downstream tasks.
+There is also a trend that models tend to perform better on environmental (Carbon Emission) and social (Population, POI) indicators than on economic (GDP, Night Light, House Price) indicators.
+This is possibly because environmental and social indicators usually have more distinct geographic distribution characteristics.
+For instance, neighboring regions may have similar climates, vegetation cover, and population densities, which are important factors for environmental and social indicators and can be effectively captured through spatial embedding.
+On the other hand, economic indicators are more complex and influenced by various factors, including policies, economic activities, and market sentiments, making them harder to model.
+Additionally, due to the multi-class nature of POI, models may require more complex structures to understand the underlying relationships between labels, thus resulting in lower performance.
+Appendix D Implementation Details
+In our experiment, Adam optimizer is chosen to minimize the training loss during parameter learning.
 A grid search on hyperparameters is conducted, where search ranges for learning rate and batch size are set as 2e-6, 2e-5, 2e-4, 2e-3, 2e-2 and 4, 8, 16, 32, 64, respectively.
-The data augmentation strategy follows the setting of (Radford et al., 2021).
+The data augmentation strategy follows the setting of [61].
 The weight hyperparameter $\alpha$ and $\beta$ of $\mathcal{L}_{Total}$ are set to 0.5, 0.5 respectively.
-We run all the models on NVIDIA A800 GPUs.
-We implement the aggregation method $Aggr$ through MLP layers.
-To ensure a fixed length for network processing, we set the max street-view sequence length within one as 25, and sequences shorter than that are achieved by padding with zeros.
-The fusion method is Addition.
-We utilize the current state-of-the-art GeoCLIP (Vivanco et al., 2023) as our location encoder.
-For the semantic segmentation model to decouple visual elements with street-view images, the architecture for the semantic model is ResNet18dilated + PPM_deepsup.
-The original dataset contains 150 categories, from which we select 38 categories and group them into 13 street-view features following.
-(Fan et al., 2023b).
-The detailed categories can be seen in Appendix C.
-4.3. RQ1: Performance Evaluation
-To further validate the effectiveness of the proposed UrbanVLP, we conduct comparisons with existing state-of-the-art methods using all datasets.
-The comprehensive results are presented in Table 2.
-As we can see, UrbanVLP significantly surpasses the performance of baseline methods employing both single-granularity and multi-granularity imagery, achieving a substantial margin of improvement.
-It outperforms the best baseline, StructualUrban (Li et al., 2022c), by 4.2%, 2.6%, 2.6%, 6.0%, 1.2% and 4.7% in terms of $R^{2}$ for all 6 indicators: Carbon, Population, GDP, Night Light, House Price and POI in Beijing.
-The similar performance improvement can also be seen in the dataset of the other three cities.
-Besides, the average performance gain of UrbanVLP on RMSE and MAE in Beijing data are 2.3% and 2.0%, respectively.
-In order to provide a more intuitive display, we have created radar Image in Figure 2 to visualize the performance of two cities, Beijing and Shenzhen.
-The results further prove the effectiveness of our multi-granularity framework for urban region representation learning.
-It is evident that pipelines leveraging multi-granularity approaches typically demonstrate superior performance compared to those employing a single granularity level.
-This enhancement can be rationalized by the incorporation of additional fine-grained information derived from street-view imagery.
-UrbanVLP exhibits superior performance over StructuralUrban (Li et al., 2022c) and Vision-LSTM (Huang et al., 2023b), illustrating the efficacy of textual mobility, which enhances interpretability.
-Concurrently, it is observed that UrbanCLIP-SV significantly underperforms compared to UrbanCLIP, demonstrating that street-view imagery, while rich in fine-grained details, lacks macroscopic perspective, leading to suboptimal outcomes when used in isolation.
-This also underscores the irreplaceability of satellite imagery as a macroscopic visual modality.
-4.4. RQ2: Ablation Studies
-We conduct ablation studies to investigate the effectiveness of different components in UrbanVLP on UrbanSV Beijing dataset, including the Satellite branch (ST branch), Street-View branch (SV branch), and Location Encoding branch (LE branch).
-The results are demonstrated in Table 3.
-It can be observed that the incorporation of street-view imagery results in an average improvement of 3.75% in terms of $R^{2}$.
-Furthermore, while street-view imagery contributes fine-grained information, its individual impact is still outstripped by that of satellite imagery, aligning with the performance differential between UrbanCLIP and UrbanCLIP-SV in Section 4.3.
-The role of location coordinates has been underexplored in previous work.
-Although not as impactful as visual modalities, their inclusion has yielded an average improvement of 0.5% in terms of $R^{2}$.
-In our work, we leverage ShareGPT4V (Chen et al., 2023a) for text description generation.
+We run all the models on NVIDIA A800 GPUs with PyTorch.
+We implement the aggregation method $\operatorname{Aggr}(\cdot)$ through MLP layers.
+The feature fusion method f  directly utilizes additive aggregation.
+To ensure a fixed length for network processing, we set the maximum street-view sequence length within one region as 25, and sequences shorter than that are achieved by padding with zeros.
+For the semantic segmentation model to decouple visual elements with street-view images, following  [17], the architecture for the semantic model is ResNet18dilated + PPM_deepsup.
+Initially, the dataset comprises 150 segmentation categories.
+However, in line with our requirements, we narrow down the selection to 38 categories, which we further group into 13 street-view features following [17]: Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
+Appendix E More Details for Modality Representation
+For utilizing ViT as the visual encoder to process urban visual data, in particular, we reshape the input image $I\in\mathbb{R}^{H\times W\times 3}$ into a sequence of flattened 2D patches ${I}_{P}\in\mathbb{R}^{N\times({P^{2}}\cdot{C})}$, where C represents the number of channels, $(P,P)$ specifies the resolution of each image patch, and $N=HW/P^{2}$ denotes the number of patches, corresponding to the input sequence length for ViT.
+Notebly, the ViT utilizes a consistent latent vector size (denoted as $d$) across all its layers.
+Thus, we can obtain the outputs as the patch embeddings through a trainable linear projection.
+We also prepend a learnable embedding to the patch embedding sequence (i.e., $\mathbf{z}_{0}^{0}={I}_{cls}$).
+Besides, we add learnable position embedding $\mathbf{E}_{pos}$ to retain positional information.
+The process is as follows:
+$\mathbf{z}_{0}=[{I}_{cls};{I}_{p}^{1}\mathbf{W};{I}_{p}^{2}\mathbf{W};\cdots;{I}_{p}^{N}\mathbf{W}]+\mathbf{E}_{pos},$ (8)
+where $\mathbf{W}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$ and $\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times d}$ are learnable parameters.
+$\mathbf{z}_{0}$ represents the result of the linear projection, which is subsequently fed into the ViT.
+Consistent with the standard Transformer encoder, ViT includes alternating layers of multi-head self-attention $\operatorname{MSA(\cdot)}$ operation [68].
+Furthermore, layer normalization (LN) is adopted before every block, whereas residual connections are after every block:
+$\mathbf{z}_{l}=\mathrm{MSA}(\mathrm{LN}(\mathbf{z}_{l-1}))+\mathbf{z}_{l-1},\quad l=1\ldots L.$ (9)
+The text encoder follows a similar $\operatorname{MSA}$ mechanism of the visual encoder.
+The input text sequence is bracketed with $[SOS]$ and $[EOS]$ tokens, and the activation of the highest layer of Transformer at $[EOS]$ token is considered the global representation $\mathbf{z}_{T}$ of text.
+Appendix F Segmentation Ratio and Prompt Examples
+As illustrated in Figure 15(a) and (b), we demonstrate examples of the segmentation ratio of street-view images in our $\mathtt{CityScape}$ dataset.
+The segmentation ratio is decoupled by pretrained segmentation models and is used in both prompt construction and PerceptionScore computation.
+We group the segmentation categories into 13 categories following [17]: Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
+The prompt template used to generate street-view descriptions is formulated as follows: Analyze the street-view panoramic image in [City] in a comprehensive and detailed manner.
+The coordinate of the street-view image is [Longitude, Latitude].
+The segmentation ratio of the street-view image is [Segment Information].
+Figure 15(c) provides illustrations of some examples of our prompts used for description generation.
+Figure 15: Prompt Examples.
+Appendix G The Choice of Image-to-Text LMM
+In our work, we leverage ShareGPT4V [11] for text description generation.
 Here we conduct experiments on current open-sourced LMM models, and compute the PerceptionScore of them.
 The results are presented in Table LABEL:table:lmms.
-As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which are used as the Image-to-Text models (Yan et al., 2023).
-Table 4. Perception score of representative LMMs.
+As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which is used as the Image-to-Text models in [74].
+Additionally, we note that there is study [86] argue that InstructBLIP [14] exhibits state-of-the-art (SOTA) performance in image-to-text generation.
+However, in our experiment, we found that InstructBLIP is unsuitable for handling geospatial coordinates as prompt content.
+Incorporating such coordinates significantly degrades the quality of the generated text, leading to poor content output.
+Due to these limitations, we have excluded it from our comparative analysis.
+Table 3: Perception score of representative LMMs.
 Models Average Generated-Description Scores
-LLaVA-v1.5 (Liu et al., 2023a) 0.674
-mPLUG-Owl (Ye et al., 2023) 0.662
-LLaMA-Adapter V2 (Gao et al., 2023) 0.698
-MiniGPTv2 (Chen et al., 2023b) 0.693
-ShareGPT4V (Chen et al., 2023a) 0.714
-Figure 5.
-Illustration of the quality of our generated description example.
-The prompt for (c) is :”From scale 1 to 10, decide how similar the image is to the description”, whereas the prompt for (d) is: ”Generate a panoramic street-view image based on the description”.
-Figure 6. Perception Score distribution on UrbanSV dataset for four representative cities.
-4.5. RQ3: Qualitative Analysis
-To visually illustrate the quality of the generated description, we present an example in Figure 5.
-The generated description of the street-view image (a) is depicted in (b).
-Subsequently, we employ GPT4V (Yang et al., 2023) to assess the quality score of the generated description, which is rated as 7 out of 10, indicating the effectiveness of the generated text.
-Furthermore, GPT4V also provides specific areas for improvement in (c).
-In (d), we utilize GPT4V to generate an image based on the description in (b).
-It is evident that the generated image bears a resemblance to the original one.
-In order to ascertain the efficacy of ShareGPT4V, we computed the distribution of PerceptionScores for descriptions generated by ShareGPT4V, as depicted in Figure 6.
-Notably, only less than 0.02% of image-text pairs scored below 0.5, thereby evidencing the high quality of the descriptions produced.
-Figure 7. Examples of predictions from UrbanCLIP and our proposed UrbanVLP.
-Here we also show some predicted results in Figure 7 on UrbanSV Beijing dataset.
-As we can see, satellite image (a) and (b), despite sharing similar layouts, (a) encompasses residential and campus areas, while (b) pertains to industrial zones.
-Their socio-economic characteristics are entirely different, which could be reflected by socio-enconomic indicators.
-The carbon emissions and GDP of campus and residential area are significantly lower than those of industrial park, whereas the disparities in population and housing prices are not as pronounced, possibly due to the unique characteristics of school district housing near educational institutions.
-Night light intensity is similar across the two area due to the fact that the majority of nocturnal economic activities within factories occur indoors, unlike in Central Business Districts (CBDs) where extensive illumination is generated.
-The same principle applies to school and residential area as well.
-As observed, the predictions made by UrbanCLIP do not effectively differentiate between the two, with the downstream metrics in Figure 7 (b) still inclined towards predicting a homogenized socio-economic functional area similar to that of (a).
-In contrast, the results from UrbanVLP are capable of distinctly distinguishing the socio-economic attributes of the two in terms of carbon emission, GDP, and house price.
-5. Conclusion and future work
-Urban region representation learning plays a significant role in helping us understand the patterns, structures, and dynamics of cities.
-UrbanVLP, for the first time, explores the differences between street-view and satellite images from a semantic granularity perspective, as well as their roles in modeling urban region representation.
-A text generation and calibration mechanism is also proposed to ensure high-quality description generation.
-It has achieved state-of-the-art results on the constructed UrbanSV and downstream dataset.
-In the future, the further research direction can be two-fold: Add additional modalities such as POIs to enhance information richness and enable the design of better model architectures such as MAE (He et al., 2022) to efficiently utilize existing data.
-Appendix A Detailed introduction of downstream dataset
-We provide the detailed introduction and data source of our downstream benchmark.
-  • Carbon Emissions: Extracted from the Open-source Data Inventory for Anthropogenic CO2 (ODIAC), this dataset offers granular carbon emission figures in 2022 that are in alignment with the spatial resolution of our satellite imagery, which is one square kilometer per image.
-The unit is ton.
-  • Population: Sourced from WorldPop, this data provides population distribution data in 2020. The unit is #citizens.
-  • GDP: This data reflects the economic development situation in China in terms of Gross Domestic Product (GDP) figures to add an economic layer to our urban region analysis.
-  • Night Light: Nighttime light can, to some extent, characterize the intensity of human activities, which is of great significance for urban development studies.
-we take the nightlight data from (Zhong et al., 2022) in 2020.
-  • House Price: The price of housing is influenced by a variety of socio-economic factors.
-By observing changes in housing prices, it is possible to discern, to a certain extent, the state of economic development.
-We collect data from Anjuke (https://www.anjuke.com) in 2023.3.
-The unit is yuan/$m^{2}$.
-We take the average of housing prices within an urban region.
-  • POI: The distribution of different types of POIs (Points of Interest) within urban areas can serve as a social indicator, as they reflect the social and economic structure, population activities, and lifestyles of the city.
-We collect POIs (location and category) from the web.
-The data has 14 categories: Dining and Cuisine, Leisure and Entertainment, Sports and Fitness, Business and Residential, Healthcare, Financial Institutions, Tourist Attractions, Lifestyle Services, Shopping and Consumption, Automobile Related, Hotel Accommodation, Transport Facilities, Science, Education and Culture, Companies and Enterprises.
-We take aggregation of the number of POIs of each category within each region.
-We construct this benchmark to comprehensively evaluate the learned urban region representation.
-We uniformly partition the downstream dataset into train:val:test=6:2:2.
-These indicators constitute the downstream task dataset, covering economic (GDP, Night Light, House Price), social (Population, POI), and environmental (Carbon Emission) indicators.
-The data preprocessing setting adheres to (Yan et al., 2023) with refinement.
-Appendix B In-depth overview of baseline methods
-The single granularity imagery-based urban region representation learning methods include:
-  • AutoEncoder (Kramer, 1991): An architectural framework for neural networks that is designed to extract representations from unlabeled satellite imagery with reconstruction loss.
-  • ResNet-18 (He et al., 2016): The well-known residual neural network pre-trained on ImageNet (Deng et al., 2009), which can directly extract visual features of satellite imagery with the knowledge acquired from natural imagery.
-  • PG-SimCLR (Xi et al., 2022): A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning.
-We don’t use PG-SimCLR to predict POI for a fair comparison in our study.
-  • UrbanCLIP (Yan et al., 2023): A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
-  • UrbanCLIP-SV (Yan et al., 2023): Original UrbanCLIP is pre-trained on satellite image-text pairs.
-Here we also verify its performance on our proposed UrbanSV dataset, utilizing the same aggregation method to acquire regional features.
-We also apply multiple baselines for multi-granularity urban imagery-based representation learning:
-  • Vision-LSTM (Huang et al., 2023b): A vision long short-term memory neural network which could process variable-length street-view image within each region.
-The integration of satellite and street-view imagery enables a more effective capture of comprehensive urban representation, achieving reliable recognition results.
-  • StructualUrban(Li et al., 2022c): The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images.
-We term this method StructualUrban for simplicity.
-Appendix C Segmentation Ratio and Prompt Examples
-As illustrated in Figure 8, we demonstrate examples of the segmentation ratio of street-view images in our UrbanSV dataset.
-The segmentation ratio is decoupled by pre-trained segmentation models and is used in both prompt construction and perception score computation.
-We group the segmentation categories into 13 categories following (Fan et al., 2023b): Light Vehicle; Façade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
-Figure 9 illustrates some exmaples of our prompts used for description generation.
-Figure 8. Segmentation Ratio.
-Figure 9. Prompt Examples.
-Appendix D RQ4: Practicality
-Figure 10. UrbanVLP system.
+LLaVA-v1.5 [48] 0.674
+mPLUG-Owl [77] 0.662
+LLaMA-Adapter V2 [20] 0.698
+MiniGPTv2 [10] 0.693
+ShareGPT4V [11] 0.714
+Appendix H More Experimental Analysis
+H.1 More Analysis for Text Calibration
+In our paper, we selected the state-of-the-art LMM ShareGPT4V, which demonstrates performance comparable to GPT4V in image captioning.
+In order to ascertain the efficacy of ShareGPT4V, we computed the distribution of PerceptionScore for descriptions generated by ShareGPT4V, as depicted in Figure 14.
+Notably, only less than 0.02% of generated texts scored below 0.5, thereby evidencing the high quality of the descriptions produced.
+Table 4: UrbanVLP performance utilizing mPLUG-Owl with and without Text Calibration. mPO is an abbreviation for mPLUG-Owl, and TC stands for Text Calibration.
+Model Carbon Population GDP
+$\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
+UrbanVLP_mPO_wo TC 0.687 0.551 0.544 0.643 0.580 0.458 0.494 0.655 0.447
+UrbanVLP_mPO 0.714 0.538 0.473 0.683 0.557 0.430 0.506 0.640 0.435
+UrbanVLP 0.769 0.477 0.369 0.714 0.523 0.411 0.537 0.684 0.416
+As a result, the experimental results comparing with and without calibration do not adequately reflect the effectiveness of our calibration method.
+To further demonstrate the effectiveness of our method, we select mPLUG-Owl [77], one of the representative LMMs compared in our paper, with the worst performance.
+After calibration, 15.3% of the text with perceptionscores below 0.5 was filtered out.
+The results validate the effectiveness of our calibration method, and indicate that more relevant and noise-free textual information could align better with image features, leading to a more coherent and meaningful visual representation.
+H.2 Analysis about Time Complexity
+Modeling and research in urban areas can provide valuable references and assistance to policymakers, builders, and citizens.
+Therefore, practical implementation should be considered.
+We conduct time complexity analysis using FPS (Frames Per Second), a common indicator for inference speed in computer vision area.
+We report FPS for satellite image in Figure 16(b).
+In fact, introducing street view features for better and finer-grained effects will inevitably sacrifice a portion of the time efficiency.
+However, compared to [41]’s graph structure inference, our linear probing setting ensures the minimum time overhead, also with better performance.
+For making model decisions during deployment, to pursue the trade-off between performance and speed, UrbanVLP is the optimal choice.
+Figure 16: (a) Hyperparameter for loss combination. (b) FPS comparison of different mdoels. (c) Expetiments for maximum of the number of street-view images.
+H.3 Analysis about Hyperparameters
+The optimal hyperparameters for Equation 7 are actually determined by grid search.
+In Figure 16(a), we present some results of different combinations of hyperparameter values .
+It can be seen that the settings we chose in the paper ($\alpha$=0.5, $\beta$=0.5) yield the best experimental results.
+We also study the maximum value of street-view images during training, which is set as 25 in our paper.
+As is shown in Figure 16(c), as the number of street-view images increases, the R² value continuously rises, reaching its peak at 25.
+Therefore, our setting is the optimal choice.
+Appendix I Practicality
+Figure 17: UrbanVLP system.
 To demonstrate the practical applications of our model in real-life scenarios, we have developed a system bearing the same name UrbanVLP.
-The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban indicator framework using the Mapbox platform (Mapbox, [n.
-d.]).
+The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban indicator framework using the Mapbox platform [56].
 This system renders a visual representation of urban areas through satellite imagery, complemented by street-level photographs for an interactive exploration experience.
 Figure 17 showcases the user-friendly interface of the map, highlighting functionalities that facilitate zooming, searching for locations, and navigating through various sectors.
-Each urban region captured by a satellite image, encompassing a 1km by 1km area, is marked by a blue dot, which is encircled by yellow points that denote the locations of street-view images accessible within that region.
+Each urban region captured by a satellite image, encompassing a 1km $\times$ 1km area, is marked by a blue dot, which is encircled by yellow points that denote the locations of street-view images accessible within that region.
 Engaging with these points allows users to uncover a suite of urban metrics, such as carbon emissions, population, GDP, night light, house price, and POI.
@@
 Ultimately, UrbanVLP is designed to distill complex urban data into a visual and user-friendly format, thereby offering a holistic and detailed view of urban landscapes and their key indicators.
+Appendix J Limitations and Social Impact
+Due to the easier accessibility of the dataset, our work primarily focuses on first-tier and domestic cities, which introduces a certain bias and can be regarded as a limitation of our study.
+Moreover, the model’s performance is significantly influenced by image resolution; higher quality image data is required to achieve optimal performance.
+Additionally, the introduction of new modalities inevitably leads to a decrease in inference efficiency, which should be a consideration for practical deployment.
+Moreover, the inability to automatically correct low-quality text highlights an area for future improvement.
+Our work can be deployed in practical applications, as demonstrated by the example provided in the paper, thereby aiding urban planning and decision-making processes.
+However, it is crucial to note that satellite and street-view images might inadvertently contain private information.
+The street-view images we collected have been blurred by the Baidu Map to obscure relevant information, such as license plates, thereby significantly reducing associated risks.
+Nevertheless, we strongly advise users to perform a comprehensive analysis of the data prior to deploying our model in practical contexts.

### v3

@@
-UrbanVLP: Multi-Granularity Vision-Language Pretraining for Urban Region Profiling
+UrbanVLP: Multi-Granularity Vision-Language Pretraining for Urban Socioeconomic Indicator Prediction
 Abstract
-Urban region profiling aims to learn a low-dimensional representation of a given urban area while preserving its characteristics, such as demographics, infrastructure, and economic activities, for urban planning and development.
+Urban socioeconomic indicator prediction aims to infer various metrics related to sustainable development in diverse urban landscapes using data-driven methods.
 However, prevalent pretrained models, particularly those reliant on satellite imagery, face dual challenges.
 Firstly, concentrating solely on macro-level patterns from satellite data may introduce bias, lacking nuanced details at micro levels, such as architectural details at a place.
-Secondly, the lack of interpretability in pretrained models limits their utility in providing transparent evidence for urban planning.
+Secondly, the text generated by the precursor work UrbanCLIP, which fully utilizes the extensive knowledge of LLMs, frequently exhibits issues such as hallucination and homogenization, resulting in a lack of reliable quality.
 In response to these issues, we devise a novel framework entitled UrbanVLP based on Vision-Language Pretraining.
 Our UrbanVLP seamlessly integrates multi-granularity information from both macro (satellite) and micro (street-view) levels, overcoming the limitations of prior pretrained models.
-Moreover, it introduces automatic text generation and calibration, elevating interpretability in downstream applications by producing high-quality text descriptions of urban imagery.
-Rigorous experiments conducted across six urban indicator prediction tasks underscore its superior performance.
-1 Introduction
-Urban Region Profiling, a scholarly endeavor within the domain of GeoAI [45, 54] and smart city [85], refers to the process of transforming urban regions into a low-dimensional vector representation using deep neural networks.
-Rooted in the escalating global prominence of urban environments and the imperative for urban planning, this academic discipline aims to capture and encode the multifaceted attributes of urban areas, such as demographic information, economic activity, environmental factors, infrastructure, and social dynamics, into a format that can be easily analyzed and utilized by computational models [73, 74].
-Given the widespread accessibility of satellite imagery on platforms such as Google Maps [42, 51, 73], coupled with its wealth of information on regional features (e.g., road networks, building density, and vegetation coverage), a predominant majority of initiatives harness satellite imagery as the foundational modality for learning urban region representations [88].
-Through a comprehensive review of the current literature, we summarize mainstream methods for urban region profiling into two categories: a) Task-specific supervised learning commonly realizes urban region profiling in a fully supervised task, e.g., identifying poverty levels [21, 59, 4, 5], crop yields [80, 70, 63, 57, 53, 78], and commercial activeness [24, 51].
+Moreover, it introduces automatic text generation and calibration, providing a robust guarantee for producing high-quality text descriptions of urban imagery.
+Rigorous experiments conducted across six socioeconomic indicator prediction tasks underscore its superior performance.
+Codehttps://github.com/CityMind-Lab/UrbanVLP
+Introduction
+Urban Socioeconomic Indicator (USI) Prediction, a scholarly pursuit in the realm of urban computing (Zou et al.
+2024), deploys data-driven methodologies to forecast socioeconomic metrics (e.g., GDP, population, carbon emission).
+Rooted in the escalating global prominence of urban environments and the imperative for judicious urban planning, this academic discipline finds motivation in enhancing the efficacy of policy-making, optimizing resource allocation, and mitigating challenges endemic to sustainable development (Xi et al.
+2022; Yan et al.
+2024).
+Figure 1: USI prediction frameworks. Compared to existing arts, we present the first attempt to introduce multi-granular visual information and high-quality calibrated texts.
+Given the widespread accessibility of satellite imagery on platforms such as Google Maps (Li et al.
+2022c; Liu et al.
+2023d; Xi et al.
+2022), coupled with its wealth of information on regional features (e.g., road networks, building density, and vegetation coverage), a predominant majority of initiatives harness satellite imagery as the foundational modality for learning urban region representations and making predictions (Zhang et al.
+2024).
+Through a comprehensive review of the current literature, we summarize mainstream methods for USI prediction into two categories:
+  • Task-specific supervised learning commonly realizes USI prediction in a fully supervised task, e.g., identifying poverty levels (Han et al.
+2020; Ayush et al.
+2020, 2021), crop yields (Rußwurm and Körner 2020; Martinez et al.
+2021; M Rustowicz et al.
+2019; Yeh et al.
+2021), and commercial activeness (He et al.
+2018; Liu et al.
+2023d).
 However, its inherent task specificity (i.e., reliant on ample labeled data), may impede the model’s capacity for broad generalization to other downstream tasks.
-b) Vision pretraining, also known as a type of Self-Supervised Learning [28], aims to learn general visual features from urban satellite imagery, which are subsequently fine-tuned on a specific task for enhancing performance  [29, 6, 73, 51].
-The inclusion of more auxiliary data, e.g., Points of Interests (POIs) [19, 29, 82, 40], road networks [46, 47, 12] and human mobility [71, 64, 44, 39], leads to richer, more accurate, and more useful representations of urban areas.
+  • Vision pretraining, also known as a type of Self-Supervised Learning (Jaiswal et al.
+2020), aims to learn general visual features from urban satellite imagery, which are subsequently fine-tuned on a specific task for enhancing performance  (Bai et al.
+2023; Xi et al.
+2022; Liu et al.
+2023d).
+The inclusion of more auxiliary data, e.g., Points of Interests (POIs) (Zhang et al.
+2021; Li et al.
+2024b) and road networks (Liang et al.
+2018, 2021; Chen et al.
+2024), leads to richer, more accurate, and more useful representations of urban areas.
 For clarity, Figure 1 (a-b) depict a sketch of these two streams of approaches.
-Despite the success of using satellite imagery for urban region profiling, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture).
-Previous approaches in Figure 1(a-b) mainly focus on single granularity, neglecting finer-grained visual clues.
+Despite the success of utilizing satellite imagery for USI prediction, urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture).
+Previous frameworks in Figure 1 (a-b) mainly focus on single granularity, neglecting finer-grained visual clues.
 Nonetheless, as the saying goes, “the devil is in the details.”.
-By zooming into micro levels using corresponding street-view images, a more nuanced and fine-grained understanding emerges, as shown in the left hand side of Figure 1(d).
-Therefore, the integration and alignment of multi-granularity information, both macro-level and micro-level, remain ongoing challenges in urban region profiling that require further exploration.
-Figure 1: Frameworks for urban region profiling.
-Compared to existing literature, we present the first attempt to introduce multi-granular visual information and high-quality calibrated texts for this task.
-Furthermore, while urban planners rely on explainable evidence for service deployment and urban planning, the vision-pretraining-based models for urban region profiling mostly lack interpretability.
-They typically learn a mapping from single/multi-modal data to the representations, posing a significant challenge in comprehending the intrinsic factors influencing the prediction outcomes.
-In light of the recent surge in Large Language Models (LLMs) [58, 84, 55], UrbanCLIP [74] stands out as an innovative solution, which effectively integrates textual modality as auxiliary knowledge into urban region profile, as depicted in Figure 1(c).
-However, the text generation process of UrbanCLIP raises several critical concerns.
-i) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery.
+By zooming into micro levels using corresponding street-view images, a more nuanced and fine-grained understanding emerges, as shown in the Figure 1(d) Left.
+Therefore, the integration and alignment of multi-granularity information remain ongoing challenges in USI prediction that require further exploration.
+In recent years, by leveraging the extensive knowledge embedded in LLMs (Manvi et al.
+2023) and the inherent interpretability of text modality, a wide range of tasks across various domains are empowered.
+Textual data can serve as an effective auxiliary tool in multimodal learning across a wide range of scenarios, such as geo-localization (Li et al.
+2024a) and recommender systems (Gao et al.
+2024).
+While in USI prediciton, the efficacy of leveraging textual modality for semantic enrichment is still less explored.
+UrbanCLIP (Yan et al.
+2024) stands out as the innovative solution, which generates text description by a pretrained LLM for satellite imagery and achieves learning urban region representations through natural language supervision.
+Nevertheless, the text generation process of UrbanCLIP raises several critical concerns: i) Hallucination: The generation of textual content sometimes deviates from or introduces information not present in the input satellite imagery.
 ii) Homogenization: The generated descriptions appear overly simplified and general, potentially leading to homogenization and inadequate differentiation.
 To better align with the intricacies of satellite (or street-view) imagery, we necessitate a more powerful approach that goes beyond the intuitive text generation method in UrbanCLIP, ensuring a more faithful and explainable representation of urban regions.
-In this paper, we present a Vision-Language Pretraining framework (i.e., UrbanVLP) for urban region profiling.
+In this paper, we present a Vision-Language Pretraining framework (i.e., UrbanVLP) for urban socioeconomic indicator prediction.
 As depicted in Figure 1(d), our model elaborately integrates multi-granularity information from both satellite (macro-level) and street-view (micro-level) imagery to produce comprehensive urban region representations, while simultaneously harnessing the interpretability inherent in high-quality textual descriptions.
 Targeting the first challenge, we introduce a novel Multi-Granularity Cross-Modal Alignment module, which utilizes dual-branch contrastive learning to establish alignment between information derived from two semantic granularities.
-To address the second issue (i.e., lack of interpretability), we devise an Automatic Text Generation together with a Calibration mechanism to uphold text quality standards.
-Given the impracticality of generating a large number of descriptions using GPT4V [75] or Gemini [18] due to prohibitive expense, we utilize a closely comparable GPT4V [75] model called ShareGPT4V [11].
+To address the second issue (i.e., hallucination and homogenization in LLM generated texts), we devise an Automatic Text Generation together with a Calibration mechanism to uphold text quality standards.
 To guarantee the quality of LLM-generated descriptions, we propose a reference-free metric called PerceptionScore, motivated by the human evaluation system.
-In summary, our major contributions lie in the following aspects:
+In summary, our contributions lie in the following aspects:
   • Multi-Granularity Cross-Modal Alignment.
 We explore the role of two distinct visual data modalities at various semantic granularities – satellite imagery and street-view imagery.
-We integrate street-view data to inject fine-grained semantic information through token-level contrastive learning.
-By aggregating and fusing local features within a specific region with satellite imagery data, we facilitate the development of a robust and comprehensive representation of urban landscapes.
+We inject fine-grained semantic information by integrating street-view data through token-level contrastive learning.
   • Automatic Text Generation and Calibration.
-Powered by image-to-text LLMs, we generate text descriptions and implement a robust evaluation mechanism based on a newly-proposed reference-free metric, ensuring the fidelity between the generated texts and the corresponding image content.
-  • Empirical Evidence.
-We evaluate the proposed UrbanVLP on six urban indicator prediction tasks (e.g., GDP, population, and carbon emission) in diverse urban landscapes.
-Extensive experiments demonstrate that our UrbanVLP outperforms existing approaches by an average improvement of 3.95% on the $R^{2}$ metric, while preserving good interpretability based on generated text.
-  • New Benchmark and System.
-We will open-source the first multi-granular vision-text benchmark ($\mathtt{CityScape}$) for urban region profiling upon paper notification, encompassing 6 downstream tasks.
-Moreover, we have deployed a web platform to verify the model’s practicality (see Appendix I).
-2 Related work
-We formally define the problem of urban region profiling.
-Given a satellite images $I^{sa}_{g}$, a set of street-view images $\mathcal{I}^{sv}_{g}$ belonging to the target coverage, and their corresponding latitude and longitude pair $\mathcal{L}_{g}$.
+Powered by image-to-text LLMs, we generate text descriptions and implement a robust evaluation mechanism based on a new reference-free metric, ensuring the fidelity between the generated texts and the corresponding image content.
+  • A New Benchmark & Empirical Evidence.
+We plan to open-source the first vision-text and multi-granularity urban dataset upon paper notification, consisting of six downstream tasks across the socio-economy.
+Extensive experiments demonstrate that our UrbanVLP outperforms existing approaches by an average improvement of 3.95% on the $R^{2}$ metric, as illustrated in Figure 2.
+Figure 2: $R^{2}$ results in Beijing and Shenzhen.
+Figure 3: Overall framework of our proposed UrbanVLP.
+Related Work
+We formally define the problem of USI prediction.
+Given a satellite image $I^{sa}_{g}$, a set of street-view images $\mathcal{I}^{sv}_{g}$ belonging to the target coverage, and their corresponding latitude and longitude pair $\mathcal{L}_{g}$.
 The objective is to employ a learning function $\mathcal{F}$ to map them to representation vectors $\mathbf{e}_{g}=\mathcal{F}(I^{sa}_{g},\mathcal{I}^{sv}_{g},\mathcal{L}_{g})$.
-The representation can then be further utilized in downstream tasks, such as inferring urban indicators $\mathbf{Y}_{g}$ for any given region $g$.
-We assume that during the training phase, automated additional text descriptions are available.
+The representation can then be further utilized in downstream tasks, such as inferring socioeconomic indicators $\mathbf{Y}_{g}$ for the given region $g$.
 The detailed formulation and definitions related to this task are provided in Appendix A.
-Next, we introduce the related work from three aspects.
-Urban Region Profiling.
-In literature, many studies have focused on learning task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility [41, 51, 73].
-For example, Urban2Vec [72] integrated street-view imagery and POI data to learn neighborhood embeddings.
-Some contrastive learning approaches like PG-SimCLR [73] and UrbanCLIP [74] have shown success in representing urban regions through satellite images.
+Urban Socioeconomic Indicator Prediction.
+In literature, many studies have focused on learning task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility (Liu et al.
+2023d; Xi et al.
+2022).
+For example, Urban2Vec (Wang, Li, and Rajagopal 2020) integrated street-view imagery and POI data to learn neighborhood embeddings.
+Some contrastive learning approaches like PG-SimCLR (Xi et al.
+2022) and UrbanCLIP (Yan et al.
+2024) have shown success in representing urban regions through satellite images.
 However, the aforementioned works exclusively considered a single type of urban imagery, overlooking the potential synergy between satellite imagery and street-view imagery, which can complement each other.
-Recently, some efforts [7, 27, 41, 9, 33, 8, 27] have begun to explore the simple combination of satellite imagery and street-view imagery.
-Vision-LSTM [27] integrates diverse street-level images into the satellite image domain using an LSTM-based network for precise urban village identification.
-Li et al.
-[41] propose utilizing street segments for region modeling and predicting socio-economic indicators by merging street-view and satellite images through a multi-stage pipeline.
-Nevertheless, prior studies mostly overlook distinct features in satellite and street-view images for varied granularities.
-To this end, we propose a unified multi-granularity cross-modal alignment framework for comprehensive urban representation in this study.
 Vision-Language Pretraining (VLP).
 VLP aims to jointly encode vision and language in a fusion model.
-Early works on VLP can be broadly categorized into single-stream [13, 31, 35] and two-stream [66, 52] methodologies, with the former using a unified architecture for modality integration and the latter employing separate encoders before merging modalities.
-Milestone work CLIP [61] and its variants [43, 76] highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval.
-Recent works [67, 2] shift towards leveraging LLMs knowledge for vision-language representation learning.
-In urban region profiling, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information.
-Our groundbreaking work introduces multiple types of text descriptions as supplements, aiming to learn more comprehensive and interpretable urban region representations.
-Large Multimodal Model (LMM).
-As LLMs undergo rapid evolution, a faction within the research community is increasingly focused on integrating visual knowledge into these models, such as CLIP [61] and ALIGN [30].
-Early studies [36, 37] enhanced CLIP with refined data strategies for more diverse datasets, which have proven effective for fundamental visual tasks [38, 50, 81] but have shown limitations for complex tasks such as visual question answering.
-Recently, new models like MiniGPT-4 [10], LLaVA [49], and InstructBLIP [14] have improved the understanding of complex problems by expanding parameters and training data, called large multimodal models (LMMs).
-In this paper, we comprehensively evaluate the text generation quality of representative LMMs model and propose an automated text description generation and calibration technique to ensure effective image description capabilities with LMMs.
-Figure 2: Overall framework of our proposed UrbanVLP.
-3 Methodology
-We present an overview of our proposed framework in Figure 2, which is decoupled into two stages.
-In the pretraining stage, we devise an automatic text generation and calibration module using ShareGPT4V [11] to generate textual descriptions for street-view images with geographical and visual prompts.
-Introducing the PerceptionScore to ensure high-quality text, we then develop a multi-granularity cross-modal alignment framework through multi-level contrastive learning and fine-grained information injection.
-In the fine-tuning stage, we use frozen encoders from Stage 1 to extract data features, fine-tune a lightweight MLP for precise predictions, and incorporate optional street-view imagery and geographical coordinates to enrich the analysis.
-3.1 Automatic Text Generation and Calibration
+A milestone work CLIP (Radford et al.
+2021) and its variants (Li et al.
+2021; Yao et al.
+2022) highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval.
+Recent works (Tsimpoukelli et al.
+2021; Alayrac et al.
+2022) shift towards leveraging LLMs knowledge for VLP.
+In USI prediction, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information.
+Methodology
+Figure 3 shows our framework with two stages.
+In the pretraining stage, we first devise an automatic text generation and calibration module using ShareGPT4V (Chen et al.
+2023b) to generate textual descriptions for street-view images with geographical and visual prompts.
+To guarantee the quality of the generated text, we introduced a novel metric called PerceptionScore.
+Then a multi-granularity cross-modal alignment framework is designed to utilize multi-level contrastive learning and fine-grained information injection.
+During the fine-tuning stage, we employ frozen encoders from Stage 1 to extract features and fine-tune a lightweight MLP for accurate predictions.
+Automatic Text Generation and Calibration
 Text Generation.
-For each street-view image, we utilize advanced Large Multimodal Models (LMMs) due to their robust cross-modal knowledge storage capabilities, to provide comprehensive textual descriptions.
+For each street-view image, we use advanced Large Multimodal Models (LMMs) due to their robust cross-modal knowledge retention abilities, to provide comprehensive textual descriptions.
 However, this confronts two primary challenges.
-Firstly, due to substantial financial costs associated with API usage, employing closed source LMMs like GPT4V [75] or Gemini [18] for generating tens of thousands of text descriptions is impractical.
-Secondly, street-view images often encompass diverse detailed information, necessitating a well-designed prompt template for obtaining high-quality textual descriptions.
-To address these issues, we first utilize a closely related GPT4V model, known as ShareGPT4V [11], to generate street-view image descriptions.
-This model is a powerful language model with vision capabilities trained on 100k GPT4V generated captions, demonstrating captioning capabilities comparable to GPT4V.
-In the process of designing text prompt templates, the proportion of segmented elements in street-view images reflects the distribution of elements, aiding LMMs in evaluating the importance of each element.
-To achieve this, we employ a pretrained segmentation model to decouple visual elements and calculate the proportion of segmentation for each element.
-Simultaneously, urban and geospatial coordinates of street-view images are included as prompts to guide the expression of additional urban information.
-Template and segmentation ratio details can be found in Appendix F.
+Firstly, due to substantial financial costs associated with API usage, employing closed source LMMs like GPT4V (Yang et al.
+2023) or Gemini (Fu et al.
+2023) for generating tens of thousands of text descriptions is impractical.
+Secondly, street-view images often encompass diverse details, necessitating a well-designed prompt template for obtaining high-quality textual descriptions.
+To address these issues, we first employ an open-source model closely related to GPT-4V, known as ShareGPT4V (Chen et al.
+2023b), to generate the street-view image descriptions.
+This model is a powerful LLM with vision capabilities trained on 100k GPT4V-generated captions, revealing captioning ability comparable to GPT4V.
+Moreover, in the process of designing text prompt templates, we believe that the proportion of each element in street-view images reflects their relative abundance, thereby facilitating LMMs in evaluating their significance.
+Therefore, we employ a pretrained segmentation model (CSAILVision 2024) to decouple visual elements and calculate the proportion of segmentation for each element.
+As evidenced by existing research (Fan et al.
+2023b), the accuracy of current segmentation methods is already adequate to recognize various visual components.
+Simultaneously, geospatial coordinates of street-view images are incorporated as prompts to aid in the generation of more precise textual descriptions.
+Template and segmentation ratio details can be found in Appendix H.
 Text Calibration.
-As existing works [32, 16, 74] demonstrated that the quality of the textural descriptions is crucial for enhancing model performance.
-However, previous research [74] primarily employed simplistic rules and manual refinement processes for text description of satellite images, leading to potential unresolved hallucination issues [62] in generated text descriptions.
-Besides, rule-based methods are overly general and fail to address specific issues in each image description as they are tailored to particular cases.
-On the other hand, manual rewriting is highly labor-intensive.
-Therefore, an effective method for evaluating text quality and calibrating the alignment between text and image content is necessary to ensure a high level of LMM-inherent knowledge injection.
-In contrast to the common evaluation methods used to assess the quality of descriptions generated by LMMs, which typically rely on manually annotated image descriptions as references [65, 83, 34] to measure content overlap, we introduce an automatic mechanism designed to simulate human evaluation systems.
-This mechanism places greater emphasis on semantic similarity rather than merely assessing typical content overlap, thereby introducing PerceptionScore as a novel quality metric.
-Specifically, it comprises two components: text semantic quality and visual recall quality.
-The former approach directly adopts the efficient implementation of CLIPScore [25], which is based on CLIP’s [61] strong zero-shot capabilities and demonstrates high correlation with human semantic judgment.
-However, CLIPScore solely provides a measure of textual quality and fails to indicate whether specific elements in the image are omitted in the text (visual recall) [26].
+Existing works (Fan et al.
+2023a; Yan et al.
+2024) have already demonstrated that the quality of the textural descriptions is crucial for enhancing model performance.
+However, previous research (Yan et al.
+2024) primarily employs simplistic rules and manual refinement processes for refining text description of satellite imagery, leading to potential unresolved hallucination problems (Rawte, Sheth, and Das 2023).
+Besides, the rule-based methods are overly general and fail to address specific issues in each image description as they are tailored to particular cases.
+On the other hand, manual rewriting is highly labor-intensive, which limits its scalability.
+Therefore, an effective method for assessing the quality of generated text and calibrating the alignment between the text and image is essential to ensure a high level of LMM-inherent knowledge integration.
+Compared to classical methods that assess LMM-generated descriptions by relying on manually-annotated references to measure content overlap (Zhang et al.
+2020; Lee et al.
+2020), we devise an automatic mechanism to simulate human evaluation systems.
+This mechanism places greater emphasis on semantic consistency rather than merely assessing typical content overlap, by introducing a novel quality metric called PerceptionScore.
+Concretely, it comprises two parts: text semantic quality and visual recall quality.
+The former one directly adopts the efficient implementation of CLIPScore (Hessel et al.
+2021), which is based on CLIP’s strong zero-shot capabilities and demonstrates high correlation with human semantic judgment.
+However, CLIPScore solely measures textual quality without consideration of the omission of specific visual elements in the text, thereby lacking an assessment of visual recall (Hu et al.
+2023).
+[…]
+$I^{\prime}=\operatorname{SDXL}(I),$ (1)
+$\texttt{CycleScore}(I,I^{\prime})=1-\operatorname{MAE}(\operatorname{Seg}(I),\operatorname{Seg}(I^{\prime})),$ (2)
+$\texttt{PerceptionScore}(I,T)=(\texttt{CLIPScore}(I,T)+$
+$\mathop{\texttt{CycleScore}}(I,I^{\prime}))/2.$ (3)
 Figure 4: The procedure of CycleScore calculation.
-[…]
-$\begin{split}&I^{\prime}=\operatorname{SDXL}(I),\quad\texttt{CycleScore}(I,I^{\prime})=1-\operatorname{MAE}(\operatorname{Seg}(I),\operatorname{Seg}(I^{\prime})),\\ &\texttt{PerceptionScore}(I,T)=(\texttt{CLIPScore}(I,T)+\mathop{\texttt{CycleScore}(I,I^{\prime})})/2.\end{split}$ (1)
-3.2 Multi-Granularity Cross-modal Alignment
+Multi-Granularity Cross-modal Alignment
 Modality Representation.
-Utilizing automated generation techniques, we obtain a dataset of high-quality image-text pairs, represented as $D=(I,T)$.
-Here, $I$ denotes satellite images $I^{st}$ or street-view images $I^{sv}$, while $T$ represents satellite text descriptions $T^{st}$ or street-view text descriptions $T^{sv}$.
-We first deploy Vision Transformer (ViT) [15] as an image encoder to process images derived from satellite and street-view sources.
-Finally, we use the output $\mathbf{z}_{l}^{0}$ corresponding to the special tokens ${I}_{cls}$ inserted at the beginning as the final output of the image encoder, rewrite as $\mathbf{z}_{I}$ for image representation.
-Concurrently, we use basic Transformer-Encoder [68] as a text encoder by default to comprehend and encode the semantics of textual descriptions in parallel.
-Furthermore, considering that the geospatial location information associated with street-view images can aid in pinpointing the precise geospatial coordinates of the area, we utilize the open-sourced GeoCLIP’s [69] location encoder to effectively instill geospatial information.
-This location encoder captures longitude and latitude $\mathcal{L}_{g}$, generating semantically rich, high-dimensional geospatial features, denoted as $\mathbf{z}_{L}$.
+Utilizing automated text generation methods, we obtain a dataset of high-quality image-text pairs $D=(I,T)$.
+Here, $I$ represents satellite images $I^{st}$ or street-view images $I^{sv}$, while $T$ indicates satellite text descriptions $T^{st}$ or street-view text descriptions $T^{sv}$.
+We first deploy Vision Transformer (ViT) (Dosovitskiy et al.
+2021) as the image encoder to process images derived from satellite and street-view sources.
+Then we use the output $\mathbf{z}_{l}^{0}$ corresponding to the global tokens ${I}_{cls}$ inserted at the beginning as the final output of the image encoder, rewrite as $\mathbf{z}_{I}$ for image representation.
+Concurrently, we leverage the basic Transformer-Encoder (Vaswani et al.
+2017) as the text encoder to comprehend and encode the semantics of textual descriptions in parallel.
+Furthermore, considering that the geospatial location information associated with street-view images can assist in pinpointing the precise geospatial coordinates of the area, we utilize the open-sourced GeoCLIP’s (Vivanco, Nayak, and Shah 2023) location encoder to capture longitude and latitude $\mathcal{L}_{g}$, generating semantically informative geospatial features, denoted as $\mathbf{z}_{L}$.
+Additional details can be found in Appendix G.
 Modality Alignment.
-The efficient alignment of visual and textual modalities is crucial for accurately modeling regional representations at both the coarse-grained global level and the fine-grained local level.
-To capture global visual information, in addition to the overall visual features for the region derived from satellite images, the clustering aggregation of location-level clues gathered within the region can additionally inject fine-grained details.
-Therefore, we further incorporate fine-grained information from street-view images, which are simultaneously fused with the corresponding geographical features.
-This approach aids in learning a robust and comprehensive representation of urban areas.
+To capture the comprehensive visual information, beyond deriving global visual features for the region from satellite images, the aggregation of location-level clues within the region can enhance the injection of details.
+Therefore, we further incorporate fine-grained information from street-view images, along with the corresponding geographical features.
 The formalized process is as follows:
-$\mathbf{z}_{g_{st}}=\textit{f}~{}(\mathbf{z}_{I_{st}},~{}\operatorname{Aggr}(\mathbf{z}_{I_{sv}}^{1},\mathbf{z}_{I_{sv}}^{2},...,\mathbf{z}_{I_{sv}}^{m}),~{}\operatorname{Aggr}(\mathbf{z}_{L}^{1},\mathbf{z}_{L}^{2},...,\mathbf{z}_{L}^{m})),$ (2)
-where $\mathbf{z}_{I_{st}}$ represents the satellite visual features for the queried region, $\mathbf{z}_{I_{sv}}^{k}$ denotes the $k$-th street-view visual feature in that region, and $\mathbf{z}_{L}^{k}$ represents the $k$-th street-view location feature in the same region.
-The variable $m$ indicates the count of street-view scenes within the region.
+$\mathbf{z}_{g_{st}}=\textit{f}~{}(\mathbf{z}_{I_{st}},~{}\operatorname{Aggr}(\mathbf{z}_{I_{sv}}^{1},...,\mathbf{z}_{I_{sv}}^{m}),~{}\operatorname{Aggr}(\mathbf{z}_{L}^{1},...,\mathbf{z}_{L}^{m})),$ (4)
+where $\mathbf{z}_{I_{st}}$ denotes the satellite visual features for the query region, $\mathbf{z}_{I_{sv}}^{k}$ is the $k$-th street-view visual feature in that region, and $\mathbf{z}_{L}^{k}$ represents the $k$-th street-view location feature in the same region.
+$m$ indicates the count of street-view scenes within the region.
 $\operatorname{Aggr}(\cdot)$ denotes the aggregation method.
 The function f  signifies the feature fusion method.
-To establish a global-level semantic correspondence between satellite visual data and their associated textual descriptions, we employ a joint contrastive optimization approach for the image and text modal encoders in the satellite branch.
-For the $i$-th satellite image-text pair $(I^{st}_{i},T^{st}_{i})$ in a mini-batch, we contrast the image-text pairs against others in the samples, which aims to maximally preserve the mutual information between the authentic pairs in the latent space.
-The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms: $\mathcal{L}_{{CG}}^{Image\rightarrow Text}$ and $\mathcal{L}_{{CG}}^{Text\rightarrow Image}$, which measure the similarity between the visual and textual embeddings, respectively, defined as: The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms which measure the similarity between the visual and textual embeddings, respectively, defined as:
-$\mathcal{L}_{CG}=-\frac{1}{N}(\sum_{i}^{N}{\log\frac{\exp(I_{g_{i}}^{\top}T_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(I_{g_{i}}^{\top}T_{g_{j}}/\tau)}}}+\sum_{i}^{N}{\log\frac{\exp(T_{g_{i}}^{\top}I_{g_{i}}/\tau)}{\sum_{j=0}^{N}{\exp(T_{g_{i}}^{\top}I_{g_{j}}/\tau)}}}),$ (3)
+To establish a global-level semantic correspondence between the satellite features and their associated textual features, we employ a joint contrastive optimization approach for the image and text encoders in the satellite branch.
+For the $i$-th satellite image-text pair $(I^{st}_{i},T^{st}_{i})$ in a mini-batch, we contrast the image-text pairs against others within the samples, which aims to maximally preserve the mutual information between the pairs in latent space.
+The global contrastive loss function $\mathcal{L}_{CG}$ is composed of two terms: $\mathcal{L}_{{CG}}^{Image\rightarrow Text}$ and $\mathcal{L}_{{CG}}^{Text\rightarrow Image}$, which measures the similarity between the visual and textual embeddings, respectively, defined as:
+$\mathcal{L}_{CG}=-\frac{1}{N}\Bigg{(}$ $\sum_{i}^{N}\log\frac{\exp(I_{g_{i}}^{\top}T_{g_{i}}/\tau)}{\sum_{j=0}^{N}\exp(I_{g_{i}}^{\top}T_{g_{j}}/\tau)}$ (5)
+$+\sum_{i}^{N}\log\frac{\exp(T_{g_{i}}^{\top}I_{g_{i}}/\tau)}{\sum_{j=0}^{N}\exp(T_{g_{i}}^{\top}I_{g_{j}}/\tau)}\Bigg{)},$
 where $I_{g_{i}}$ and $T_{g_{j}}$ are the normalized embedding of satellite representation $\mathbf{z}_{g_{st}}$ in the $i$-th pair and that of textual representation $\mathbf{z}_{T_{st}}$ in the $j$-th pair, respectively.
 Besides, $N$ is the batch size, and $\tau$ is the temperature for contrastive learning.
-For capturing the local-level information, given that the street-view branch is expected to capture fine-grained details to enrich region embedding at ground level, approaches like [61, 30] rely exclusively on the similarity of each modality’s global feature, thus lacking the capability to capture finer-level information such as the relationship between visual objects and textual tokens [76].
-We leverage a fine-grained interaction mechanism to implement cross-modal alignment.
+The street-view branch aims to enrich region embedding with fine-grained local-level information.
+However, previous approaches like (Radford et al.
+2021; Jia et al.
+2021) rely exclusively on global feature similarity within each modality, disregarding the necessity for fine-grained alignment, such as the correspondence between visual objects and textual tokens (Yao et al.
+2022).
+To address this challenge, we leverage a fine-grained interaction mechanism to implement cross-modal alignment at the local level.
 Specifically, we utilize token-level maximum similarity between visual and textual tokens to direct the contrastive objective.
-Initially, we compute the similarity between each visual token and all textual tokens, and then leverage the maximum value to calculate the average similarity of all image tokens to textual tokens.
-The similar approach is also applied to text-to-image.
-$\operatorname{SIM}(v_{i},t_{i})=\frac{1}{l_{1}}(\sum_{k_{1}=1}^{l_{1}}{\operatorname{argmax}_{k_{2}\in[0,l_{2})}(v_{ik_{1}}^{\top}t_{ik_{2}})}),\vspace{-0.5em}$ (4)
+We first compute the similarity between each visual token and all textual tokens, and then leverage the maximum value to calculate the average similarity of all image tokens to textual tokens.
+The similar approach is also applied to text-to-image process.
+$\operatorname{SIM}(v_{i},t_{i})=\frac{1}{l_{1}}(\sum_{k_{1}=1}^{l_{1}}{\operatorname{argmax}_{k_{2}\in[0,l_{2})}(v_{ik_{1}}^{\top}t_{ik_{2}})}),$ (6)
 $\operatorname{SIM}(t_{i},v_{i})=\frac{1}{l_{2}}(\sum_{k_{2}=1}^{l_{1}}{\operatorname{argmax}_{k_{2}\in[0,l_{1})}(t_{ik_{1}}^{\top}v_{ik_{2}})}),$ (7)
 where $v_{i}$ and $t_{j}$ denote the normalized embedding of street-view representation $\mathbf{z}_{I_{sv}}$ in the $i$-th pair and that of textual representation $\mathbf{z}_{T_{sv}}$ in the $j$-th pair, respectively.
-The fine-grained token-level representation can be optimized through:
-$\mathcal{L}_{CL}=-\frac{1}{N}(\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(v_{i}^{\top},t_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(v_{i}^{\top},t_{j})/\tau)}}}+\sum_{i}^{N}{\log\frac{\exp(\operatorname{SIM}(t_{i}^{\top},v_{i})/\tau)}{\sum_{j=0}^{N}{\exp(\operatorname{SIM}(t_{i}^{\top},v_{j})/\tau)}}}).$ (6)
-3.3 Pretraining & Fine-Tuning
-Pretraining Stage. The overall objective of UrbanVLP can be defined as the simultaneous optimization of the aforementioned two losses:
-$\mathcal{L}_{Total}=\alpha\mathcal{L}_{{CG}}+\beta\mathcal{L}_{{CL}}.$ (7)
-where $\alpha$ and $\beta$ are weight hyperparameter.
-Through backpropagation optimization, We achieve multi-granularity cross-modal urban image-text input data alignment, resulting in a robust encoder.
-Fine-Tuning Stage.
-During the fine-tuning stage, we employ a linear probing approach [23].
-This process begins by extracting frozen features $\mathbf{e}_{st},\mathbf{e}_{sv},\mathbf{e}_{p}$ from the pretrained encoder for satellite images, street-view images, and street-view positions (The latter two are optional).
-Subsequently, these features are fused, and a minimalist classifier (MLP) is trained on top to fine-tune the prediction of urban metrics, denoted as $\mathbf{Y}_{i}=MLP(\mathbf{e}_{sa},\mathbf{e}_{sv},\mathbf{e}_{p})$.
-As depicted in Figure 2, these encoders have adapted to capturing high-level abstractions from extensive visual data, enabling the MLP to manipulate static features and study the model’s inference capabilities.
-Note that in downstream tasks, text information is unnecessary as additional text generation time impedes real-time application.
-4 Experiments
-4.1 Experimental Setup
-Dataset & Task Description.
-In this paper, given the current lack of open-sourced datasets in the research community, we introduce a new benchmark dataset named $\mathtt{CityScape}$, which will be released upon paper notification.
-$\mathtt{CityScape}$ is unique in that it comprises a dual-category structure encompassing both satellite and street-view image components, each paired with corresponding high-quality textual descriptions.
-For the satellite image-text pairs, we adhere to [74] to utilize core area data of four cities in China.
-A comprehensive overview of $\mathtt{CityScape}$, along with relevant statistics, is provided in Appendix B.
-Table 1: Urban indicators prediction results in four datasets. The best results are in bold and the second-best results are underlined.
-Methods UrbanVLP StructuralUrban [41] Vision-LSTM [27] UrbanCLIP-SV [74] UrbanCLIP [74] PG-SimCLR [73] ViT [15]
+The fine-grained token-level representation can be optimized via:
+$\mathcal{L}_{CL}=-\frac{1}{N}\Bigg{(}$ $\sum_{i}^{N}\log\frac{\exp(\operatorname{SIM}(v_{i}^{\top},t_{i})/\tau)}{\sum_{j=0}^{N}\exp(\operatorname{SIM}(v_{i}^{\top},t_{j})/\tau)}$ (8)
+$+\sum_{i}^{N}\log\frac{\exp(\operatorname{SIM}(t_{i}^{\top},v_{i})/\tau)}{\sum_{j=0}^{N}\exp(\operatorname{SIM}(t_{i}^{\top},v_{j})/\tau)}\Bigg{)}.$
+Methods UrbanVLP StructuralUrban Vision-LSTM UrbanCLIP-SV UrbanCLIP PG-SimCLR ViT
 Metric $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
@@
 $1^{\text{st}}$ount 56 1 0 1 5 0 0
-Baselines for Comparison.
-We compare our method with seven recent baselines in the field of urban imagery-based socioeconomic prediction.
-The single granularity imagery-based urban region representation learning methods include: ViT [15], PG-SimCLR [73], UrbanCLIP [74], UrbanCLIP-SV [74].
-We also apply multiple baselines for multi-granularity urban imagery-based representation learning: Vision-LSTM [27] and StructuralUrban  [41].
-The in-depth overview of each baseline method is presented in the Appendix C.
+Table 1: Socioeconomic indicators prediction results in four datasets. The best results are in bold and the second-best results are underlined.
+Pretraining & Fine-Tuning
+Pretraining Stage. The overall objective of UrbanVLP can be defined as the joint optimization of the above two losses:
+$\mathcal{L}_{Total}=\alpha\mathcal{L}_{{CG}}+\beta\mathcal{L}_{{CL}}.$ (9)
+where $\alpha$ and $\beta$ are hyperparameters for a trade-off. Through backpropagation optimization, we achieve multi-granularity cross-modal alignment, resulting in robust encoders.
+Fine-Tuning Stage.
+As depicted in Figure 3 Right, during the fine-tuning stage, we employ a linear probing approach (He et al.
+2022) which begins by extracting $\mathbf{e}_{st},\mathbf{e}_{sv},\mathbf{e}_{p}$ features from the pretrained encoder for satellite images, street-view images, and street-view positions.
+Subsequently, these features are fused together, and a minimalist classifier (MLP) is trained on top to fine-tune the prediction of urban metrics, denoted as $\mathbf{Y}_{i}=MLP(\mathbf{e}_{st},\mathbf{e}_{sv},\mathbf{e}_{p})$.
+It is noteworthy that in downstream tasks, textual information is unnecessary since the knowledge embedded in the text has already been imparted to other modality encoders through pretraining, and additional text generation time hinders real-time application.
+Experiments
+Experimental Setup
+Dataset & Task Description.
+Given the current lack of open-sourced datasets in the research community, we introduce a new benchmark dataset named $\mathtt{CityView}$, which will be released upon paper notification.
+$\mathtt{CityView}$ is unique in that it comprises a dual-category structure encompassing both satellite and street-view image components, each paired with corresponding high-quality textual descriptions.
+We adhere to (Yan et al.
+2024) to cover core area data of four cities in China.
+A comprehensive overview of $\mathtt{CityView}$, along with relevant statistics, is provided in Appendix C.
+Baselines.
+Following the established practice (Yan et al.
+2024; Xi et al.
+2022; Liu et al.
+2023d) in this area, we compare our method with seven recent baselines in the field of imagery-based USI prediction.
+The single-granularity urban imagery-based methods include:
+ViT (Dosovitskiy et al. 2021), PG-SimCLR (Xi et al. 2022), UrbanCLIP (Yan et al. 2024), UrbanCLIP-SV (Yan et al. 2024).
+We also apply multiple baselines for multi-granularity urban imagery-based USI prediction: Vision-LSTM (Huang et al.
+2023) and StructuralUrban (Li et al.
+2022c).
+They are all trained using data from our $\mathtt{CityView}$ dataset.
+The in-depth overview of baselines is presented in Appendix E.
 Evaluation Metrics.
-The performance of our predictions is evaluated using a trio of standard metrics: the coefficient of determination denoted as $R^{2}$, the root mean squared error (RMSE), and the mean absolute error (MAE).
+We follow (Yan et al.
+2024) to evaluate our method in terms of: the coefficient of determination $R^{2}$, the root mean squared error (RMSE), and the mean absolute error (MAE).
 An increase in $R^{2}$, along with a decrease in RMSE and MAE values, signifies improved model accuracy.
-More implementation details are provided in Appendix D.
-4.2 Performance Evaluation
-Figure 4: $R^{2}$ results in Beijing.
-To further validate the effectiveness of the proposed UrbanVLP framework, we conduct comparisons with existing state-of-the-art methods on our proposed datasets.
-The comprehensive results are presented in Table 1, and we can draw the following conclusions: UrbanVLP significantly surpasses the performance of baseline methods employing both single-granularity and multi-granularity imagery, achieving a substantial margin of improvement.
-It can be seen that UrhanVLP outperforms the best baseline, StructuralUrban [41], by 3.3%, 2.3%, 2.5%, 4.9%, 1.0% and 2.4% in terms of $R^{2}$ for all 6 indicators: Carbon, Population, GDP, Night Light, House Price and POI in Beijing.
-The similar performance improvement can also be seen in the dataset of the other three cities.
-Besides, the average performance gain of UrbanVLP on RMSE and MAE in Beijing dataset are 1.8% and 2.1%, respectively.
-In order to provide a more intuitive display, we have created Radar Images in Figure 4 to visualize the performance of Beijing.
-The results further prove the effectiveness of our framework for urban region representation learning.
-Multi-granularity cross-modal alignment effectively incorporates fine-grained information, demonstrating superior performance.
-We can observe that pipelines leveraging multi-granularity approaches typically demonstrate superior performance compared to those employing a single granularity level.
-This enhancement can be rationalized by the incorporation of additional fine-grained information derived from street-view imagery.
-Concurrently, it is noticed that UrbanCLIP-SV underperforms compared to UrbanCLIP, demonstrating that street-view imagery, while rich in fine-grained details, lacks macroscopic perspective, leading to suboptimal outcomes when used in isolation.
-This also underscores the irreplaceability of satellite imagery as a macroscopic visual modality.
-Compared to other modalities, the textual modality facilitates a more explicit and comprehensive understanding of the region.
-In baseline models, there are works such as Vision-LSTM [27], which introduces trajectories, and PGSimCLR [73], which incorporates POI distributions to integrate external information and enhance regional representations.
-However, the superior performance of our UrbanVLP illustrates the efficacy of textual mobility as a high-information-density modality [23], which fully leverages the inherent knowledge of LLMs and explicitly enhances interpretability.
-4.3 Ablation Studies
-Figure 5: Ablation study on $\mathtt{CityScape}$-$\mathtt{Beijing}$.
-We conduct ablation studies to investigate the effectiveness of different components in UrbanVLP on Beijing dataset, including the Satellite branch (ST branch), Street-View branch (SV branch), and Location Encoding branch (LE branch).
-More ablation experiments can be found in Appendix H.
-The effectiveness of incorporating street-view branch.
-One of the main contributions of this work is the incorporation of multi-granularity information that reflects the urban spatial hierarchy.
-The street-view branch specifically represents micro-level modeling.
+Implementation Details. The implementation details are provided in Appendix F.
+Performance Evaluation
+To evaluate our UrbanVLP framework, we conduct comparisons with existing state-of-the-art methods on our proposed $\mathtt{CityView}$ datasets.
+Table 1 presents the overall results, from which we can obtain the following findings:
+1) UrbanVLP significantly outperforms the baselines which employ single/multi-granularity imagery.
+It can be seen that UrbanVLP surpasses the best baseline (StructuralUrban) by 3.3%, 2.3%, 2.5%, 4.9%, 1.0% and 2.4% in terms of $R^{2}$ for all six indicators: Carbon, Population, GDP, Night Light, House Price and POI in Beijing.
+Similarly, we can witness consistent improvements in the other three cities.
+In addition, the average reduction of UrbanVLP in terms of RMSE and MAE on the Beijing dataset are 1.8% and 2.1%, respectively.
+For a more intuitive display, we visualize the overall performance on Beijing and Shenzhen in Figure 2, where the results further prove the versatility of our framework for USI prediction.
+2) Integrating multi-granularity imagery can lead to superior performance against baselines.
+This merit can be rationalized by the incorporation of additional fine-grained information derived from street-view imagery.
+Meanwhile, UrbanCLIP-SV underperforms UrbanCLIP, revealing that street-view imagery (while rich in details) lacks a macro view, resulting in suboptimal outcomes when used in isolation.
+This also underscores the irreplaceability of satellite imagery serving as a macro visual modality.
+3) Compared to other modalities (e.g., POIs), the textual modality facilitates a more comprehensive understanding of the region.
+We include various baselines for comparison, such as PGSimCLR which incorporates POI distributions to integrate external information and enhance regional representations.
+Despite their promising results, our UrbanVLP illustrates the efficacy of integrating textual data as a information-compact modality (He et al.
+2022), which fully leverages the inherent knowledge of LLMs and explicitly enhances interpretability in the training phase.
+Ablation Studies
+As shown in Figure 5, we conduct ablation studies to examine each component in UrbanVLP on the $\mathtt{CityView}$-Beijing dataset, including the Satellite branch (ST branch), Street-View branch (SV branch), and Location Encoding branch (LE branch).
+More discussion can be found in Appendix J.
+Figure 5: Ablation study on $\mathtt{CityView}$-$\mathtt{Beijing}$.
+Effects of Street-View Branch.
+One of our contributions is the incorporation of multi-granularity information that reflects the urban spatial hierarchy.
 It can be observed that the incorporation of street-view branch results in an average improvement of 3.56% in terms of $R^{2}$.
-The presence of the street-view branch introduces the capture of rich detailed information, thereby enhancing the precision of the modeling.
-Furthermore, it is also can be seen that while street-view imagery provides fine-grained clues, its individual impact is still overshadowed by that of satellite imagery.
-This aligns with the performance differential between UrbanCLIP and UrbanCLIP-SV discussed in the Section 4.2, underscoring the indispensable role of macro-level modeling
-The effectiveness of textual modality.
-In Table 1, we demonstrate the function of the textual modality by comparing UrbanVLP with a standard ViT-based model, which shares the same configuration as the unimodal visual encoder of UrbanVLP (encompassing both satellite and street-view branches).
-We then utilize the visual features extracted without textual augmentation to predict downstream urban indicators.
+This enhancement is attributed to the street-view branch’s ability to capture rich, detailed information, thereby improving modeling precision.
+Effects of Textual Modality.
+In Table 1, we demonstrate the function of the textual modality by comparing UrbanVLP with a standard ViT-based model, which shares the same configuration as the unimodal visual encoder of UrbanVLP.
+We then utilize the visual features extracted without textual augmentation to predict downstream USI.
 As we can see, the lack of textual information leads to substantial performance degradation, underscoring the critical role of textual modalities in attaining a comprehensive visual representation.
-Similar findings were also reported in [74].
-Exploration of the impact of location coordinates.
-The role of location coordinates has been underexplored in previous work of urban region representation learning.
-However, its importance has already been emphasized in work related to Geo-Localization [69, 79, 3].
+Similar findings were also reported in (Yan et al.
+2024).
+Effects of Location Coordinates.
 In Figure 5, we evaluate the performance of the model without the Location Encoding branch.
 Although not as impactful as visual modalities, the inclusion of geospatial locations has yielded an average improvement of 1.5% in $R^{2}$.
-However, when relying solely on location information, the performance is suboptimal.
-This indicates that at the current stage, location data can only serve as an auxiliary enhancement and lacks the capacity for standalone modeling due to its deficiency in contextual semantic information.
-Figure 6: Illustration of the quality of our generated description example.
-Figure 7: Examples of predictions from UrbanCLIP and our proposed UrbanVLP.
-4.4 Qualitative Analysis
-We further investigate the quality of generated texts and the predictive performance of UrbanVLP in practice. Due to page limitations, more experimental analysis can be found in the Appendix.
+It is noteworthy that we utilize a well-pretrained location encoder and freeze it within our framework.
+Consequently, its generalizability is ensured, preventing overfitting to specific cities in $\mathtt{CityView}$ dataset.
+Qualitative Analysis
+We further investigate the quality of generated texts and the predictive performance of UrbanVLP in practice. More empirical analysis can be found in Appendix J.
 Illustration of the quality of our generated descriptions.
@@
 The generated description of the street-view image (a) is depicted in (b).
-Subsequently, we employ GPT4V [75] to assess the quality score of the generated description, which is rated as 7 out of 10, indicating the effectiveness of the generated text.
+Subsequently, we employ GPT4V (Yang et al.
+2023) to assess the quality score of the generated description, which is rated as 7 out of 10, indicating the effectiveness of the generated text.
 Furthermore, GPT4V also provides specific areas for improvement in (c).
@@
 It is evident that the generated image bears a resemblance to the original one.
-Figure 8: Our web platform. Zoom in or see Figure 17 for a better view.
+Figure 6: An example of generated text descriptions.
 Case Study for Predicted Results.
-Here we also show some predicted results in Figure 7 on $\mathtt{CityScape}$-$\mathtt{Beijing}$ dataset.
-As we can see, satellite images (a) and (b), although sharing similar layouts, differ significantly in their land use.
-(a) encompasses residential and campus areas, whereas (b) is indicative of industrial zones.
-Their socio-economic characteristics are entirely different, which could be reflected by socio-enconomic indicators.
-The carbon emissions and GDP of campus and residential area are significantly lower than those of industrial park, whereas the disparities in population and housing prices are not as pronounced, possibly due to the unique characteristics of school district housing near educational institutions.
-Night light intensity is similar across the two area due to the fact that the majority of nocturnal economic activities within factories occur indoors, unlike in Central Business Districts (CBDs) where extensive illumination is generated.
-The same principle applies to school and residential area as well.
-As observed, the predictions made by UrbanCLIP do not effectively differentiate between the two, with the downstream metrics in Figure 7(b) still inclined towards predicting a homogenized socio-economic functional area similar to that of (a).
-In contrast, the results from UrbanVLP are capable of distinctly distinguishing the socio-economic attributes of the two in terms of carbon emission, GDP, and house price.
-To showcase practical applications, we further develop a web-based system (Figure 8 and 17), with more details and discussion in Appendix I.
-Visualization of Region Representations.
+Here we also show some predicted results in Figure 7 on $\mathtt{CityView}$-$\mathtt{Beijing}$ dataset.
+Satellite images (a) and (b), though similar in layout, differ significantly in land use.
+(a) encompasses residential and campus areas, whereas (b) represents industrial zones, leading to entirely different socioeconomic characteristics.
+The carbon emissions and GDP of campus and residential area are significantly lower than those of industrial parks, whereas the disparities in population are not as pronounced, possibly due to the unique characteristics of school district housing near educational institutions.
+As observed, UrbanCLIP’s predictions fail to effectively differentiate between the two, as the downstream metrics in Figure 7(b) still lean towards predicting a homogenized socioeconomic area similar to (a).
+In contrast, the results from UrbanVLP are capable of distinctly distinguishing the socioeconomic attributes of the two in terms of carbon emission and GDP.
+Figure 7: Case study of UrbanCLIP (UC) and our UrbanVLP (UV).
+Visualization of Region Representations
 In this section, we map the region representations learned in UrbanVLP into a two-dimensional space using the PCA algorithm in Figure 8.
-This process aims to visualize how the semantic meanings of the regions are manifested within the high-dimensional representation space.
 As we can see, the three images on the left and on the right belong to different clusters, each exhibiting intra-cluster similarities and inter-cluster differences.
-Specifically, UrbanVLP effectively models regions into high-dimensional space, wherein satellite images with similar architectural layouts demonstrate spatial similarity.
-Consequently, this leads to a reasonable, accurate, and semantically rich representation of the regions.
+Therefore, UrbanVLP could effectively model regions into high-dimensional space, wherein satellite images with similar architectural layouts demonstrate spatial similarity.
+Our framework learns a reasonable, accurate, and semantically rich region representation for USI prediction.
 Figure 8: Representation space visualization.
-Figure 10: PCA visualization of urban imagery.
-Figure 10 provides a further visualization with a population indicator, compared to UrbanCLIP.
-We use the dot color to indicate the population (log) in the corresponding regions of Beijing.
-As shown by the red dashed box, the regions modeled by UrbanVLP exhibit more significant clustering characteristics and greater distinction in high-dimensional space.
-Additionally, the lower parts of the two main band-shaped branches in the figure are also more distinctly separated.
-Points that are spatially close also have similar population values.
-5 Conclusion and Future Work
-Urban region representation learning plays a significant role in helping us understand the patterns, structures, and dynamics of cities.
+Figure 9: Transferability test on $R^{2}$ between UrbanVLP and UrbanCLIP, on the Carbon indicator across 4 cities.
+Transferability Study
+In this study, we explore the practical application of USI prediction within a transfer learning context (Park et al.
+2022; Liu et al.
+2023d).
+Specifically, We use visual encoders trained on source city data and fine-tune them on target city data.
+We experiment with different pairs of source and target cities and present the carbon emission prediction in Figure 9.
+As we can see, on 16 source-target city pairs, UrbanVLP achieves an average $R^{2}$ of 0.574, while that of UrbanCLIP is 0.495.
+Additionally, the $R^{2}$ values show greater similarity between Beijing and Shanghai, as well as between Shenzhen and Guangzhou.
+This observation aligns with the geographical proximity of Beijing to Shanghai and Guangzhou to Shenzhen.
+These findings confirm the robust transferability of our UrbanVLP model in urban areas, despite the previously mentioned differences in geological and demographic characteristics among the selected cities.
+Analysis about Hyperparameters
+In Figure 10 (a), we explore different combinations of hyperparameter values for loss combination.
+It can be seen that the settings we choose in the paper ($\alpha$=0.5, $\beta$=0.5) yield the best experimental results.
+We also study the maximum value of street-view images within one region during training, which is set as 25 in our paper.
+As is shown in Figure 10 (b), the $R^{2}$ value steadily increases with the number of street-view images, peaking at 25.
+Therefore, our setting is the optimal.
+Figure 10: (a) Hyperparameter for loss combination. (b) Experiments for the maximum number of street-view images.
+Practicality
+To showcase the practical applications, we further develop a web-based system (Figure 19 in Appendix), with more details and discussion in Appendix K.
+This system renders a visual representation of urban areas through satellite imagery, complemented by street-level photographs for an interactive exploration experience.
+It enhances social planning and sustainable development by providing a data-driven platform that deepens understanding of socioeconomic indicators.
+Conclusion and Future Work
+USI prediction plays a significant role for understanding societal patterns and dynamics.
 UrbanVLP, for the first time, explores the differences between street-view and satellite images from a semantic granularity perspective, as well as their roles in modeling urban region representation.
 A text generation and calibration mechanism is also proposed to ensure high-quality description generation.
-It has achieved state-of-the-art results on the constructed $\mathtt{CityScape}$ and downstream dataset.
-In the future, the further research direction can be two-fold: 1) Adding additional modalities such as POIs to enhance information richness, and 2) enabling the design of better model architectures such as MAE [23] to efficiently utilize existing data source.
-Appendix / supplemental material
-Appendix A Formulation
-Definition 1 (Urban Region) We follow previous research [73, 74] and evenly divide the target area (e.g. a city) into $L$ urban regions.
+It has achieved state-of-the-art results on the constructed $\mathtt{CityView}$ dataset.
+Future research could incorporate additional modalities such as POIs, road networks, and building footprints to enhance information richness and introduce a broader perspective.
+Acknowledgments
+This work is mainly supported by the National Natural Science Foundation of China (No.
+62402414).
+This work is also supported by the Guangzhou-HKUST(GZ) Joint Funding Program (No.
+2024A03J0620), Guangzhou Municipal Science and Technology Project (No.
+2023A03J0011), the Guangzhou Industrial Information and Intelligent Key Laboratory Project (No.
+2024A03J0628), and a grant from State Key Laboratory of Resources and Environmental Information System, and Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No.
+2023B1212010007).
+A    Formulation
+Definition 1 (Urban Region) We follow previous research (Xi et al. 2022; Yan et al. 2024) to evenly divide the target area (e.g. a city) into $L$ urban regions.
 Definition 2 (Urban Imagery) Diverse visual images, including satellite and street-view images, facilitate an intuitive analysis of urban regions.
 Specifically, satellite images provide a coarse-grained overview of the structural composition of a region.
-Each input satellite image w.r.t the urban area $g$ can be represented as $I^{sa}_{g}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the length and width.
+Each input satellite image w.r.t the urban area $g$ can be represented as $I^{st}_{g}\in\mathbb{R}^{H\times W\times 3}$, where $H$ and $W$ denote the length and width.
 Street-view images are captured by vehicles from map providers during urban traversals, offering detailed visual information on the fine-grained aspects within a region.
@@
 Each individual street-view image is represented as $I^{{sv}}_{g}\in\mathbb{R}^{H\times W\times 3}$.
-Definition 3 (Text Description) Textual descriptions of urban area $g$ (including satellite text $T^{sa}_{g}$ and street-view text $T^{sv}_{g}$) provide precise means for urban area analysis.
-Such text can be manually generated or produced using image annotation tools.
+Definition 3 (Text Description) Text descriptions of urban area $g$ (including satellite text $T^{st}_{g}$ and street-view text $T^{sv}_{g}$) provide precise tools for urban area analysis.
+Such texts can be manually generated or produced using image annotation.
 Particularly, leveraging advanced LLMs enables detailed and insightful analytical descriptions of given area images.
-Notably, incorporating spatial context (e.g.
-vegetation coverage and POI) significantly reflects the functional characteristics of the region.
-Definition 4 (Urban Indicators) Urban indicators assess various aspects of a region, including population, economy, society, and public services, from multiple perspectives.
+Definition 4 (Urban Socioeconomic Indicators) Urban socioeconomic indicators assess various aspects of a region, including population, economy, society, and public services, and so on.
 A set of $L$ urban regions with $K$ indicators is represented as $\mathbf{Y}\in\mathbb{R}^{L\times K}$.
-In this paper, we comprehensively use population (count), GDP (in million Chinese Yuan), carbon emissions (in tons), night lights (intensity), house prices (in Chinese Yuan), and points of interest, i.ePOI, (count for different categories) as ground-truth urban indicators.
-Appendix B More Details about Datasets & Tasks
-B.1 CityScape Dataset
-We construct the $\mathtt{CityScape}$ Dataset using street-view geo-tagged images collected from Baidu Map API and generated text descriptions by LMM.
+In this paper, we comprehensively use population, GDP, carbon emissions, night lights, house prices , and points of interest, i.ePOI, as ground-truth urban indicators.
+B    Detailed Related Work
+Urban Socioeconomic Indicator Prediction.
+In literature, many studies have focused on learning task-specific region representations from various urban data, especially urban imagery due to its consistent updates and easy accessibility (Li et al.
+2022d; Liu et al.
+2023d; Xi et al.
+2022).
+For example, Urban2Vec (Wang, Li, and Rajagopal 2020) integrated street-view imagery and POI data to learn neighborhood embeddings.
+Some contrastive learning approaches like PG-SimCLR (Xi et al.
+2022) and UrbanCLIP (Yan et al.
+2024) have shown success in representing urban regions through satellite images.
+However, the aforementioned works exclusively considered a single type of urban imagery, overlooking the potential synergy between satellite imagery and street-view imagery, which can complement each other.
+Recently, some efforts (Barbierato et al.
+2020; Huang et al.
+2023; Li et al.
+2022d; Cao et al.
+2023; Law, Paige, and Russell 2019; Cao et al.
+2018; Huang et al.
+2023) have begun to explore the simple combination of satellite imagery and street-view imagery.
+Vision-LSTM (Huang et al.
+2023) integrates diverse street-level images into the satellite image domain using an LSTM-based network for precise urban village identification.
+Li et al.
+(2022d) propose utilizing street segments for region modeling and predicting socioeconomic indicators by merging street-view and satellite images through a multi-stage pipeline.
+Nevertheless, prior studies mostly overlook distinct features in satellite and street-view images for varied granularities.
+To this end, we propose a unified multi-granularity cross-modal alignment framework for comprehensive urban representation in this study.
+Vision-Language Pretraining (VLP).
+VLP aims to jointly encode vision and language in a fusion model.
+Early works on VLP can be broadly categorized into single-stream (Chen et al.
+2020; Kim, Son, and Kim 2021; Li et al.
+2020) and two-stream (Tan and Bansal 2019; Lu et al.
+2019) methodologies, with the former using a unified architecture for modality integration and the latter employing separate encoders before merging modalities.
+Milestone work CLIP (Radford et al.
+2021) and its variants (Li et al.
+2021; Yao et al.
+2022) highlight the efficacy of contrastive learning in cross-modal downstream tasks, such as zero-shot learning and cross-modal retrieval.
+Recent works (Tsimpoukelli et al.
+2021; Alayrac et al.
+2022) shift towards leveraging LLMs knowledge for vision-language representation learning.
+In urban region profiling, the potential of VLP paradigm remains untapped, with limited exploration of the benefits of textual information.
+Our groundbreaking work introduces multiple types of text descriptions as supplements, aiming to learn more comprehensive and interpretable urban region representations.
+Large Multimodal Model (LMM).
+As LLMs undergo rapid evolution, a faction within the research community is increasingly focused on integrating visual knowledge into these models, such as CLIP (Radford et al.
+2021) and ALIGN (Jia et al.
+2021).
+Early studies (Li et al.
+2022a, 2023) enhanced CLIP with refined data strategies for more diverse datasets, which have proven effective for fundamental visual tasks (Li et al.
+2022b; Liu et al.
+2023c; Zhang et al.
+2022) but have shown limitations for complex tasks such as visual question answering.
+Recently, new models like MiniGPT-4 (Chen et al.
+2023a), LLaVA (Liu et al.
+2023b), and InstructBLIP (Dai et al.
+2023) have improved the understanding of complex problems by expanding parameters and training data, called large multimodal models (LMMs).
+In this paper, we comprehensively evaluate the text generation quality of representative LMMs model and propose an automated text description generation and calibration technique to ensure effective image description capabilities with LMMs.
+C    More Details about Datasets & Tasks
+CityView Dataset
+We construct the $\mathtt{CityView}$ Dataset using street-view geo-tagged images collected from Baidu Map API following previous works (Liu et al.
+2023d) and generated text descriptions by LMM.
+We collect and generate satellite data following the practice in  (Yan et al.
+2024).
   • Street-View Panoramic Image Collection.
-The Street-View image sampling process utilized the Baidu Maps API [^1] https://lbsyun.baidu.com/, facilitating the extraction of road network data within our predefined sampling area.
-Using the ArcGIS [^2] https://www.esri.com/en-us/arcgis/products/develop-with-arcgis/overview tool, we exported the road network data and established a sampling interval of 500 meters, which allowed for the comprehensive collection of street-view images at every half-kilometer point along the road network, ensuring thorough coverage that complements the satellite imagery’s scope.
-Following the road network sampling, panoramic street-view images were collected based on the sampling points.
-These images, with a resolution of 2048x664, also far exceed the 256x256 resolution of satellite imagery [74], ensuring clarity and ample details for further analysis.
-    As is shown in Table 2, a total of 47,582 street-level images were collected, whose data volume is 2.7 times larger than that of the satellite dataset, capturing the intricacies of the urban landscape and providing a street-level view that enhances the overall urban analysis.
-    In figure 11, we showcase our sampling process for the road network, which is used to collect street-view imagery at strategically selected locations.
-We initially acquire road network data from the online open-source website and subsequently import it into ArcGIS with a sampling interval of 500 meters.
-Following this, we gather the latitude and longitude coordinates of the sampling points within our dataset area.
-These coordinates are then used to invoke the Baidu API to collect street-view imagery.
-Additionally, not all sampled points have corresponding street view images, thereby resulting in a final count of street view images that is less than the number of sampled points.
+The Street-View image sampling process utilized the Baidu Map API [^1] https://lbsyun.baidu.com/, after the extraction of road network data within our predefined sampling area.
+Using the ArcGIS [^2] https://www.esri.com/en-us/arcgis/products/develop-with-arcgis/overview tool, we first exported the road network data.
+We then set a sampling interval of 500 meters, allowing for the collection of street-view images at every half-kilometer point along the road network.
+This approach ensured thorough coverage of the corresponding area, as demonstrated in Figure 11.
+Based on the sampling points, we collected panoramic street-view images at a resolution of 2048x664, which significantly exceeds the 256x256 resolution of satellite imagery (Yan et al.
+2024).
+This higher resolution ensures greater clarity and provides ample detail for further analysis.
+    As is shown in Table 2, a total of 47,582 street-view images were collected, whose data volume is 2.7 times larger than that of the satellite data, capturing the intricacies of the urban landscape and providing a street-level view that enhances the overall urban analysis.
   • Corresponding Text Description Generation.
-The textual data corresponding to each street-view image is generated from ShareGPT4V [11], which is recognized for its superior text generation capabilities compared to other contemporary open-source LMM models [48, 14, 77, 20, 10, 11], offering comprehensive and high-quality textual information.
+The textual data corresponding to each street-view image is generated from ShareGPT4V (Chen et al.
+2023b), which is recognized for its superior text generation capabilities compared to other contemporary open-source LMM models (Liu et al.
+2023a; Dai et al.
+2023; Ye et al.
+2023; Gao et al.
+2023; Chen et al.
+2023a, b).
 The quality of the generated data is guaranteed by our proposed PerceptionScore.
-We filter out the texts with PerceptionScore lower than 0.5.
+We filter out the texts with PerceptionScore lower than 0.6.
+Figure 11: Road Network Sampling for Street-View Images on Beijing.
 Table 2: Urban imagery dataset statistics.
@@
 Shenzhen 22.45°N, 113.75°E 22.84°N, 114.62°E 10,167 4,324
-Figure 11: Road Network Sampling for Street-View Images on Beijing.
-B.2 Downstream Dataset & Tasks
+Downstream Dataset & Tasks
 In our research, the Downstream Dataset plays a crucial role in gauging the practical implications of our multimodal framework.
-We collect six representative urban indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI.
+We collect six representative urban socioeconomic indicators: Carbon Emission, Population, GDP, Night Light, House Price, and POI.
 Each indicator illuminates a significant dimension of urban vitality and sustainability.
@@
   • Carbon Emissions: Extracted from the Open-source Data Inventory for Anthropogenic CO2 (ODIAC), this dataset offers granular carbon emission figures in 2022 that are in alignment with the spatial resolution of our satellite imagery, which is one square kilometer per image.
-The emissions are quantified in tons.
+The emissions are quantified in tons on a monthly basis.
   • Population: The quantum of population serves as an indicator of a region’s sociodemographic characteristics.
@@
 The unit is #citizens.
-  • GDP: This dataset encapsulates China’s economic development by incorporating Gross Domestic Product (GDP) figures, thereby infusing an economic dimension into our analysis of urban regions.
+  • GDP: Our dataset encapsulates China’s economic development by incorporating Gross Domestic Product (GDP) figures, thereby infusing an economic dimension into our analysis of urban regions.
 The unit is million Chinese Yuan.
   • Night Light: Nighttime light can, to some extent, characterize the intensity of human activities, which is of great significance for urban development studies.
-we take the nightlight data from [87] in 2020.
-  • House Price: The price of housing is is subject to the influence of numerous socio-economic factors.
+We take the nightlight data from (Zhong, Yan, and Li 2022) in 2020.
+  • House Price: The price of housing is subject to the influence of numerous socioeconomic factors.
 By observing fluctuations in housing prices, it is possible to discern, to a certain extent, the state of economic development.
 We collect data from Anjuke (https://www.anjuke.com) in March 2023 with yuan/$m^{2}$ unit.
-Individual house price represents a locational attribute.
-To extrapolate a regional-level characteristic, we take the average of housing prices across an urban region.
+Since individual house prices represent locational attributes, we extrapolate a regional-level characteristic by averaging housing prices across an urban region.
   • POI: The spatial distribution of various types of POIs (Points of Interest) within urban areas is a potent social indicator, as it reflects the underlying social and economic architecture, as well as the population’s activities and urban lifestyles.
 Consequently, to glean insights into these urban dynamics, we collect data of POIs from web sources, encompassing both their locations and categories.
-The collected data has 14 categories: Dining and Cuisine, Leisure and Entertainment, Sports and Fitness, Business and Residential, Healthcare, Financial Institutions, Tourist Attractions, Lifestyle Services, Shopping and Consumption, Automobile Related, Hotel Accommodation, Transport Facilities, Science, Education and Culture, Companies and Enterprises.
+The collected data has 14 categories: Dining and Cuisine; Leisure and Entertainment; Sports and Fitness; Business and Residential; Healthcare; Financial Institutions; Tourist Attractions; Lifestyle Services; Shopping and Consumption; Automobile Related; Hotel Accommodation; Transport Facilities; Science, Education and Culture; Companies and Enterprises.
 To capture the unique attributes of each area, we aggregate the number of POIs by category within each region.
-We construct this benchmark with the aim of conducting a thorough evaluation of the urban region representations that have been derived through our learning framework.
-We uniformly partition the downstream dataset into train:val:test=7:1:2.
-These indicators constitute the downstream task dataset, covering economic (GDP, Night Light, House Price), social (Population, POI), and environmental (Carbon Emission) indicators.
-The data preprocessing setting adheres to [74] with refinement.
+In accordance with established conventions (Liu et al.
+2023d; Yan et al.
+2024; Xi et al.
+2022), we uniformly and randomly partition the downstream dataset into train:val:test=7:1:2.
+The data preprocessing setting adheres to (Yan et al.
+2024) with refinement.
 For some indicators data with relatively small differences in values, we have taken the logarithm of these values to facilitate training.
-B.3 Downstream Dataset Analysis
+Downstream Dataset Analysis
+To provide readers with an intuitive understanding, we visualize the distribution of three downstream socioeconomic indicators for Beijing in Figure 12.
+The distribution of different indicators exhibits similar trends, indicating that Beijing has a clear “center-periphery” structure, with the core area being densely populated and economically developed, while the peripheral areas have relatively sparse economic and demographic data.
 Figure 12: Spatial distribution of Popultion, GDP and NightLight in Beijing.
-Here we visualize the distribution of some downstream socio-economic indicator data on Beijing in Figure 12.
-The distribution of different indicators exhibits similar trends, indicating that Beijing has a clear ’center-periphery’ structure, with the core area being densely populated and economically developed, while the peripheral areas are relatively sparse.
 Beyond geographical spatial differences, the concentration trends and variability of the data are critical perspectives that merit attention.
-In Figure 14 we demonstrate these patterns by illustrating the normalized values of different socio-economic indicators in Beijing.
-Outliers with extreme maximum and minimum values will be filtered out during our processing.
+In Figure 13 we demonstrate these patterns by illustrating the normalized values of different socioeconomic indicators in Beijing.
+Outliers with extreme maximum and minimum values will be filtered out during our pre-processing.
 As we can see, the values of Carbon Emission and Night Light are relatively concentrated, showing minimal variability.
@@
 Figure 13: Data distribution of Carbon Emission, Popultion, GDP and Night Light in Beijing.
-Figure 14: PerceptionScore distribution on $\mathtt{CityScape}$ dataset for 4 representative cities.
-Appendix C In-depth Overview of Baseline Methods
-The single granularity imagery-based urban region representation learning methods include:
-  • ViT [15]: ViT is a successful paradigm of applying transformers to computer vision, where images are segmented into patches for fixed-length modeling.
-When sufficient data is available for pre-training, ViT’s performance can surpass that of CNNs [22].
+D    Motivation Demonstration for Multi-granularity Modeling
+Here, we provide a more detailed example to illustrate our motivation for urban multi-granularity modeling.
+Urban environments exhibit a spatial hierarchy in reality, from a macro region level to a micro location level (e.g., architectural details, street furniture, storefront).
+As depicted in Figure 14, while the satellite images (a) and (b) possess remarkably similar layouts, their urban indicators considerably differ.
+Upon zooming into micro levels using the corresponding street-view images (c) and (d), a more nuanced and fine-grained understanding becomes apparent.
+We can observe that their differences arise from the varying surrounding environments (i.e., educational buildings vs.
+business districts).
+Exclusively capturing macro-level patterns through satellite imagery may introduce substantial bias, falling short in furnishing more comprehensive information across diverse urban levels.
+Figure 14: Single-granularity vs. Multi-granularity modeling.
+E    In-depth Overview of Baseline Methods
+The single granularity imagery-based USI prediction methods include:
+  • ViT (Dosovitskiy et al.
+2021): ViT is a successful paradigm of applying transformers to computer vision, where images are segmented into patches for fixed-length modeling.
+When sufficient data is available for pre-training, ViT’s performance can surpass that of CNNs (He et al.
+2016).
 In our experiment, we leverage ViT-B as the baseline for comparison.
-  • PG-SimCLR [73]: A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning.
+  • PG-SimCLR (Xi et al.
+2022): A contrastive learning framework which introduces geographic information (i.e., POI) into urban region representation learning.
 In our study, we do not employ PG-SimCLR to predict POI to ensure a fair comparison.
-  • UrbanCLIP [74]: A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
-  • UrbanCLIP-SV [74]: Original UrbanCLIP is pretrained on satellite image-text pairs.
-Here we also verify its performance on our proposed $\mathtt{CityScape}$ dataset, utilizing the same aggregation method to acquire regional features.
-We also apply multiple baselines for multi-granularity urban imagery-based representation learning:
-  • Vision-LSTM [27]: A vision long short-term memory neural network which could process variable-length street-view image within each region.
+  • UrbanCLIP (Yan et al.
+2024): A model that utilizes an image-text contrastive learning-based approach for learning robust representations from urban imagery, intending to capture the complexity and diversity of urban areas.
+  • UrbanCLIP-SV (Yan et al.
+2024): Original UrbanCLIP is pretrained on satellite image-text pairs.
+Here we also verify its performance on street-view image-text pairs of our proposed $\mathtt{CityView}$ dataset, utilizing the same aggregation method to acquire regional features.
+We also apply multiple baselines for multi-granularity USI prediction:
+  • Vision-LSTM (Huang et al.
+2023): A vision long short-term memory neural network which could process variable-length street-view image within each region.
 The integration of satellite and street-view imagery, along with trajectories, facilitates a more effective capture of comprehensive urban representations, yielding reliable recognition results.
-  • StructuralUrban[41]: The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images.
+In our experiments, we include only satellite and street-view data, without incorporating trajectory data.
+  • StructuralUrban (Li et al.
+2022d): The graph-based framework for urban region profiling leveraging street segments as containers to adaptively fuse the features of multi-level urban images.
 We term this method StructuralUrban for simplicity.
-We present a visual representation of the performance in the main paper Figure 4 to effectively demonstrate the comprehensive capabilities of our model across various downstream tasks.
-There is also a trend that models tend to perform better on environmental (Carbon Emission) and social (Population, POI) indicators than on economic (GDP, Night Light, House Price) indicators.
-This is possibly because environmental and social indicators usually have more distinct geographic distribution characteristics.
-For instance, neighboring regions may have similar climates, vegetation cover, and population densities, which are important factors for environmental and social indicators and can be effectively captured through spatial embedding.
-On the other hand, economic indicators are more complex and influenced by various factors, including policies, economic activities, and market sentiments, making them harder to model.
-Additionally, due to the multi-class nature of POI, models may require more complex structures to understand the underlying relationships between labels, thus resulting in lower performance.
-Appendix D Implementation Details
-In our experiment, Adam optimizer is chosen to minimize the training loss during parameter learning.
+F    Implementation Details
+In our experiment, Adam optimizer (Kingma and Ba 2015) is chosen to minimize the training loss during the parameter learning.
 A grid search on hyperparameters is conducted, where search ranges for learning rate and batch size are set as 2e-6, 2e-5, 2e-4, 2e-3, 2e-2 and 4, 8, 16, 32, 64, respectively.
-The data augmentation strategy follows the setting of [61].
+The data augmentation strategy follows the setting of (Radford et al.
+2021).
 The weight hyperparameter $\alpha$ and $\beta$ of $\mathcal{L}_{Total}$ are set to 0.5, 0.5 respectively.
-We run all the models on NVIDIA A800 GPUs with PyTorch.
+We run all the models on NVIDIA A800 GPUs with PyTorch 1.13.1 on Ubuntu 22.04.
+The number of epochs for the pretraining and fine-tuning stages is set to 10 and 100, respectively, with an early stopping strategy.
 We implement the aggregation method $\operatorname{Aggr}(\cdot)$ through MLP layers.
-The feature fusion method f  directly utilizes additive aggregation.
+The feature fusion method f  directly utilizes additive aggregation (the experiments regarding the choice of fusion method are provided in Appendix J).
 To ensure a fixed length for network processing, we set the maximum street-view sequence length within one region as 25, and sequences shorter than that are achieved by padding with zeros.
-For the semantic segmentation model to decouple visual elements with street-view images, following  [17], the architecture for the semantic model is ResNet18dilated + PPM_deepsup.
+For the semantic segmentation model to decouple visual elements with street-view images, following the successful precedent set by  (Fan et al.
+2023b), the architecture for the semantic model is ResNet18dilated + PPM_deepsup.
 Initially, the dataset comprises 150 segmentation categories.
-However, in line with our requirements, we narrow down the selection to 38 categories, which we further group into 13 street-view features following [17]: Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
-Appendix E More Details for Modality Representation
-For utilizing ViT as the visual encoder to process urban visual data, in particular, we reshape the input image $I\in\mathbb{R}^{H\times W\times 3}$ into a sequence of flattened 2D patches ${I}_{P}\in\mathbb{R}^{N\times({P^{2}}\cdot{C})}$, where C represents the number of channels, $(P,P)$ specifies the resolution of each image patch, and $N=HW/P^{2}$ denotes the number of patches, corresponding to the input sequence length for ViT.
+However, in line with our requirements, we narrow down the selection to 38 categories, which we further group into 13 street-view features following (Fan et al.
+2023b): Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
+G    More Details for Modality Representation
+For utilizing ViT as the visual encoder to process urban visual data, we reshape the input image $I\in\mathbb{R}^{H\times W\times 3}$ into a sequence of flattened 2D patches ${I}_{P}\in\mathbb{R}^{N\times({P^{2}}\cdot{C})}$, where C represents the number of channels, $(P,P)$ specifies the resolution of each image patch, and $N=HW/P^{2}$ denotes the number of patches, corresponding to the input sequence length for ViT.
 Notebly, the ViT utilizes a consistent latent vector size (denoted as $d$) across all its layers.
 Thus, we can obtain the outputs as the patch embeddings through a trainable linear projection.
-We also prepend a learnable embedding to the patch embedding sequence (i.e., $\mathbf{z}_{0}^{0}={I}_{cls}$).
+We also prepend a learnable embedding to the patch embedding sequence to serve as the overall representation (i.e., $\mathbf{z}_{0}^{0}={I}_{cls}$).
 Besides, we add learnable position embedding $\mathbf{E}_{pos}$ to retain positional information.
@@
 $\mathbf{z}_{0}=[{I}_{cls};{I}_{p}^{1}\mathbf{W};{I}_{p}^{2}\mathbf{W};\cdots;{I}_{p}^{N}\mathbf{W}]+\mathbf{E}_{pos},$ (10)
-where $\mathbf{W}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$ and $\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times d}$ are learnable parameters.
-$\mathbf{z}_{0}$ represents the result of the linear projection, which is subsequently fed into the ViT.
-Consistent with the standard Transformer encoder, ViT includes alternating layers of multi-head self-attention $\operatorname{MSA(\cdot)}$ operation [68].
-Furthermore, layer normalization (LN) is adopted before every block, whereas residual connections are after every block:
+where $\mathbf{W}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$
+Consistent with the standard Transformer encoder, ViT includes alternating layers of multi-head self-attention $\operatorname{MSA(\cdot)}$ operation (Vaswani et al.
+2017).
+Furthermore, layer normalization (LN) is adopted before every block, whereas residual connections are applied after every block:
 $\mathbf{z}_{l}=\mathrm{MSA}(\mathrm{LN}(\mathbf{z}_{l-1}))+\mathbf{z}_{l-1},\quad l=1\ldots L.$ (11)
@@
 The input text sequence is bracketed with $[SOS]$ and $[EOS]$ tokens, and the activation of the highest layer of Transformer at $[EOS]$ token is considered the global representation $\mathbf{z}_{T}$ of text.
-Appendix F Segmentation Ratio and Prompt Examples
-As illustrated in Figure 15(a) and (b), we demonstrate examples of the segmentation ratio of street-view images in our $\mathtt{CityScape}$ dataset.
+H    Segmentation Ratio and Prompt Examples
+As illustrated in Figure 15 (a) and (b), we demonstrate examples of the segmentation ratio of street-view images in our $\mathtt{CityView}$ dataset.
 The segmentation ratio is decoupled by pretrained segmentation models and is used in both prompt construction and PerceptionScore computation.
-We group the segmentation categories into 13 categories following [17]: Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
-The prompt template used to generate street-view descriptions is formulated as follows: Analyze the street-view panoramic image in [City] in a comprehensive and detailed manner.
+We group the segmentation categories into 13 categories following (Fan et al.
+2023b): Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
+The prompt template for generating street-view descriptions is formulated as follows: Analyze the street-view panoramic image in [City] in a comprehensive and detailed manner.
 The coordinate of the street-view image is [Longitude, Latitude].
 The segmentation ratio of the street-view image is [Segment Information].
-Figure 15(c) provides illustrations of some examples of our prompts used for description generation.
+Figure 15 (c) provides illustrations of some examples of our prompts for description generation.
 Figure 15: Prompt Examples.
-Appendix G The Choice of Image-to-Text LMM
-In our work, we leverage ShareGPT4V [11] for text description generation.
+I    The Choice of Image-to-Text LMM
+In our work, we leverage ShareGPT4V (Chen et al.
+2023b) for text description generation.
 Here we conduct experiments on current open-sourced LMM models, and compute the PerceptionScore of them.
-The results are presented in Table LABEL:table:lmms.
-As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which is used as the Image-to-Text models in [74].
-Additionally, we note that there is study [86] argue that InstructBLIP [14] exhibits state-of-the-art (SOTA) performance in image-to-text generation.
+The results are presented in Table 3.
+As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which is used as the Image-to-Text models in (Yan et al.
+2024).
+Additionally, we note that there is study (Zhong et al.
+2024) argue that InstructBLIP (Dai et al.
+2023) exhibits state-of-the-art (SOTA) performance in image-to-text generation.
 However, in our experiment, we found that InstructBLIP is unsuitable for handling geospatial coordinates as prompt content.
@@
 Due to these limitations, we have excluded it from our comparative analysis.
-Table 3: Perception score of representative LMMs.
-Models Average Generated-Description Scores
-LLaVA-v1.5 [48] 0.674
-mPLUG-Owl [77] 0.662
-LLaMA-Adapter V2 [20] 0.698
-MiniGPTv2 [10] 0.693
-ShareGPT4V [11] 0.714
-Appendix H More Experimental Analysis
-H.1 More Analysis for Text Calibration
-In our paper, we selected the state-of-the-art LMM ShareGPT4V, which demonstrates performance comparable to GPT4V in image captioning.
-In order to ascertain the efficacy of ShareGPT4V, we computed the distribution of PerceptionScore for descriptions generated by ShareGPT4V, as depicted in Figure 14.
-Notably, only less than 0.02% of generated texts scored below 0.5, thereby evidencing the high quality of the descriptions produced.
-Table 4: UrbanVLP performance utilizing mPLUG-Owl with and without Text Calibration. mPO is an abbreviation for mPLUG-Owl, and TC stands for Text Calibration.
+Table 3: PerceptionScore of representative LMMs. GD is short for Generated-Description.
+Models Average GD-Scores
+LLaVA-v1.5 (Liu et al. 2023a) 0.674
+mPLUG-Owl (Ye et al. 2023) 0.662
+LLaMA-Adapter V2 (Gao et al. 2023) 0.698
+MiniGPTv2 (Chen et al. 2023a) 0.693
+ShareGPT4V (Chen et al. 2023b) 0.714
+J    More Experimental Analysis
+More Analysis for Text Calibration
+In our paper, we select the state-of-the-art LMM ShareGPT4V, which demonstrates performance comparable to GPT4V in image captioning.
+Here, we compute the distribution of PerceptionScore for descriptions generated by ShareGPT4V in Figure 16.
+Moreover, in our proposed dataset, we filter out the texts with PerceptionScore lower than 0.6.
+To validate the effectiveness of this Text Calibration process, we report the performance of using raw generated descriptions (i.e., UrbanVLP w/o Text Calibration) for comparison.
+Table 4 demonstrates that UrbanVLP consistently surpasses this variant across various indicators.
+This outcome confirms the effectiveness of our Text Calibration method, suggesting that aligning more relevant and noise-free textual information with image features enhances coherence and meaning in visual representation.
+Figure 16: PerceptionScore distribution on $\mathtt{CityView}$ dataset for 4 representative cities.
+Table 4: UrbanVLP performance with and without Text Calibration. TC stands for Text Calibration.
 Model Carbon Population GDP
 $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
-UrbanVLP_mPO_wo TC 0.687 0.551 0.544 0.643 0.580 0.458 0.494 0.655 0.447
-UrbanVLP_mPO 0.714 0.538 0.473 0.683 0.557 0.430 0.506 0.640 0.435
+UrbanVLP w/o TC 0.724 0.542 0.476 0.686 0.559 0.433 0.501 0.697 0.436
 UrbanVLP 0.769 0.477 0.369 0.714 0.523 0.411 0.537 0.684 0.416
-As a result, the experimental results comparing with and without calibration do not adequately reflect the effectiveness of our calibration method.
-To further demonstrate the effectiveness of our method, we select mPLUG-Owl [77], one of the representative LMMs compared in our paper, with the worst performance.
-After calibration, 15.3% of the text with perceptionscores below 0.5 was filtered out.
-The results validate the effectiveness of our calibration method, and indicate that more relevant and noise-free textual information could align better with image features, leading to a more coherent and meaningful visual representation.
-H.2 Analysis about Time Complexity
-Modeling and research in urban areas can provide valuable references and assistance to policymakers, builders, and citizens.
-Therefore, practical implementation should be considered.
+Analysis about Time Complexity
+Researches in urban areas can provide valuable references and assistance to policymakers, builders, and citizens. Therefore, the practicality of the model should be taken into consideration.
 We conduct time complexity analysis using FPS (Frames Per Second), a common indicator for inference speed in computer vision area.
-We report FPS for satellite image in Figure 16(b).
-In fact, introducing street view features for better and finer-grained effects will inevitably sacrifice a portion of the time efficiency.
-However, compared to [41]’s graph structure inference, our linear probing setting ensures the minimum time overhead, also with better performance.
+The result is reported in Figure 17 (a).
+In fact, compared to UrbanCLIP, UrbanVLP’s introduction of street-view features for improved and finer-grained effects will inevitably sacrifice a portion of the time efficiency.
+However, compared to StructuralUrban’s graph structure inference, our linear probing setting ensures the minimum time overhead, also with better performance.
 For making model decisions during deployment, to pursue the trade-off between performance and speed, UrbanVLP is the optimal choice.
-Figure 16: (a) Hyperparameter for loss combination. (b) FPS comparison of different mdoels. (c) Expetiments for maximum of the number of street-view images.
-H.3 Analysis about Hyperparameters
-The optimal hyperparameters for Equation 7 are actually determined by grid search.
-In Figure 16(a), we present some results of different combinations of hyperparameter values .
-It can be seen that the settings we chose in the paper ($\alpha$=0.5, $\beta$=0.5) yield the best experimental results.
-We also study the maximum value of street-view images during training, which is set as 25 in our paper.
-As is shown in Figure 16(c), as the number of street-view images increases, the R² value continuously rises, reaching its peak at 25.
-Therefore, our setting is the optimal choice.
-Appendix I Practicality
+We also present a practical deployment example of UrbanVLP in Appendix K.
+Figure 17: (a) FPS comparison of different models. (b) Comparison of the quality of generated texts of UrbanVLP and UrbanCLIP.
+Discussion on the Enhancements in Text Generation Quality by UrbanVLP
+As we discussed in the Introduction, the text generation process of UrbanCLIP raises several critical concerns: i) Hallucination and ii) Homogenization.
+To assess the comparative improvement of UrbanVLP in addressing these issues, we calculate three metrics for the generated text in both UrbanCLIP and UrbanVLP: (1) the average text length, (2) the mean text similarity, and (3) the average PerceptionScore.
+First of all, the average length of the text description in UrbanCLIP is quantified at 18.1 words.
+This relatively short length may lead to insufficient information being conveyed.
+In contrast, the average length of the text description in UrbanVLP is 144.1 words, nearly eight times longer, providing ample space to describe the content of the images in greater detail, thereby covering more nuanced information.
+Furthermore, we calculate the mean text similarity of all the generated texts of UrbanCLIP and UrbanVLP.
+As depicted in Figure 17 (b) on the left side, the mean text similarity of UrbanCLIP texts is 0.28, while that of UrbanVLP is 0.19.
+This indicates that our Automatic Text Generation and Calibration mechanism significantly mitigates the issue of homogenization, resulting in more diversified text descriptions for images with different content.
+Additionally, on the right side of Figure 17 (b), we evaluate the quality of generated descriptions through our proposed PerceptionScore.
+As we can see, the average scores of UrbanCLIP texts and UrbanVLP texts are 0.675 and 0.714, respectively.
+This indicates that UrbanVLP effectively mitigates the issue of inconsistency (hallucination) between generated text and visual content.
+Ablation study for fusion method
+To verify the effectiveness of our addition fusion method of satellite image features, street-view image features and location features, we experiment with different types fusion method, including (1) Concatenation: we directly concat the three kinds of features; (2) MLP: we feed the three features into a multi-layer perceptron (MLP) for fusion, resulting in a single output.
+As indicated in Table 5, the differences among several distinct fusion methods are relatively minor, which may be attributed to the overall complexity of the model, resulting in minimal impact from the various fusion techniques.
+Consequently, we opt for the most straightforward and high-performing approach, the addition method, for implementation.
+Table 5: Ablation study on $\mathtt{CityView}$-$\mathtt{Beijing}$ dataset. The best results are in bold.
+Model Carbon Population GDP
+$\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
+Concat 0.759 0.487 0.389 0.705 0.535 0.429 0.526 0.693 0.424
+MLP 0.761 0.483 0.374 0.710 0.527 0.415 0.532 0.688 0.418
+Addition (Ours) 0.769 0.477 0.369 0.714 0.523 0.411 0.537 0.684 0.416
+More Visualization of Region Representations
+Figure 18 provides a visualization with a population indicator, compared to UrbanCLIP.
+We use the dot color to indicate the population (log) in the corresponding regions of Beijing.
+As shown by the red dashed boxes, the regions modeled by UrbanVLP exhibit more significant clustering characteristics and greater distinction in high-dimensional space.
+Additionally, the lower parts of the two main band-shaped branches are also more distinctly separated.
+Points that are spatially close also have similar population values.
+Figure 18: PCA visualization of urban imagery.
+K    Practicality
 Figure 19: UrbanVLP system.
 To demonstrate the practical applications of our model in real-life scenarios, we have developed a system bearing the same name UrbanVLP.
-The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban indicator framework using the Mapbox platform [56].
-This system renders a visual representation of urban areas through satellite imagery, complemented by street-level photographs for an interactive exploration experience.
+The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban socioeconomic indicator framework using the Mapbox platform (Mapbox 2024).
 Figure 19 showcases the user-friendly interface of the map, highlighting functionalities that facilitate zooming, searching for locations, and navigating through various sectors.
-Each urban region captured by a satellite image, encompassing a 1km $\times$ 1km area, is marked by a blue dot, which is encircled by yellow points that denote the locations of street-view images accessible within that region.
+Each urban region, captured by a satellite image and covering a 1km $\times$ 1km area, is marked by a blue dot.
+Yellow points surround the blue dot, indicating the locations of street-view images accessible within that region.
 Engaging with these points allows users to uncover a suite of urban metrics, such as carbon emissions, population, GDP, night light, house price, and POI.
@@
 Ultimately, UrbanVLP is designed to distill complex urban data into a visual and user-friendly format, thereby offering a holistic and detailed view of urban landscapes and their key indicators.
-Appendix J Limitations and Social Impact
-Due to the easier accessibility of the dataset, our work primarily focuses on first-tier and domestic cities, which introduces a certain bias and can be regarded as a limitation of our study.
-Moreover, the model’s performance is significantly influenced by image resolution; higher quality image data is required to achieve optimal performance.
-Additionally, the introduction of new modalities inevitably leads to a decrease in inference efficiency, which should be a consideration for practical deployment.
-Moreover, the inability to automatically correct low-quality text highlights an area for future improvement.
-Our work can be deployed in practical applications, as demonstrated by the example provided in the paper, thereby aiding urban planning and decision-making processes.
+We have also prepared a demonstration video in the supplementary material to showcase the utility of our system.
+L    Limitations and Social Impact
+Due to the easier accessibility of the data, our work primarily focuses on Chinese first-tier cities, which introduces a certain bias and can be regarded as a limitation of our study.
+Moreover, the model’s performance is significantly influenced by image resolution.
+Higher quality image data is required to achieve optimal performance.
+Additionally, the inability to automatically correct low-quality text highlights an area for future improvement.
+Our work can be deployed in practical applications, as demonstrated by the example provided in the paper, thereby aiding in urban planning and decision-making processes.
 However, it is crucial to note that satellite and street-view images might inadvertently contain private information.
-The street-view images we collected have been blurred by the Baidu Map to obscure relevant information, such as license plates, thereby significantly reducing associated risks.
+The street-view images we collected have been blurred to obscure relevant information, such as license plates, thereby significantly reducing associated risks.
 Nevertheless, we strongly advise users to perform a comprehensive analysis of the data prior to deploying our model in practical contexts.